A Bibliography of Papers in Lecture Notes in Computer Science (2012): Volumes 7300–7349

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 1.03

Title word cross-reference

2 [551]. 3 [760, 555, 545, 535, 182, 879, 881, 755, 546, 450, 551, 476]. ^{+} [385]. ^{2} [405]. ^{n} [174]. ^{♯} [254]. a ∈ R^{N} [447]. α [895]. c [41]. ℓ^{2} [454]. k [442, 610, 58, 12, 723]. L^{2} [447]. L_{1/2} [585]. n [170, 799, 205]. P [907]. S(x, a) [447]. S_{R}(a) [447]. T [907]. u [665]. Z [174].

- [907]. -Automata [8]. -Certainty [41]. -Clique [58]. -gram-Based [205]. -Nearest [442]. -Petri [12]. -Reconstructibility [895]. -RPMS [665]. -Simple [177]. -Synchronous [799]. -Systems [907].

.NET [362].

/M [264]. /N [264].

101companies [82]. 10th [982, 961]. 11th [958, 983, 949]. 128/ [789]. 12th [962, 974, 975, 976, 977]. 13th [984]. 15th [960]. 160 [789]. 16th [944, 945]. 17th [978, 951]. 19th [957].

2012 [154]. 24th [979, 969]. 256-Bit [788]. 25th [986, 953]. 26th [956].

3-Manifolds [169]. 3.1 [239]. 33rd [988]. 3D [735].

4th [970, 952].

50th [947]. 5th [981, 950, 985].

65th [943]. 6th [964, 948, 968].

802.16 [268]. 8th [989, 955, 990].

9th [965, 966, 967, 963].

ABCD [737]. Ability [297]. Abnormalities [747]. Abstract [309, 959]. Abstracting [510]. Abstraction [674, 897, 401]. Abstractions [89, 375]. Abstracts [916]. ABZ [959]. Accelerated [703, 717]. Accelerating [233, 57]. Accelerator [232]. Acceptance [838]. Access [512, 845, 90, 414]. Accessibility [644]. Accounting [356]. Accumulators [786]. Accuracy [830]. Acetabular [551]. Achieve [542]. ACNS [982]. ACO [613]. ACO-Based [613]. Acoustic [766, 860]. Acquisition [831, 755]. Action [630, 800, 375]. Actions [194]. Active [23, 38, 41, 733, 318]. Activities [474, 687, 822, 52]. Activity [49, 291]. Acyclic [180]. Ad [869, 198]. Ada [951, 159, 158, 153, 151, 154, 155]. Adaptation [120, 374, 924, 199]. Adaptive [564, 241, 300, 532, 765, 581, 408]. Additive [591]. Additively [655]. Administration [828]. Advanced [962, 729, 750, 931, 232, 986, 969]. Advances [953, 954, 944, 945, 972, 973]. Advantages [34]. Advertisement [54]. Advice [2]. Affect [443, 838, 296, 289]. Affective [292, 293]. Affinity [223]. Agadir [981]. Again [720]. against [844]. Agent [490, 290, 968, 481, 494, 496, 483, 968]. Agents [486, 288]. Aggregate [250]. Aggregating [505]. Aggressive [27]. Agile [824, 823]. Agility [821]. Agreements [487]. Agriculture [662]. Ahp [637]. AI [953]. Aided [794, 341]. Algebra [281, 384]. Algebraic [806]. Algorithm [652, 598, 607, 427, 566, 436, 608, 656, 739, 614, 526, 27, 331, 729, 869, 570, 580, 585, 603, 579, 12, 741, 174, 611, 891, 581, 648, 144, 345, 569, 565, 780]. Algorithms [856, 884, 904, 801, 899, 210, 965]. Aliasing [253]. Aligning [515]. Alignment [717]. Alignments [537]. Alive [943]. Allocation [434, 484, 605]. Allowing [791]. Alloy [959, 313, 314, 315, 316, 317, 318]. ALM [240]. Alternating [452]. Am [645]. Ambient [964]. Ambiguous [699]. among [676]. Amplification [785]. AMSTA [968]. Analysing [389]. Analysis [676, 455, 186, 443, 634, 377, 307, 383, 471, 272, 266, 966, 967, 880, 378, 468, 249, 674, 659, 622, 641, 773, 47, 324, 258, 530, 393, 881, 746, 750, 876, 349, 787, 717, 300, 870, 467, 491, 632, 240, 813, 268, 67, 236, 669, 262, 277, 69, 690, 108, 642, 205, 497, 601, 390]. Analytical [280, 957]. Analyze [79]. Analyzed [645]. Analyzing [846, 823]. Android [849, 840, 421]. AndroidLeaks [849]. ANFIS [580]. Angle [647, 200]. Animated [774]. Annotation [911, 596]. Annotations [511]. Anonymity [834]. ANSI [392]. ANSI-C [392]. Answering [913, 600]. Ant [611]. Anti [440, 394]. Anti-Bayesian [440]. Anti-patterns [394]. Antipatterns [378]. Anytime [729]. API [156]. Apollonian [431]. App [665]. Application [307, 383, 48, 631, 659, 83, 479, 72, 737, 238, 647, 268, 663, 661, 68, 858, 988, 67]. Application-Level [68]. Applications [957, 978, 119, 782, 411, 960, 279, 513, 849, 475, 653, 968, 986, 233, 521, 503, 527, 974, 975, 976, 977, 152, 609, 236, 606, 990, 835, 784, 987]. Applicative [808]. Applied [81, 769, 986, 982]. Applying [186, 200]. Approach [670, 520, 274, 629, 756, 617, 158, 816, 660, 903, 72, 387, 932, 746, 770, 29, 478, 870, 453, 298, 745, 636, 445, 679, 695, 228, 595, 205, 590, 675, 510]. Approaches [213, 925]. Approximation [181]. AR-Tactile [359]. Arabic [455]. Arbitrary [178, 728]. Archetypes [151]. Architecture [231, 825, 835]. Architectures [116, 906, 816, 226, 240, 235, 677]. Archive [571]. Areas [754, 633]. Arguments [87]. Arithmetic [314]. Art [77, 4, 329, 809]. Artifact [154]. Artificial [652, 617, 986, 953, 651, 611, 525, 518]. Asia [944, 945]. ASM [310, 311, 312]. ASMTA [957]. Aspect [120, 683, 933, 935]. Aspect-Oriented [120, 935]. ASRS [656]. asse [635]. Assessing [638, 297, 235, 677]. Assessment [138, 453, 748]. Assignment [869]. Assisted [971, 230, 614, 72, 95, 740]. Assisting [416, 742]. Association [191, 46]. Associative [706, 617, 540, 525]. Assumption [469]. Assumptions [346]. Astronomical [719]. Asymptotic [279, 16]. Asynchronous [254, 119]. Atelier [323]. Atlas [557]. Atlas-Based [557]. Attachment [426]. Attack [848, 418, 39]. Attacking [318]. Attacks [775, 790]. Attention [430, 300]. Attested [840]. attrezzato [635]. Attribute [879]. Attributed [58]. Attributes [607, 855, 493]. Auditable [834]. Auditory [209]. Augmenting [160]. August [987]. Australia [954]. Austria [985]. Authenticated [778, 833, 773]. Authentication [407, 841]. Auto [224, 609]. Auto-scoping [224]. Auto-tuning [609]. AutoFutures [71]. Automata [807, 278, 5, 8, 13, 172, 19]. Automated [100, 805, 939, 671, 814, 113, 323, 124, 642]. Automatic [389, 739, 526, 854, 860, 737, 890, 546, 71, 741, 365, 600, 584, 525, 228, 305, 586]. Automatically [849, 357]. Automation [885]. Automotive [516, 923, 165]. Aveiro [966, 967]. Averaged [705]. Aware [143, 120, 521, 361]. AWESOM [365]. Axiomatic [326]. Axis [619, 621, 733, 740]. Axis-Parallel [619].

B [959, 319, 398, 159, 385, 118, 323, 161, 399, 397]. B&B [655, 609]. Back [20, 75, 114, 128, 149, 166, 184, 220, 284, 381, 422, 437, 561, 810, 852, 886, 920, 264]. Backdoors [334]. Bacterial [566]. Badger [934]. Bag [205]. Bag-of-Words [205]. Bahia [974, 975, 976, 977]. Balanced [265]. Balancing [579]. Ban [201]. Band [771]. Bands [386]. Bank [455]. Banner [54]. Based [237, 186, 612, 805, 549, 760, 939, 388, 295, 929, 109, 848, 119, 102, 768, 782, 844, 490, 670, 24, 556, 472, 417, 897, 138, 564, 624, 27, 781, 640, 903, 653, 550, 658, 407, 593, 575, 860, 441, 737, 707, 195, 393, 762, 841, 926, 746, 750, 473, 869, 489, 700, 571, 570, 603, 579, 104, 759, 544, 157, 464, 466, 764, 904, 847, 467, 198, 855, 151, 52, 623, 600, 651, 154, 304, 236, 366, 918, 445, 201, 37, 613, 891, 587, 663, 581, 868, 216, 397, 576, 865, 606, 599, 925, 574, 148, 205, 592, 557]. Based [345, 858, 602, 594, 569, 857, 565, 601, 675, 591, 439, 790, 817, 673]. Basics [371]. Basis [647]. Bass [860]. Batch [858]. Battery [280, 350]. Battery-Shaped [350]. Bayesian [752, 197, 485, 440, 857]. Beam [647]. Beating [552]. Bee [611]. Behaved [570, 580]. Behavior [687, 850, 349, 31, 67, 95, 353]. Behavioral [351]. Behaviour [888, 482, 397]. Behaviour-Based [397]. Beijing [956]. Being [354]. Belief [846]. Benchmark [238]. Benefits [504, 642]. BEP [634]. Berkeley [587]. Bertinoro [962, 952]. between [190, 367, 665, 770, 822, 162, 609, 202, 293, 933]. beyond [789, 868]. Bi [432]. Bi-directional [432]. Bidirectional [136, 135]. Big [308]. Big-Step [308]. Bigrams [495]. Bijective [769]. Binarization [764]. Binary [859, 623]. Biological [563]. Biologically [598, 448]. Biomass [649]. Biomedical [917, 746, 29, 915, 477]. Biomedicine [980]. Biometric [619, 468, 461]. Bioprosthetic [750]. Biosignals [750]. BIP [116]. Bipartite [410]. Birthday [943]. Bisection [658]. Bit [788]. Blasts [770]. Blending [265]. Blind [769, 451]. Blob [475]. Block [787, 791, 780]. Blogging [497, 720]. Blood [731]. Blueprints [357]. Body [867]. Boolean [338, 9, 335]. Boost [915]. Border [570]. Boruvka [610]. Both [229]. Botnet [408]. Boundaries [875]. Boundary [577]. Bounded [101, 893, 392]. Bounded-Exhaustive [101]. Bounds [655]. BPM [673]. BPMN [117]. Brain [734, 216, 557]. Brain-Machine [216]. Branching [281]. Brazil [974, 975, 976, 977]. Break [330, 409]. Breaking [538, 774]. Bridging [100, 933]. Brisbane [954]. Browser [312]. BT* [729]. Bubbles [411]. Buffer [267, 799]. Bug [96]. Bugs [392]. Building [32, 661]. Burn [742]. Business [127, 515, 504, 516, 138, 506, 627, 673].

C [254, 736, 405, 87, 91, 392, 891, 819]. C# [81]. C-Means [736]. C-Nets [891]. CAC [269]. CAiSE [969]. Calculating [801]. Calculi [795]. Calculus [281, 496, 214, 806, 868]. Calculus-Based [868]. Calibrated [551]. Calibration [548, 364]. Call [264, 265]. Call-Back [264]. Camera [472, 622, 473, 757, 467]. Camera-Projector [622]. Can [778, 842]. Canada [965, 953]. Canadian [953]. Cancellation [227]. Cancer [914]. Candidate [916]. Capability [499]. Capacities [272]. Capacity [273, 903, 387]. CAPTCHA [774]. Capturing [67]. Capuano [638]. Cardiac [553, 554].

Carlo [527]. Carte [145]. Cascade [736]. Cascading [685]. Case [672, 225, 384, 556, 927, 516, 889, 637, 84, 160, 828, 632, 154, 918, 930, 865, 933, 626]. Case-Based [918, 865]. Cases [694]. Cast [441]. Castel [638]. Catalog [682]. Categorical [296]. Categorization [706, 38]. Catheterization [554]. CD [8]. CD-Systems [8]. CDCL [331]. Cell [178, 47, 280]. Cellular [172]. Center [264]. Centered [411, 838]. Central [645]. Centric [639, 368, 421]. Centroid [707]. Centroid-Based [707]. Certain [431]. Certainty [41]. Certification [815]. Certified [674, 173]. CFD [233, 661]. Chain [671, 276, 435]. Chains [649, 702]. Challenges [121, 503, 923, 483]. Chance [597]. Chance-Constrained [597]. Change [866]. Chania [979, 958]. Channel [768, 869, 771, 791, 767]. Chaos [570]. Character [538]. Characterization [768, 766, 771]. Charts [92]. Cheat [397]. Check [455]. Checking [100, 78, 901, 158, 251, 99, 392, 317, 390]. Children [351]. China [986, 956, 972, 973]. Chinese [602, 601]. Chips [613]. CHOC’LATE [162]. Choice [629]. Choices [162]. Choosing [330]. Chord [860]. Chronic [742]. Cipher [791]. Ciphers [787, 780, 779]. Circle [593]. Circum [593]. City [354, 645]. CK [686]. Class [266, 147, 40, 201, 37, 32, 39]. Class-Imbalanced [40]. Classes [443, 686]. Classical [19]. Classification [442, 443, 188, 754, 24, 756, 617, 707, 729, 23, 855, 162, 22, 445, 201, 525, 560, 440, 30]. Classification-Based [445]. Classification-Tree [162]. Classifier [593, 29, 31]. Classifier-Ensemble [29]. Classifiers [542, 39]. Classify [96]. Classifying [708, 34]. Clause [343, 332]. Clean [130]. Clearing [368]. Clinical [917, 919]. Clique [58]. Clockless [839]. Close [726]. Closed [548]. Closed-Form [548]. Closing [389]. Cloud [256, 816, 567, 579, 500, 591]. Clouds [840]. Cluster [47, 632]. Cluster KDD [599]. Clustering [598, 736, 446, 728, 856, 575, 633, 201, 695, 878, 599, 493, 594, 601, 439]. Clusters [444, 432, 541]. Co [290, 47, 142, 857]. Co-evolution [142]. Co-occurrence [857]. Co-Occurring [290, 47]. Coalescent [890]. Code [98, 247, 325, 151, 154, 819, 219]. Coded [532]. Codes [120]. Coevolutionary [573]. Cogeneration [657]. Cognition [215]. Cognition-Inspired [215]. Cognitive [297, 684, 491]. Cohesion [653]. Cohort [919]. Coil [363]. Collaborative [288, 603, 445]. Collapse [538]. Collect [746]. Collection [64]. Collections [259, 926, 58, 723]. College [989]. Collisions [792]. Colony [611]. Color [752, 755, 450]. Colored [890, 904]. Coloring [424]. Coloured [903]. Combating [405]. Combination [823]. Combinatorial [341]. Combined [269, 6, 228]. Combining [913, 929, 325, 99, 469]. Combustion [659]. Commercial [500]. Commodity [835, 434]. Common [355]. Communication [50, 962, 123, 355, 489, 265, 825, 615]. Communities [497]. Community [186, 82, 354, 45, 497]. Commutative [924]. Companies [823]. Company [824, 815]. Comparative [575, 925]. Compare [696]. Comparing [162]. Comparison [500, 68]. Compensation [549, 553]. Competition [481]. Competitive [481]. Compile [88]. Compile-Time [88]. Compiler [230, 233, 391, 234, 228]. Compiler-Assisted [230]. Compilers [243].

Complaint [602]. Complete [889, 869, 316]. Completeness [183]. Complex [698, 492, 216]. Complexes [178]. Complexities [15]. Complexity [685, 893, 684, 13, 16, 648, 18]. Compliance [201]. Component [157, 151, 345]. Component-Based [157, 151]. Components [947, 400, 58, 96]. Composite [531]. Composition [121, 120, 949, 117, 509, 924, 507]. Compositions [487]. Compound [474]. Comprehensions [326]. Comprehensive [201, 601]. Compression [179, 744, 789]. Computation [211, 2, 416, 170, 276, 415, 173, 212, 644]. Computational [639, 66, 954, 175, 952, 974, 975, 976, 977]. Computations [427, 180]. Compute [174]. Computer [971, 794, 962, 295, 329, 740]. Computer-Aided [794]. Computer-Assisted [971]. Computer-Based [295]. Computers [1]. Computing [964, 833, 72, 985, 961, 579, 361, 14, 837, 333, 834]. Concealment [532]. Concentrated [176]. Concept [641, 467, 599]. Concepts [155, 62]. Conceptual [689]. Concerns [157, 507, 851]. Concise [48]. Concrete [745]. Concurrency [803, 312, 89]. Concurrent [65, 888, 395]. Condition [748]. Conference [971, 979, 957, 982, 989, 978, 951, 948, 966, 967, 970, 958, 960, 959, 963, 984, 981, 947, 983, 949, 988, 950, 968, 986, 985, 961, 953, 974, 975, 976, 977, 956, 946, 980, 969, 944, 945, 972, 973, 990]. Confidence [855]. Confidence-Based [855]. Configurable [514]. Configuration [187, 939, 515]. Confinement [94]. Conflict [507]. Conflict-Free [507]. Confluent [898]. Conformal [544]. Conformance [103, 105]. Congress [954]. Congruence [826]. Congruency [764]. Connect [863, 875]. Connected [666, 722]. Connection [429]. Connectivity [915]. conscious [297]. Considerations [658]. Consistency [256, 446]. Consolidation [271]. Constrained [833, 478, 894, 866, 651, 597]. Constraint [372, 102, 146, 766, 165, 107, 204, 342]. Constraint-Based [102]. Constraint-Driven [146]. Constraints [93]. Construct [890]. Constructing [928, 808, 122, 780]. Construction [610, 417, 983]. Constructive [108, 872]. Constructs [226]. Contact [539]. Content [411]. Context [639, 515, 120, 674, 4, 952, 368, 521, 11, 301, 52, 825]. Context-Aware [120]. Context-Supported [368]. Contextual [775, 448]. Continuous [311, 715]. Contour [733, 875]. Contourlet [762]. Contours [621]. Contrast [449]. Contrasts [557]. Contribution [307, 383]. Contributions [345]. Control [411, 845, 164, 313, 163, 74, 650, 750, 400, 414, 247, 668, 216, 219]. Controlled [12, 17, 18]. Controller [310]. Controllers [800]. Controlling [275]. Converse [841]. Cooperating [278, 5, 6]. Cooperation [303, 609]. Cooperative [509, 216]. Coordination [496, 119]. Cope [324]. Core [231, 627, 716, 240, 922, 68]. Correct [134]. Correction [450]. Correctness [804, 686]. Corrector [648]. Correlated [48]. Correlation [531, 266, 768, 418, 262]. Coscheduling [70]. Cosine [859, 744]. COSMIC [677]. Cost [192, 894, 904, 817]. Cost-Effective [904]. Cost-Sensitive [192]. Costs [504]. COTS [653]. Could [692]. Counter [392].

Counter-Examples [392]. Counterexample [337]. Coupling [653]. Coverability [899]. Coverage [101, 389, 666, 54]. Cox [585]. CPMD [866]. Creating [207]. Crete [979, 958]. Crime [645]. Crisis [486]. Criteria [101, 643, 737, 440]. Croatia [968]. Croatian [495]. Cross [188, 911, 44, 719, 38]. Cross-Identification [719]. Cross-Roads [44]. Crowd [368]. Crowding [571]. Crowdsourcing [916]. Crown [529]. Cryptanalysis [788, 419]. Cryptographic [156, 418]. Cryptography [982, 420]. Cryptosystem [790]. CSMA [270]. CSMA/CA [270]. CSP [387, 118]. CT [741]. CTIC [952]. CUDA [234]. Cues [752, 289]. Curvature [176, 877]. Curve [640, 420, 444]. Custom [362]. Customization [86]. Customized [207, 122]. Cyber [405, 863]. Cyber-Physical [863]. Cyber-Threat [405]. Cytological [739]. Czech [948, 947, 949, 950, 946].

D [760, 555, 545, 170, 535, 182, 879, 881, 755, 546, 450, 551, 476]. DAG [425]. Dalian [986]. Damage [47, 565]. Damaged [532]. Dan [809]. Dangerous [247]. Dassow [943]. Data [929, 229, 119, 917, 897, 641, 355, 728, 726, 218, 881, 40, 33, 60, 517, 755, 710, 57, 43, 644, 491, 600, 813, 498, 783, 421, 107, 914, 878, 944, 945, 718, 493, 713, 34, 989]. Data-Centric [421]. Data-Flow [929, 119]. Data-Intensive [498]. Database [979, 512, 806]. Databases [712, 680, 719, 461]. Dataset [542]. DB [587]. Deal [326]. Dealing [676]. Deblocking [765]. Debugging [110, 88]. Decision [127, 630, 629, 758, 32, 505]. Decision-Making [630, 505]. Declarative [519, 246]. Declassification [797]. Decomposing [892]. Decompositions [698]. Decoupling [274]. Dedicated [943]. Deep [557]. Defect [813]. Defensive [484]. Defined [421]. Defining [117]. Definition [622]. Definitive [372]. Degraded [633]. Degree [436]. Degrees [428]. Delaunay [618]. Delegable [845]. Deletion [177]. Delta [104]. Delta-Oriented [104]. Demand [207]. Demands [434]. Demographic [890]. Denmark [990]. Denoising [760, 476]. Dense [723]. Dependencies [676, 704, 333]. Dependency [35, 69]. Dependently [805]. Depth [471, 472]. Derivation [816, 6]. Derivations [11]. Derivative [654]. Derivative-Free [654]. Derive [584]. Derived [931]. Deriving [800]. Description [183, 499]. Descriptional [13]. Descriptions [326]. Design [962, 671, 102, 649, 223, 816, 660, 321, 359, 923, 873, 742, 303, 248, 934, 600, 525, 69, 661]. Designing [905, 156, 344, 165]. Desk [828]. Desktop [66]. Detailed [854]. Detecting [849, 856, 60, 350, 449]. Detection [531, 428, 747, 48, 44, 739, 475, 731, 452, 401, 296, 441, 737, 732, 717, 57, 546, 94, 678, 743, 457, 529, 397, 451, 408, 45, 858, 565, 289]. Detectors [469]. Determination [859, 745]. Determine [681, 488]. Determining [444]. Deterministic [8]. Develop [98]. Developing [696]. Development [388, 818, 79, 83, 637, 559, 125, 160, 665, 822, 157, 358, 915, 151, 161, 936, 823, 154, 95, 825]. Developments [404]. Device [363, 837]. Devices [420, 357, 350, 663, 842, 362]. Diagnosis [742]. Diagrams [929, 109].

Difference [190]. Differences [822, 37]. Different [575, 68]. Differential [419, 548]. Digital [880, 173, 876, 884, 172, 137]. DILS [989]. Dimension [619, 582]. Dimensional [589, 296, 690]. Dimensionality [443, 753]. Dimensions [178]. Directed [474, 340]. Direction [452]. Directional [738, 432]. Directives [233]. Disambiguation [701]. Discharging [323]. Discovering [55, 54, 718]. Discovery [217, 341, 519, 866, 909, 56, 891, 944, 945, 599, 723]. Discrete [169, 272, 566, 882, 987, 877, 365]. Discrete-Time [272]. Discriminant [470]. Discrimination [770]. Disease [885]. Diseases [748]. Disparate [915]. Display [753]. Disruptions [597]. Dissimilarity [442]. Distance [181, 880, 571]. Distinguishers [789, 780]. Distorted [531]. Distortion [444]. Distributed [612, 272, 413, 674, 5, 153, 6, 282, 666, 840, 825, 483, 718]. Distribution [589, 276, 351]. Distributions [330, 771]. Diverse [542]. DM6437EVM [765]. Do [245]. Document [706, 764, 695]. Documenting [669]. Documents [932, 690, 602, 591]. Domain [126, 279, 454, 928, 830, 738, 499, 199, 925, 594, 941, 137, 922]. Domain-Specific [126, 928, 594, 941, 137]. Dominance [801]. Don’t [850]. Door [56]. Dots [875]. Double [695]. Doubling [779]. Dresses [626]. Drifts [43]. Driven [962, 384, 136, 146, 654, 132, 157, 13, 609, 95, 941]. Driver [360]. Drivers [823]. Driving [127, 349]. DroidSense [79]. DSA [153]. DSL [371, 261]. DSLs [140]. DSmT [758]. DSP [765]. Dual [837]. Dual-OS [837]. Dualities [203]. Dubrovnik [968]. during [295, 818, 302, 290, 296, 165, 245, 219]. Dy [116]. Dy-BIP [116]. Dynamic [116, 434, 869, 433, 679, 723]. Dynamics [280].

e-Learning [587]. e-Remanufacturing [568]. EAP [409]. EAP-MD5 [409]. Earley [12]. Early [671, 731, 401, 741, 500]. Earthquakes [770]. ECG [747, 295]. ECMFA [990]. ECOOP [956]. Ecosystems [964]. EDF [164]. Edge [762]. Edges [449]. EDI [516]. Educational [300]. EEG [749]. Effect [612, 635, 700, 301]. Effective [904]. Effectiveness [298, 498]. Effects [302, 288, 820, 255, 234]. Efficiency [785, 863, 724]. Efficient [188, 77, 229, 720, 714, 170, 417, 727, 415, 420, 282, 859, 519, 890, 197, 346, 174, 678, 791, 165, 26, 86, 780, 725]. Effort [830, 679]. Egham [964]. Eigenspace [474]. EIMOS [664]. Eindhoven [798]. Elastic [577]. Elder [360]. Elderly [472]. Electric [74]. Electrical [350, 363]. Electron [881]. Electronic [74]. Elementary [798]. Elements [298]. Elicitation [816]. Elliptic [420]. ELM [200]. Embedded [939, 589, 261, 761, 837, 418]. Embedding [470, 400]. Emendation [937]. Emerge [487]. Emerging [775, 349]. EMF [77, 931, 144, 142]. Emotional [290, 196]. Emotions [294]. Empirical [138, 193, 826]. Enabled [845, 420]. Enabling [555]. Encodings [193]. Encrypted [591]. Encryption [778, 785, 833, 591]. Endogenous [134]. Endoscopic [479, 547]. Energy [634, 356, 49]. Energy-Reliant [356]. Enforcement [421]. Engagement [294, 298]. Engine [719, 909, 693, 679]. Engineering [676, 672, 962, 812, 828, 986, 873, 946, 969, 925, 209, 941].

Engines [903]. English [586]. Enhance [297, 518]. Enhanced [555, 608]. Enhancement [755, 759, 627]. Enhancements [781]. Enhancing [664, 260, 85, 732, 298, 171]. Ensemble [25, 29, 43, 542, 22]. Ensembles [207]. Enterprise [502]. Entities [3]. Entity [705, 29]. Entrez [909]. Entropy [468]. Entry [250]. Envelopes [834]. Environment [653, 567, 579, 361, 827, 936, 298]. Environment-Aware [361]. Environments [351, 484, 304, 500]. Equation [279, 716]. Equations [7, 391]. Equivalent [788]. ERP [681]. Error [88]. Escape [612]. Escrow [776]. Essays [943]. Establishing [829]. Estimates [830]. Estimating [818, 367, 724]. Estimation [883, 553, 176, 877, 757, 679, 817]. Euclidean [786]. Europe [951]. European [956, 990, 645]. Evacuation [486]. Evaluating [110, 248, 713, 345]. Evaluation [81, 637, 182, 488, 222, 458, 199, 560, 819, 601]. Evaluations [558]. Evaluator [402]. Evaluators [122]. Evasion [39]. Event [319, 384, 398, 159, 928, 871, 49, 519, 161, 69, 399, 397, 105, 518]. Event-B [319, 398, 159, 161, 399, 397]. Event-Driven [384]. Events [350]. Eventual [256]. Everyday [356]. Evolution [247, 245, 142]. Evolutionary [211, 212, 213, 210]. Evolved [24]. Evolving [218, 433, 22]. Exact [48]. examining [331]. Examples [392]. Exchange [343, 803, 773]. Exclusion [935]. Executable [325, 922]. Execution [237, 163, 687, 99, 842]. Exhaustive [101]. Expandable [588]. Expansion [697]. Expansions [584]. Experience [302, 830]. Experiences [829]. Experimental [471]. Experimenting [836]. Experiments [231]. Expert [830]. Experts [453]. Explain [809]. Explicit [463, 85]. Exploiting [197]. Exploration [289]. Explorative [642]. Exploratory [294]. Exploring [302, 293]. Exponential [440, 572]. Expression [467]. Expressions [110]. Extended [528, 335, 675]. Extending [123, 326, 226, 315, 780]. Extensibility [244]. Extension [258, 246]. Extensions [880, 809, 245]. External [681, 571]. Extra [85]. Extra-Functional [85]. Extractable [782]. Extracting [695]. Extraction [455, 595, 394, 602]. Eye [732].

Face [463, 468, 470, 623, 469]. Face-Part [469]. Facebook [851]. Facial [854, 467, 465]. Fact [148]. Factor [125, 773]. Factorization [703]. Factors [299, 205]. Factory [364]. Failed [335]. Failed-Literal [335]. Failures [283]. Fair [487]. Fake [87, 691]. Fall [472, 534]. Fallen [407]. Family [141, 440]. FAQ [700]. Farming [865]. Fast [610, 78, 436, 836, 556, 883, 824, 584]. Fast-Checking [78]. Fault [401, 787]. Faulty [342]. FCFS [266]. Fear [645]. Feature [43, 94, 743, 448, 642, 547, 439]. Features [754, 854, 860, 931, 594]. Feed [408]. Feed-Forward [408]. Feedback [750, 277, 606]. Feedback-Based [606]. Few [126]. Fibers [745]. Fiction [148]. Field [351, 277]. Filter [654, 651, 765]. Filter-Based [651]. Filtering [720, 879, 603, 477, 857]. Filters [531, 928]. Find [436]. Finding [911, 58, 283]. Findings [738]. Fine [421]. Fine-Grained [421]. Fingerprint [537, 466]. Fingerprinting [365]. Finite [273, 267, 108, 19].

Finite-Buffer [267]. Finnish [828]. Fire [485]. Firefly [739]. First [802, 102, 987]. First-Past-the-Post [802]. Fish [652, 651, 865]. Fish4Knowledge [499]. Fixing [763]. Fixtures [559]. Flexibility [127]. Flock [598]. Flooded [720]. Flow [929, 229, 119, 674, 733]. Fly [937, 901]. Flying [216]. FocalTest [102]. Focus [371, 456]. Focused [984]. Folding [339]. Folksonomies [62]. Following [540]. Food [602]. Footprint [259]. Footprints [896]. Forced [269]. Forces [733]. Forecast [716]. Forecasting [25]. Foreground [452]. Forensics [137]. Forest [485]. Forests [179, 33, 34]. Fork [229]. Fork-Join [229]. Form [548, 699]. Formal [962, 389, 78, 963, 159, 321, 160, 9, 391, 161, 203, 399, 380, 962]. Formalism [907]. Formalizing [319, 254]. Formally [102]. Formula [9]. Formulae [338]. Formulas [111, 335]. Formulation [548]. Forward [408]. Fostering [824]. Foundation [40]. Foundations [990]. FPGA [761]. Fragment [311]. Frames [919]. Framework [558, 111, 326, 249, 83, 292, 258, 395, 884, 596, 168, 31, 152, 232, 216, 353, 460, 497]. Frameworks [307, 383]. France [957]. Free [654, 507]. Frequent [712, 862, 62]. Front [21, 42, 522, 543, 562, 583, 604, 625, 646, 667, 688, 709, 730, 751, 772, 793, 811, 832, 853, 874, 887, 908, 921, 942, 63, 76, 97, 115, 129, 150, 167, 185, 206, 221, 242, 263, 285, 306, 327, 348, 369, 382, 403, 423, 438, 459, 480, 501]. Fuel [47]. Full [550]. Fully [502]. Function [447, 49]. Functional [676, 694, 85, 686, 377, 507, 933]. Functions [169, 839, 647, 789, 784]. Functors [808]. Fusion [758, 762, 858]. Fusion-Based [858]. Fusion/ Classification [758]. Future [503, 502, 283]. Fuzzy [735, 631, 736, 214, 629, 770, 508, 215, 534, 216]. Fypercomputations [14].

Gadgeteer [362]. Gains [287]. Gait [471]. Galleries [708]. Game [302, 430, 869, 303, 300, 301, 298, 304, 122]. Game-Based [304]. Game-Like [298]. Game-Theoretic [430]. Games [802, 424, 24, 494, 397]. Gamma [771]. Gap [100, 933]. Gated [268]. Gaussian [736, 450, 469]. Gdansk [969]. Gender [471, 574, 465]. Gender-Hierarchy [574]. Gene [916]. Gene-Mutation [916]. General [70]. Generalized [170, 470, 4, 340, 771]. Generated [357]. Generating [704, 506, 435, 144]. Generation [309, 389, 106, 694, 132, 814, 282, 107, 518]. Generic [884, 29]. GeneTegra [914]. Genetic [656, 614, 539]. Genome [911, 717]. Genre [707]. Geographic [639, 641, 697]. Geographical [643]. Geometric [881]. Geometrically [531, 272]. Geometry [622, 796, 987, 464]. GeoSocial [711]. Geospatial [596]. Geostatistics [641]. Geovisualization [641]. Germany [988]. Gesture [867, 560]. Get [720]. Getting [196]. GIS [642]. GIS-integrated [642]. Give [842]. Glaucoma [731]. Gliwice [980]. Global [654, 658, 651, 611, 606]. Globally [825]. GMM [623]. Go [912, 868]. Goal [499, 161]. Goal-Oriented [161]. Going [353]. Going-Out [353]. Good [698, 204]. GPGPU [616, 606]. GPS [352]. GPU [610, 607, 624, 883, 703, 717]. GPU-Accelerated [703, 717]. GPU-Based [624]. GPUs [233]. Gradient [556, 733, 465]. Grained [421]. gram [205].

Grammar [6]. Grammars [4, 11, 12, 17, 18]. Graph [965, 429, 374, 73, 412, 801, 723, 51, 439, 725]. Graph-Based [439]. Graphical [213]. Graphics [57]. Graphs [610, 911, 436, 721, 862, 58, 435, 724, 410]. Gray [601]. Gray-Clustering [601]. Greece [979, 958]. Grenoble [957]. Grey [569]. Grey-Based [569]. Grid [616]. Groningen [978]. Group [36, 193]. Grouping [524, 857]. Groups [173]. Growing [883, 824, 445, 821]. Growth [712]. GSD [827]. GSR [295]. Guidance [555, 360, 552]. Guide [372]. Guided [939, 337, 554]. GWT [587].

H264 [765]. H264/AVC [765]. Habitat [358]. Hacking [358]. Haemodynamic [748]. Hairpin [10]. Halifax [965]. Hamburg [988]. Hand [475, 750, 473, 363, 464, 536]. Hand-Geometry [464]. Hand-Sign [536]. Hand-Worn [363]. Handheld [350, 363, 361]. Handling [157, 165, 574]. Handoff [269]. Haralick [754]. Hardware [775, 225, 163]. Hash [782, 407]. Hash-Based [407]. Hazard [602]. Hazards [585]. HCBRG [445]. Health [189, 917]. HeapMotiv [302]. Heart [552, 748]. Henkin [338]. Heterogeneous [59, 955, 230, 43, 840]. Heuristic [538, 473]. Heuristically [542]. Hidden [354, 890, 826, 560]. Hierarchical [23, 551, 445, 878, 493, 51]. Hierarchy [574]. High [754, 616, 582, 436, 622]. High-Dimension [582]. Higher [791, 780]. Higher-Order [791, 780]. Hints [300]. HiPoLDS [413]. Histories [247]. History [412]. Hoare [794]. Hoc [869, 198]. Home [356, 645]. Homes [357, 358]. Homogeneous [396, 743]. Homological [179]. Homology [180, 170, 182, 173]. Horst [658]. Housing [628]. HRT [731]. HTTP [408]. Hu [592]. Huatulco [970]. Human [471, 190, 639, 474, 533, 534, 209]. Human-Centric [639]. Humans [489]. Hybrid [319, 901, 665, 349, 240, 366, 611, 34, 51, 497]. Hypergraph [424]. HyRAL [788].

I/O [222, 69]. i2b2 [918]. IC3 [328]. ICCSA [974, 975, 976, 977]. ICIAR [966, 967]. ICISP [981]. ICM [758]. ICMT [950]. ICSI [972, 973]. Identification [919, 719, 781, 642, 342]. Identifying [432, 89, 673]. Identities [776]. Identity [463, 782, 781]. Identity-Based [782, 781]. IDS [44]. IEA [986]. IEA/AIE [986]. IEEE [954, 268]. IFIP [964]. IFM [963]. II [225, 967, 975, 945, 973]. III [976]. IKEA [632]. Illuminant [757]. Image [752, 952, 526, 456, 885, 550, 762, 708, 354, 884, 554, 527, 453, 745, 458, 202, 878, 592, 966, 967, 981]. Image-Guided [554]. Image-Mining [885]. Images [531, 754, 479, 739, 454, 731, 173, 732, 753, 168, 449, 741, 172, 457, 532, 15, 551, 734, 476, 477]. Imaging [744, 757]. Imbalance [32]. Imbalanced [40]. Immune [617]. Impact [676, 778, 830, 686, 147]. Impedance [748]. Implementation [719, 420, 89, 600, 821]. Implementations [159, 235]. Implementing [164]. Implicit [286]. Important [205]. Impossible [419]. Imprecise [713]. Improve [463, 706, 697, 918, 539]. Improved [332, 571, 581, 576, 601, 518].

Improvement [984, 814, 815, 813]. Improvements [276]. Improves [448]. Improving [163, 559, 812, 828, 537, 872]. Improvisation [289]. Impulse [477]. Impulse-Noise [477]. In-the-Wild [358]. Including [225]. Incomplete [484]. Inconsistencies [934]. Incorporating [488, 647, 557]. Increasing [465]. Incremental [564, 27, 707, 104, 347, 855, 298, 37]. Increments [442]. Indentations [456]. Independent [85, 797]. Individual [505]. Indoor [768, 365, 366]. Induced [722]. Industrial [986, 930, 925, 933]. Industry [401]. Inertia [572]. Inference [848, 257, 251, 890, 197]. Infinite [273]. Inflective [495]. Influence [624, 68]. Influential [53]. Information [463, 189, 471, 978, 639, 53, 917, 545, 521, 869, 484, 702, 697, 969, 918, 171, 51, 602, 971, 964, 980]. Informativeness [495]. Informed [321]. Infrared [752, 460, 615]. Infrastructural [635]. Infrastructure [904, 864]. Infrastructures [726]. Initialization [252]. Inner [784]. Inner-Product [784]. Innovation [824, 626]. Innovative [906]. Input [13, 813]. Input-Driven [13]. Inspection [864]. Inspired [598, 215, 566]. Instance [520, 36]. Instances [315]. Instantiation [322]. Institutes [917]. Instruction [606]. Instrumentations [121]. Insults [287]. Insurance [240]. Integral [780]. Integrate [636]. Integrated [308, 231, 152, 642, 963]. Integrating [77, 153, 717, 915, 931, 493]. Integration [463, 989, 929, 917, 860, 662, 492, 863, 914]. Integrity [843]. Intel [231]. Intelligence [986, 953, 567, 954, 972, 973, 568]. Intelligent [612, 964, 294, 986, 286, 488, 600, 291, 958]. Intel(R) [231]. Intensive [498]. Inter [516, 910]. Inter-Ontology [910]. Inter-organizational [516]. Interaction [292, 290, 288, 563, 644, 590, 342]. Interactions [302, 539]. Interactive [398, 351, 361]. Interactivity [664]. Interchange [927]. Interclass [266]. Interesting [55]. Interface [322]. Interfaces [222]. interference [92]. Internal [140, 928]. International [971, 979, 957, 964, 982, 962, 989, 965, 978, 951, 948, 966, 967, 958, 955, 960, 959, 963, 984, 981, 952, 947, 983, 949, 988, 950, 968, 986, 985, 961, 987, 974, 975, 976, 977, 946, 980, 969, 972, 973]. InterOnto [910]. Interoperation [665]. Interplay [838]. Interpolation [621, 477]. Interpretation [644]. Interpreters [122]. Intersection [836]. Interval [655, 283, 609]. Interventions [971]. Intra [556]. Intra-operative [556]. Intrinsic [589]. Introducing [227]. Introduction [558, 379]. Intrusion [48, 44]. intrusive [70]. Invalidations [511]. Invariant [867, 592]. Invariants [879]. Invasive [72, 748]. Inventory [274, 650]. Invisible [51]. IPCAI [971]. IPTV [588]. Iris [460]. Irregular [605]. Issues [880, 640]. Istanbul [987]. Italian [631]. Italy [971, 962, 955, 960, 959, 963, 952, 635, 626]. Items [55]. Itemset [712]. Iterative [585, 595, 550]. ITIB [980]. IV [977]. IWOMP [955].

J2ME [420]. J2ME-Enabled [420]. Java [259, 420, 402, 246, 124]. JavaScript [260, 261, 262]. Jetpack [258]. Job [656]. Join [229, 719, 725]. Joint [785, 829]. July [990]. June [971, 979, 957, 964, 982, 962, 989, 965, 978, 951, 948, 966, 967, 970, 958, 955, 960, 959, 963, 984, 981, 983, 949, 988, 968, 986, 985, 961, 954, 974, 975, 976, 977, 956, 946, 980, 969, 944, 945, 972, 973]. Jürgen [943].

K-Nearest [478]. Kan [809]. Keeping [872]. Kernel [760, 609]. Kernel-Based [760]. Kernelized [736]. KES [968]. KES-AMSTA [968]. Key [788, 773, 154, 318, 776, 402]. Keys [788]. Keystroke [848]. Keywords [595]. Kgs [990]. Kindergarten [351]. Knapsack [652]. Knowledge [913, 217, 490, 820, 596, 829, 131, 944, 945, 599, 626]. Known [370]. Kuala [944, 945]. Kummer [790]. Kummer-Based [790]. Kynoid [421].

Label [35, 707, 168]. Labeled [105]. Labeling [693]. Labial [766]. Lack [830]. Landform [607]. Landmark [46]. Language [706, 978, 372, 119, 413, 694, 704, 928, 9, 506, 38, 248, 702, 135, 95, 155, 142, 819, 89, 936]. Languages [126, 141, 82, 140, 932, 10, 148, 510, 941, 943]. Laparoscopy [545, 544]. Laplacian [760]. Large [50, 428, 352, 727, 849, 257, 721, 241, 915, 724, 725]. Latent [186]. Lattice [417, 599]. Lattice-Based [417]. Launch [663]. Launching [851]. Law [436]. Laws [795]. Layer [575, 732, 869]. Lazy [207]. LBlock [419]. Leader [598]. Leaks [849]. Lean [823]. Learned [481, 826]. Learner [297, 302, 290, 288, 293]. Learner-Support [288]. Learning [187, 598, 189, 332, 207, 190, 698, 217, 490, 564, 27, 35, 36, 882, 292, 296, 441, 488, 489, 23, 38, 299, 208, 301, 699, 28, 41, 287, 763, 298, 584, 304, 26, 213, 202, 587, 293, 96, 408, 858, 219]. Least [200]. Lectures [962, 954]. Left [733]. Leftmost [11]. Legacy [120, 930, 673]. Length [779]. Length-Doubling [779]. Lengthening [10]. Lens [535]. Lesions [478]. Less [370]. Lesson [112]. Lessons [481, 826]. Level [404, 164, 687, 462, 448, 451, 68, 937, 143]. Levels [401]. Leveraging [351, 51]. LGIP [465]. libOMP [229]. Libraries [257]. Library [156, 103, 303]. LiDAR [529]. Life [356, 989, 49, 240]. Lifetime [666]. Lifetime-Maximizing [666]. Lightweight [496, 940, 932, 840, 255]. Like [298]. Limb [740]. Limited [855]. Limits [283]. Line [901, 387, 497]. Linear [900, 787, 333, 108, 39, 45, 523]. Linearized [452]. Lines [127, 104]. Link [59]. Linkage [36, 51]. Links [432, 910]. Liquor [601]. List [541]. Literal [335]. Litmus [694]. Live [545, 358]. Liver [735]. Load [579]. Loans [581]. Local [188, 330, 630, 803, 470, 550, 623, 584, 448, 465]. Localization [854, 546, 633]. Localized [439]. Locally [614, 457]. Location [643, 366, 713, 740]. Locations [578]. Locators [124]. Lock [257]. Lockdown [835]. Log [517]. Logic [194, 846, 770, 107, 108, 317]. Logical [80, 816, 511, 699]. Logics [794, 324]. Logs [519, 912, 518]. Long [820, 547]. Long-Term [820, 547]. Lookup [86]. Loop [765, 228]. Loops [237]. Lossy [784]. Love [432]. Low [404, 589, 449, 448]. Low-Dimensional [589]. Low-Level [404, 448]. Lower [655, 740]. LTL [901]. Lumping [278]. Lumpur [944, 945]. Lyngby [990]. Lynx [339].

MAC [869]. Machine [496, 598, 929, 190, 882, 292, 199, 216, 142].

Machines [309, 959, 147, 579, 89, 933, 935]. Macro [187]. Madrid [984, 983]. Maintainability [670, 677]. Maintaining [680]. Maintenance [571, 679]. Make [330]. Making [630, 505, 497]. Malaysia [944, 945]. Malicious [687, 691]. Malware [844, 687]. Mammographic [738]. Man [775]. Man-in-the-Middle [775]. Management [979, 384, 490, 513, 163, 72, 820, 726, 828, 521, 596, 829, 632, 318]. Managing [932]. Manifolds [169, 589]. Manipulation [533]. Manufacturing [626]. Many [231]. Map [539, 264]. Mapping [699, 93, 749]. Mappings [912]. MapReduce [725]. Maps [170, 530, 357]. Margin [27]. Margin-Based [27]. Marketing [627]. Markov [276, 890, 435, 277, 560]. MARS [494]. MARTE [668]. MAS [161]. Mass [641]. Masses [244]. Massive [33]. Master [638]. Matching [181, 520, 556, 271, 61, 678, 458, 145]. Materials [341]. Math [298]. Mathematical [660, 987]. Mathematics [983]. MATLAB [249, 683, 866]. Matrix [703, 527]. Matter [20, 75, 114, 128, 149, 166, 184, 220, 284, 381, 422, 437, 561, 810, 852, 886, 920, 21, 42, 522, 543, 562, 583, 604, 625, 646, 667, 688, 709, 730, 751, 772, 793, 811, 832, 853, 874, 887, 908, 921, 942, 63, 76, 97, 115, 129, 150, 167, 185, 206, 221, 242, 263, 285, 306, 327, 348, 369, 382, 403, 423, 438, 459, 480, 501, 734]. Matting [752]. Max [281, 447]. Maximal [876, 898]. Maximization [657]. Maximizing [666]. MaxSAT [193, 195]. May [948, 952, 947, 949, 950, 953, 946, 944, 945]. MCPR [970]. McSAF [249]. MD [989]. MD5 [409]. MDE [371, 374, 932]. Me [842]. Mean [176, 277]. Mean-Field [277]. Means [736, 825]. Measure [675]. Measurements [768, 755, 744, 529]. Measures [680]. Measuring [79, 84, 290, 498, 205]. Mechanical [740]. Mechanism [123, 89, 198]. Mechanisms [837, 69]. Mechanized [394]. Medial [621]. Medical [475, 744, 476]. MedIntegraWeb [665]. MEDLINE [916]. Meet [245]. Meeting [831, 923]. Mehrotra [648]. Mehrotra-Type [648]. Melanomas [737]. Membrane [14]. Memetic [573]. Memories [525]. Memory [225, 67]. Mental [294]. Merge [275]. Mesh [760, 179]. Message [91]. Messages [516]. Meta [207, 88, 941]. Meta-Learning [207]. Meta-Modelling [941]. Meta-programs [88]. Metacognitive [293]. Metadata [704, 708]. Metamodel [675]. MetaTutor [290]. Method [760, 736, 608, 901, 452, 492, 733, 651, 462, 529, 86, 584]. Methodology [766, 162]. Methods [962, 78, 963, 645, 759, 890, 877, 94, 647, 22, 204, 547]. Metrics [685, 670, 831, 686, 683, 684, 819, 677]. Mexican [970]. Mexico [970]. MIC [231]. Micro [720, 497]. Micro-Blogging [497, 720]. Microgrids [634]. Microscopy [881]. Middle [775]. Middleware [153, 662]. Migrating [930]. Migration [496]. Mimic [350]. Min [447]. Minimal [899]. Minimum [593]. Mining [712, 44, 516, 191, 355, 728, 885, 47, 46, 696, 40, 703, 862, 517, 52, 600, 944, 945, 62, 175, 892]. Mirror [15]. Mirrors [243]. Mislabeling [31]. Mitigating [775]. Mitral [552]. Mixed [24, 34]. Mixture [736, 450, 453]. MLICC [707]. Mobile [269, 664, 411, 79, 355, 420, 367, 665, 484, 198, 52, 494, 663, 767, 842].

Mobility [354]. modal [550, 864]. Modality [360]. Mode [629]. Model [134, 100, 549, 143, 377, 939, 77, 962, 207, 379, 426, 906, 141, 136, 736, 373, 927, 608, 901, 158, 138, 479, 614, 140, 312, 430, 132, 653, 838, 814, 650, 769, 950, 553, 264, 147, 585, 104, 157, 924, 847, 414, 540, 699, 99, 934, 931, 938, 392, 268, 930, 131, 581, 216, 476, 317, 380, 133, 144, 925, 148, 858, 219, 390, 941, 817, 139, 756]. Model-Based [549, 939, 104, 216, 925, 817]. Model-Checking [100]. Model-Driven [962, 136, 157]. Model-Sensitive [144]. Model-to-Model [148]. Modeling [957, 116, 378, 146, 897, 313, 270, 820, 393, 444, 49, 376, 366, 213, 679, 668, 216, 353, 935]. Modelling [486, 352, 159, 838, 280, 218, 867, 491, 215, 529, 399, 689, 590, 510, 941, 990]. Models [937, 965, 110, 424, 716, 80, 504, 516, 947, 374, 888, 401, 926, 506, 514, 923, 519, 890, 644, 450, 505, 551, 930, 131, 277, 560, 144, 817]. Modern [226]. Modes [6]. Modified [873]. Modular [807, 653]. Modularisation [124]. Modulation [448]. Modules [183]. Moment [879]. Moments [592]. Monitoring [295, 349, 711, 636, 68, 864]. Monocular [545]. Monotone [276]. Monte [527]. MoProSoft [815]. MOPSO [571]. Morocco [981]. Morphing [202]. Morphological [878]. Morphology [987]. Morse [169]. Motif [866]. Motifs [60]. Motion [549, 848, 553, 544]. Motivated [448]. Motivation [302, 301]. Motivational [299]. Mozilla [258]. MPC [983]. MPI [240]. MPI/OpenMP [240]. MR [734]. MRgHIFU [549]. MRI [556, 733, 557]. MSEPT [946]. MST [610]. MULE [393]. MULE-Based [393]. Multi [496, 937, 143, 416, 607, 181, 490, 434, 716, 520, 24, 756, 35, 513, 550, 773, 308, 32, 707, 968, 753, 89, 467, 453, 240, 936, 609, 922, 213, 679, 695, 494, 690, 872, 460, 483, 925, 68, 574, 39, 864]. Multi-Agent [490, 968, 494, 496, 483]. Multi-camera [467]. Multi-class [32, 39]. Multi-commodity [434]. Multi-core [716, 240, 922, 68]. Multi-dimensional [690]. Multi-document [695]. Multi-Domain [925, 922]. Multi-equation [716]. Multi-factor [773]. Multi-GPU [607]. Multi-Label [707, 35]. Multi-language [89, 936]. Multi-level [937, 143]. Multi-modal [550, 864]. Multi-model [756]. Multi-objective [213]. Multi-optimization [574]. Multi-Party [416]. Multi-player [24]. Multi-regression [679]. Multi-scale [181]. Multi-stage [460]. Multi-Start [872]. Multi-step [308]. Multi-strategy [520, 453]. Multi-tenant [513]. Multi-threaded [609]. Multi/Hyperspectral [753]. Multiactor [637]. Multiagent [208, 305]. Multibiometric [462]. Multicast [318]. Multichannel [746]. Multicomponent [756]. Multicore [607, 65, 66, 74, 73, 226, 946]. Multidimensional [652, 630, 636]. Multifocal [135]. Multilayer [408]. Multilingual [696, 702]. Multimodal [296]. Multimoded [152]. Multiparty [91]. Multiplayer [397]. Multiple [719, 386, 36, 653, 60, 478, 537, 323, 250, 202, 557]. Multiplication [527]. Multipole [584]. Multiprocessor [152]. Multiscale [800, 738]. Multisource [758]. Multispectral [755, 458, 757]. Multistart [654]. Multivalued [177]. Music [205]. Mutation [916, 81, 570, 576]. Mutual [432, 303, 935]. My [645, 842].

’n [254]. Named [705, 29]. Named-Entity [705]. Naples [638]. National [917]. Natural [706, 978, 694, 704, 506, 358, 95, 661, 978]. Navigating [710]. Navigation [368, 360, 359, 552, 366]. NCBI [909]. Near [460]. Nearest [442, 478]. Negative [518, 859, 703]. Negotiation [496]. Neighbor [442, 368]. Neighborhood [30]. Neighbors [859]. Neighbour [478]. Nested [237, 334]. Net [888, 12]. Netherlands [978]. Nets [309, 895, 906, 897, 903, 988, 907, 893, 890, 894, 904, 902, 891, 868, 896]. Nets-within-Nets [906]. Nets-within-Nets-Formalism [907]. Network [281, 269, 982, 384, 119, 78, 53, 55, 355, 687, 869, 60, 603, 197, 915, 268, 485, 679, 613, 408, 410]. Network-Based [119]. Network-on-Chips [613]. Networked [503, 502]. Networks [50, 706, 273, 59, 431, 430, 715, 218, 393, 662, 432, 198, 711, 508, 611, 525, 394, 45, 590]. Neural [218, 611, 679, 525, 408, 219]. Neuro [629]. Neuro-Fuzzy [629]. Neuroevolution [208]. Neurogenetic [218]. Neurosurgery [556]. Newcastle [961]. News [196, 49]. Next [56]. NFC [845]. NFC-Enabled [845]. Nigra [557]. NLDB [978]. NN [610, 610]. NN-Boruvka [610]. Node [350]. Nodes [428, 53, 241]. Noise [31, 451, 477]. Noiselet [744]. Noisy [531]. Non [676, 512, 377, 297, 550, 859, 703, 92, 748, 507, 70, 171, 933]. Non-conscious [297]. Non-functional [676, 377, 507, 933]. Non-interference [92]. Non-intrusive [70]. Non-invasive [748]. Non-iterative [550]. Non-negative [859, 703]. Non-relational [512]. Non-uniform [171]. Nonlinear [769]. Nonterminal [18]. Normative [482]. Note [6]. Notification [69]. Noun [495]. Novel [25, 598, 728, 603, 554, 590]. NS [965]. NTP [815]. Nuclear [201]. Nuclear-Test-Ban [201]. Nuclei [739]. Nucleus [557]. NUMA [241, 235]. Number [443, 575, 444, 283]. Numerical [641, 477].

O [222, 69]. Object [372, 670, 81, 83, 888, 251, 686, 956, 93, 445, 252, 523]. Object-Oriented [670, 81, 83, 686, 956]. Object-Relational [93]. objective [213]. Objectives [271]. Objects [531, 880, 692, 755, 855, 245, 947]. Obligations [323]. Observation [197]. Observer [485]. Occasion [943]. occurrence [857]. Occurring [290, 47]. Oceans [710]. OCL [110, 940, 372]. OFDM [767]. Off [188, 331]. Oil [858]. Oils [659]. OLAP [44]. Old [809, 899]. OMC [44]. OMC-IDS [44]. OMP2012 [238]. On-Line [497]. On-the-Fly [937, 901]. ONCO [918]. ONCO-i2b2 [918]. One [11, 201, 19]. One-Class [201]. One-Sided [11]. One-Way [19]. Online [2, 28]. online.com [112]. Onshore [826]. Ontologies [499, 915]. Ontology [490, 749, 910, 602]. Ontology-Based [602]. OPC [928]. Open [822, 500]. Open-Source [500]. OpenMP [955, 237, 225, 229, 955, 223, 241, 222, 238, 224, 234, 236, 232, 227, 235, 228, 239]. OpenMP* [226]. Operating [664, 844, 70]. Operation [576]. Operational [308, 904, 865]. Operations [170]. operative [556]. Operator [449, 924]. Operators [178, 81, 117, 763]. Optimal [415, 763, 440]. Optimality [418]. Optimisation [809, 91]. Optimistic [605]. Optimization [616, 566, 577, 582, 654, 739, 624, 653, 658, 575, 570, 580, 651, 647, 462, 213, 581, 576, 572, 578, 569, 342, 137, 574]. Optimizations [234]. Optimize [661]. Optimizer [873, 574, 573]. Optimizing [232, 597]. Option [264]. OR-Proof [781]. OR-Use [558]. Order [791, 440, 476, 591, 780]. Organ [915]. Organization [615]. organizational [516]. Organizations [487]. Orientation [745]. Orientations [556]. Oriented [670, 120, 81, 816, 83, 686, 596, 104, 492, 414, 956, 161, 399, 935]. OSGi [85]. Other [986]. OtO [520]. OTP [775]. Outcomes [293, 851]. Outlier [57]. Outliers [856]. Outputs [202]. Outsourcing [513, 826]. Overflows [314]. Overhead [68]. Overlapping [7, 45]. Overlays [643]. Oversubscription [66]. Overview [620]. Ownership [251, 250].

P2P [643]. PaCE [119]. Pacemaker [311]. Pacific [944, 945]. PageRank [427, 434, 433]. Pages [707]. Pairwise [188]. Paisley [145]. PAKDD [944, 945]. Panel [726]. Paper [121]. Papers [987]. Paradigm [728]. Paradigms [229]. Parallel [619, 343, 607, 616, 719, 608, 344, 73, 222, 238, 91, 172, 246]. Parallelization [716, 71, 605]. Parameter [187]. Parameters [415, 580, 763]. Parametric [283, 440]. Parikh [17]. Park [989]. Parkinson [885]. Parsing [12]. Part [966, 967, 974, 975, 976, 977, 944, 945, 972, 973, 469]. Partial [550, 7, 402, 324, 315, 131]. Partially [400]. Partially-Supervised [400]. Particle [616, 577, 582, 575, 570, 580, 873, 462, 581, 576, 572, 578, 574, 569, 573]. Particles [575]. Partitioning [425, 716, 365]. Parts [523]. Party [416, 415]. Passage [586]. Passages [892]. Password [844]. Past [802]. Path [721, 492, 801, 333]. Path-Oriented [492]. PathCrawler [113, 112]. PathCrawler-online.com [112]. Paths [821]. Pathway [209]. Patients [748, 918]. Pattern [712, 218, 303, 743, 647, 56, 534, 663, 525, 440, 145, 465, 970]. Pattern-Based [663]. Patterns [911, 48, 947, 47, 623, 457, 291, 54, 69, 394]. Pause [254]. Pay [912]. Pay-as-You-Go [912]. PCA [452]. PDE [759]. PE [402]. PE-KeY [402]. Pedagogical [290, 288]. Pedestrian [366]. Pedestrians [367]. Peer [643, 287]. -160 [789]. AIE [986]. AVC [765]. CA [270]. Classification [758]. Guarantee [386]. Hyperspectral [753]. Invited [954]. M [264]. N [264]. OCL [926]. OpenMP [240]. Penetrating [724]. People [368, 358, 693]. People-Centric [368]. Peptide [1]. Perception [297]. Perceptions [850]. Perceptrons [705]. Perceptual [524]. Percolated [58]. Perfect [169, 273, 869]. Performability [269]. Performance [778, 616, 880, 378, 608, 163, 271, 222, 946, 240, 376, 458, 609, 67, 236, 69, 872, 767]. Periodic [871]. Periodically [268]. Perishable [650]. Permissions [90]. Permutation [443]. Persistence [183, 171]. Persistent [182]. Personal [189, 55]. Personalized [349]. Perspective [691, 483]. Perspectives [639]. Perturbation [542]. Pervasive [961, 361, 961]. Pescara [635]. Petri [988, 897, 903, 890, 904, 12, 902, 868, 896]. pGCL [396].

PH [282]. PH-Distributed [282]. Phase [882, 550, 831, 8, 764]. Phase-Based [550]. Philosophy [217]. Phones [367, 494]. Photometric [528]. Physical [839, 863]. PICARO [791]. Picture [4]. Pinpointing [687]. Pisa [971, 959, 963]. Pixels [753]. PKIs [776]. Placement [54]. Plagiarism [678]. Plan [638]. Planarity [883]. Planner [934]. Planning [187, 660, 903, 195, 647, 366]. Plans [321, 144]. Plant [400]. Plants [400]. Platform [512, 405, 552, 599, 842, 362, 497]. Platforms [230, 503, 835, 500]. Plausibility [406]. Play [254]. player [24]. Plenary [954]. Plenary/ Invited [954]. Plus [281]. Point [866, 171, 790]. Points [674, 177, 250, 174, 262, 875, 718]. Points-To [262, 674]. Poisson [279]. Poland [980, 969]. Polarity [692]. Policies [78, 240, 421]. Policy [413, 267]. Polish [702]. Political [690]. Polychronous [391]. Polymorphic [255]. Pool [332]. Poor [190]. Poporo [78]. Population [564]. Population-Based [564]. Portfolio [581, 345]. Portfolio-Based [345]. Porting [765]. Portugal [966, 967]. Position [121, 367]. Positioning [368]. Positions [711]. Possibility [748]. Post [802]. Posture [473, 867]. Potential [849]. Power [2, 436, 418]. PQL [246]. Practical [835]. Practicality [848]. Practice [964, 950, 792]. Pragmatics [689]. Prague [948, 947, 949, 950, 946]. Pre [556]. Pre-operative [556]. Precise [88]. Predicting [670, 304, 204, 353]. Prediction [59, 352, 861, 283, 539]. Predictive [217]. Predictor [648]. Predictor-Corrector [648]. Preemption [164]. Preemption-Level [164]. Preference [195, 692]. Preference-Based [195]. Preferences [194, 806]. Preferred [426]. Prefetching [230, 663]. Preliminary [749]. Preprocessing [882, 347, 335]. Prerequisite [67]. Prescribed [283]. Presence [257]. Preservation [412]. Preserving [398, 762, 591, 410]. Preventing [314]. Priced [894]. Priming [297]. Primitives [833]. Prior [302]. Privacy [964, 78, 411, 849, 412, 851, 410]. Private [836]. ProB [385]. Probabilistic [794, 712, 719, 715, 860, 393, 711, 213, 399]. Probabilities [364]. Probability [330, 530, 798, 197]. Problem [425, 652, 65, 339, 32, 563, 9, 570, 893, 540]. Problems [566, 582, 871, 203, 892]. Procedure [555]. Procedures [554, 620]. Proceedings [966, 967, 974, 975, 976, 977, 944, 945, 972, 973, 971, 979, 957, 964, 982, 989, 965, 978, 951, 948, 970, 958, 955, 960, 959, 963, 984, 981, 952, 947, 983, 949, 988, 950, 968, 986, 985, 961, 953, 956, 946, 980, 969, 990]. Process [127, 384, 818, 630, 984, 504, 927, 516, 816, 814, 795, 506, 514, 517, 519, 815, 813, 821, 518, 892]. Process-Oriented [816]. Processes [281, 278, 79, 812, 828, 491, 505, 293, 898]. Processing [971, 706, 978, 981, 884, 527, 702, 745, 761, 95, 878, 69, 725]. Processor [605]. Processors [57, 64, 68]. Product [818, 820, 104, 784, 984]. Product-Focused [984]. Production [766]. Products [7]. PROFES [984]. Profile [488, 668, 851]. Profiles [896]. Profit [657]. Prog&Play [301]. Program [983, 809, 683, 819]. Programmatic [339]. Programming [805, 120, 660, 650, 795, 329, 91, 956, 246, 392, 107, 155, 597]. Programs [81, 386, 674, 158, 637, 194, 402, 395, 124, 605, 88]. Project [82, 820, 829].

Projected [743]. Projection [203]. Projections [550]. Projector [622]. Projectors [361]. Projects [830]. Promise [245]. Promote [639]. Proof [781, 395, 323]. Proofs [794, 948, 782]. Propagation [204]. Properties [389, 426, 431, 85, 900, 847, 842]. Property [635, 340]. Proportional [585]. Proposal [616, 631, 817]. Proprioceptor [485]. Prospect [636]. Prosthesis [750]. Protection [844]. Protein [589, 539, 590]. Protein-Protein [590]. Protocol [773, 407, 318]. Protocols [100, 415, 846]. Provenance [847]. Provenance-Based [847]. Provers [323]. Providing [827]. Proving [805]. Proximity [79]. Proxy [843]. PSO [579]. PSO-Based [579]. Psychology [98]. Public [514]. Publication [410]. PUF [841]. PUF-Based [841]. Purely [246]. Purely-Declarative [246]. Purging [517]. Purpose [70]. Pushdown [13].

QBF [337]. QBFs [336]. QSAT [111]. Quality [681, 190, 680, 81, 831, 732, 466, 453, 601, 586]. Quality-Based [466]. Quantification [628, 94]. Quantified [335]. Quantifiers [338, 326]. Quantifying [50]. Quantitative [393, 493]. Quantum [566, 570, 580, 19]. Quantum-Behaved [570, 580]. Quantum-inspired [566]. Quarry [770]. Queries [77, 701, 714, 919, 704, 928, 715, 721, 697, 931, 693, 713]. Query [912, 52]. Querying [914]. Quest [214]. Question [913, 870, 600]. Questions [698]. Queue [272, 267, 615]. Queueing [266, 274, 264]. Queues [273, 265]. Quick [428, 663]. Quo [212].

R [248]. Radial [879, 647]. Radio [767]. Radiotherapy [647]. Railways [387]. Random [111, 4, 431, 282, 33, 722, 11, 435, 34, 30]. Randomness [785, 2]. Range [714, 473, 711, 539]. Ranges [715]. Rank [724]. Ranking [429, 912, 910, 591]. Ranks [856]. Rapport [287]. Rare [48]. Rat [219]. Rate [564, 444, 408]. Rate-Distortion [444]. Rational [628]. Ravenscar [151]. Ray [551]. RDF [725]. Re [331]. Re-examining [331]. Reachability [340]. Reaction [3]. reactive [386]. Reader [364]. Reading [586]. Real [295, 800, 83, 153, 364, 861, 165, 761, 199, 421, 547]. Real-Time [295, 800, 83, 153, 364, 861, 165, 152, 761, 421, 547]. Real-World [199]. Reasoning [940, 297, 803, 386, 160, 865]. reCAPTCHAs [538]. Reciprocal [59]. Reciprocity [50]. Recognition [463, 471, 705, 966, 967, 970, 538, 470, 472, 860, 218, 473, 29, 464, 867, 741, 533, 534, 536, 465]. Recognizing [474, 363]. Recommendation [37, 494, 588]. Recommendations [186]. Reconfiguration [153]. Reconstructibility [895]. Reconstruction [535, 622, 640, 182, 551, 875, 171]. Record [36, 872]. Record-Keeping [872]. Recovery [528]. Recursive [106]. Red [557]. Redesign [147]. Reduce [101]. Reduced [419]. Reduced-Round [419]. Reduction [619, 74, 753]. Reengineering [673]. Refinement [121, 388, 398, 321, 322, 337, 530, 938, 618]. Refinement-Based [388]. Refinement-Preserving [398]. Refining [620]. Regard [859]. Region [883, 445, 875]. Regions [463, 624]. Register [398]. Register-Voice [398].

Registration [556, 550]. Regression [631, 614, 934, 200, 679]. Regular [332, 721, 10]. Regularization [585]. Regulation [482, 304]. Rehabilitation [475]. Reiner [658]. Reinforced [745]. Reinforcement [763]. Relation [396, 674, 367]. Relational [794, 214, 93, 22, 896, 512]. Relations [916, 926, 938, 105]. Relationship [826]. Relationships [293]. Relatives [726]. Relaxing [118]. Relevance [3]. Reliable [951, 165, 837, 951]. Reliant [356]. Rely [386]. Rely/Guarantee [386]. Remanufacturing [568]. Remarks [862]. Remote [864]. Removal [170, 700]. Reordering [778]. Repair [552]. Repeating [457]. Replication [271]. RepOK [101]. Reporting [88]. Reports [389, 96]. Representation [211, 607, 48, 465]. Representations [296, 448, 878, 523]. Representative [718]. Representing [926, 93]. Republic [948, 947, 949, 950, 946]. Reputation [406, 198]. Reputation-Based [198]. Requirement [672, 831, 154]. Requirement-Based [154]. Requirements [676, 694, 812, 400, 157, 669, 933]. Research [619, 917, 885, 726, 986, 829, 749, 748, 581, 592, 594]. Resequencing [717]. Residual [659]. Resilient [161]. Resistance [791]. Resolution [754, 332, 336, 333]. Resolution-Path [333]. Resolve [934]. Resource [164, 72, 894]. Resource-Constrained [894]. Resources [66]. Respiratory [549, 553]. Responses [290]. Restoration [525]. Restrictions [118]. Retractions [177]. Retrial [265]. Retrieval [181, 700, 697, 918, 592]. Reusability [819]. Reusable [134, 884, 122]. Reuse [785, 148]. Reutilization [510]. Reveal [692]. Reverse [803, 209]. Reverse-Engineering [209]. Reversed [278]. Review [686, 618]. Reviews [691]. Revised [987]. Revising [123]. Revisited [658, 541, 390]. Revisiting [343]. Revocable [776]. Revocation [834]. Reward [277]. RFID [364, 407]. Rheocardiography [748]. Rigorous [307, 383]. RIKE [776]. Rings [786]. RIPEMD [789]. RIPEMD-128 [789]. RIPEMD-128/-160 [789]. RNA [339]. RNA-Folding [339]. ROAC [414]. Road [668]. Roads [44]. Robot [533]. Robots [615]. Robust [556, 452, 477]. Rodin [320]. Role [330, 414]. Role-Oriented [414]. Rome [955]. ROS [905]. Round [419]. Rounds [789]. Route [360, 366]. Routes [355]. Routing [570, 613]. RPMS [665]. RR [8]. Rudeness [287]. Rule [758, 46, 679, 148]. Rule-Based [148]. Rules [191, 673]. Run [87]. Run-Time [87]. Runtime [229, 384, 230, 374, 228, 68]. Runtime-Prefetching [230].

SaaS [513]. Sabotaged [826]. Safe [140, 645, 91, 835]. Safety [387, 863, 902]. Salient [753]. Salvador [974, 975, 976, 977]. Samples [443]. Sampling [273, 528, 722, 527, 213, 875]. Sandbox [130]. SAT [960, 343, 111, 339, 871, 344, 926, 346, 347]. SAT-Based [926]. Satellite [754]. Satisfaction [204]. Satisfiability [960, 334, 329]. Scala [140]. Scalable [610, 355, 344, 33, 61, 62]. Scale [352, 849, 181]. Scaling [526]. Scenario [109, 897, 488]. Scenario-Based [109, 897]. Scenes [547]. Scense [497]. Scheduler [797]. Scheduler-Independent [797]. Scheduling [871, 799, 228, 606]. Schema [912]. Schemas [135]. Scheme [230, 666, 774]. Schemes [785, 24]. School [962]. Science [217, 974, 975, 976, 977]. Sciences [989]. Scientific [500, 979]. SCIMS [521]. Sclerosis [478]. SCOPE [868]. scoping [224]. Scoring [586]. Scrum [821]. SDSS [638]. Seals [850]. Search [612, 330, 65, 727, 903, 909, 855, 52, 647, 26, 693, 872, 144]. Searching [190]. Second [476]. Secret [417]. Sector [516, 514, 626]. Secure [416, 786, 591]. Security [404, 100, 785, 964, 982, 794, 413, 781, 773, 846, 421, 835, 842]. Seed [94]. Segmentation [735, 538, 479, 456, 731, 466, 733, 445, 529, 734, 460, 557]. Segmenting [478]. Segments [876]. Segway [359]. Seismic [770]. Selected [987]. Selecting [763]. Selection [127, 643, 653, 87, 585, 918, 597, 439]. Selectors [345]. Self [454, 192, 364, 304]. Self-calibration [364]. Self-regulation [304]. Self-similarity [454]. Self-Training [192]. Semantic [186, 919, 324, 511, 596, 517, 870, 289]. Semantics [396, 308, 689, 142, 896]. Semi [441, 699]. Semi-supervised [441]. Semigroup [7]. Sensed [744]. Sensing [311, 354, 864]. Sensitive [674, 192, 144]. Sensor [715, 393, 662, 350]. Sensors [750, 350]. Sentences [699]. Sentiments [189]. Separable [655]. Separation [332, 769, 157]. Sequence [929, 436, 717]. Sequences [276]. Sequent [336]. Sequential [309, 56]. SER [612]. SER-Based [612]. Series [25, 61]. Serious [303, 301]. Server [416]. Service [272, 670, 643, 513, 828, 596, 509, 508, 124, 399, 487]. Service-Oriented [399]. Services [356, 503]. Session [91]. Set [326, 836, 31, 462, 37]. Sets [727, 900, 534, 899, 171]. Setting [199]. Settings [925]. Setup [786]. Several [862]. SFM [962]. Shading [545, 544]. Shadow [441]. Shape [183, 181, 456, 528, 876, 755, 544]. Shape-from-Motion-and-Shading [544]. Shaped [350]. Shared [630]. Sharing [411, 417, 118, 837]. Sheets [685]. Shenzhen [972, 973]. Shop [656]. Short [701, 733, 539]. Short-Axis [733]. Short-Range [539]. Shortening [10]. Shortest [801]. Showcase [214]. Shrinkage [476]. SI [565]. SI-Based [565]. Side [791]. Side-Channel [791]. Sided [11]. SIFT [458]. Sign [533, 536]. Signal [981, 770]. Signals [295, 746, 668]. Signature [785, 473]. Signatures [783]. SIMD [226]. Similarity [727, 447, 84, 859, 61, 205, 454]. Simple [177, 769, 900]. Simplex [658]. Simplified [817]. Simulated [539]. Simulation [389, 629, 138, 559, 232]. Simulations [481, 305]. Simultaneous [757]. Singapore [982]. Single [267]. Singularities [351]. Situation [868]. Situations [486]. Size [833]. Size-Constrained [833]. Sizing [799]. Skeleton [475]. Skeletonizing [172]. Skill [560]. Skills [559]. Skin [463, 530]. Skyline [718]. SkyQuery [719]. Slicing [902]. Small [657, 308, 815]. Small-Step [308]. Smaller [259]. Smalltalk [86]. Smart [364, 357, 358]. Smartphones [845]. SmartTokens [845]. Smells [682]. Smoothing [760]. SMS [700]. SMS-Based [700]. SMT [320, 341, 891]. SMT-Aided [341]. SMT-Based [891]. Snapshotable [90]. SNP [717]. SOA [673]. Social [59, 78, 53, 628, 55, 430, 521, 432, 603, 45, 410]. Socially [521]. Socially-Aware [521]. Society [502]. Socio [826]. Socio-technical [826]. Soft [153]. Software [404, 681, 939, 962, 65, 951, 818, 79, 378, 515, 984, 82, 653, 830, 831, 838, 125, 104, 924, 815, 678, 946, 99, 240, 376, 813, 245, 392, 823, 679, 825, 868, 677, 949]. Solutions [279, 640]. Solved [452]. Solver [339, 345]. Solvers [320, 344]. Solving [343, 652, 657, 871, 650, 337, 346, 540, 342]. SOM [446]. Some [426, 276]. SOR [608]. SOS [512]. Soundness [893, 894]. Sounds [860]. Source [769, 822, 247, 154, 500, 819]. Sources [913]. Southern [626]. Space [589, 906, 867]. Spaces [361, 365]. Spain [984, 983]. Spanning [179]. Sparse [560]. Sparsifying [429]. Spatial [630, 768, 426, 645, 548, 644, 532, 878]. Spatially [478]. Spatio [218, 56, 642]. Spatio- [218]. Spatio-Explorative [642]. Spatio-sequential [56]. SPEC [238]. Specific [126, 928, 142, 594, 941, 137]. Specification [111, 132, 222, 380]. Specification-Driven [132]. Specifications [800, 103, 890]. Specified [582]. Specifying [938]. Spectral [594]. Spectro [218]. Spectro-Temporal [218]. Specularity [545]. Speech [766]. Sphere [447]. Spiking [218]. Spills [858]. Split [746, 275]. Split-and-Collect [746]. Split-Merge [275]. Spread [53]. Spreadsheet [682]. Spreadsheets [136]. SQL [704]. SRF [31]. SSDBM [979]. SSIM [454]. Stability [563]. Stable [432]. Stage [705, 460]. Standard [81]. Start [872]. State [309, 77, 929, 959, 4, 276, 147, 92, 15, 933, 935]. State-Charts [92]. State-of-the-Art [77]. Stateless [8]. Statements [804]. States [293, 19]. Static [249, 613, 390]. Statistical [979, 449]. Statistics [670, 645, 440]. Statistics-Based [670]. Steady [276]. Steady-State [276]. Steel [745]. Step [102, 308]. Steps [126]. Stepwise [631]. Stereo [757]. Stereotypes [672]. Stimulation [557]. Stochastic [957, 330, 650, 280, 60]. Stockholm [951]. Stop [435]. Stopword [700]. Storage [591]. Strangers [726]. Strategic [198, 135, 632]. Strategies [820, 286, 304, 627]. Strategy [188, 564, 489, 571, 537, 551, 539, 520, 453]. Strategy-Based [489]. Streaming [61, 711]. Streams [43]. Strengthening [926]. String [940, 727, 678]. Strings [714]. Stroke [741]. Strong [334]. Structural [447, 454, 112, 113, 783, 565]. Structure [35]. Structured [913, 253]. Structuredness [504]. Structures [783, 107, 105]. Stubborn [900]. Student [300, 301, 298, 304, 291]. Student-Adaptive [300]. Studies [933]. Study [443, 672, 471, 384, 294, 556, 927, 516, 637, 193, 575, 351, 828, 850, 358, 31, 247, 632, 930, 826, 203, 925]. Style [685, 798]. Subgraphs [722, 723]. Subjects [734]. Sublinear [427]. Subsets [174]. Subspace [180]. Substantia [557]. Substructure [728]. Subthalamic [557]. Subtle [738]. Subword [16]. Suite [84, 684, 238]. Suites [101]. Sum [715]. Summaries [695]. Summarization [196, 51]. Superpixels [524]. Supervectors [623]. Supervised [400, 699, 441]. Supervision [486, 699]. Supervisory [313]. Supplier [597]. Supply [634, 649]. Support [672, 84, 288, 924, 904, 627, 865, 933, 776]. Supported [368, 568]. Supporting [88]. SURF [458]. Surface [528, 551]. Surfaces [176, 877, 743]. Surgical [559, 560]. Surrogate [614]. Sustainable [634]. Sustaining [824].

SVM [754, 623]. Swarm [652, 616, 577, 582, 575, 567, 570, 580, 873, 651, 462, 581, 972, 973, 576, 572, 578, 574, 568, 569, 573, 615]. Swarming [566]. Swarms [483]. Sweden [951]. Swedish [514]. Sweep [901]. Sweep-Line [901]. Symbolic [99, 205]. Synchronization [157]. Synchronous [799]. Synergia [232]. Syntactic [919, 704]. Syntactic-Semantic [919]. Syntax [689]. Synthesis [310, 796]. System [229, 266, 631, 490, 917, 906, 520, 800, 472, 617, 535, 622, 687, 521, 264, 488, 750, 359, 861, 742, 464, 600, 584, 462, 761, 843, 494, 587, 483, 925, 588, 864]. System-Level [687]. Systematic [558, 686]. Systems [319, 964, 512, 664, 939, 905, 962, 388, 607, 978, 639, 958, 294, 106, 844, 66, 720, 413, 398, 800, 336, 3, 657, 83, 153, 653, 838, 907, 6, 73, 968, 986, 286, 218, 8, 482, 799, 502, 238, 491, 702, 165, 969, 70, 291, 418, 216, 399, 275, 483, 305, 673].

Tables [127]. Tablets [125]. Tackling [65]. Tactile [359]. Tag [186]. Tag-Based [186]. TAP [948]. Target [666, 484]. Task [237, 697, 236, 227, 219]. Task-Based [237, 236]. Tasking [159, 235]. Tasks [224, 202]. Tax [828]. Taxi [352]. Taxonomy [708]. Teaching [299, 155]. Team [79, 24]. Team-Based [24]. TeamSkill [24]. technical [826]. Technique [25, 654, 885, 554]. Techniques [957, 527, 26, 199, 236, 918, 669, 618]. Technologies [951, 82, 968, 980]. Technology [74, 830, 98, 592]. Telecentric [535]. Teleo [386]. Teleo-reactive [386]. Teleportation [433]. Template [326, 468, 87, 544]. Template-Based [544]. Temporal [218, 108, 203, 868, 317]. tenant [513]. Term [820, 547]. Terminating [905]. Terms [655, 454, 324]. Test [309, 101, 106, 326, 694, 84, 132, 814, 131, 107, 201]. Testing [100, 929, 65, 109, 102, 960, 927, 103, 112, 113, 98, 104, 904, 99, 380, 933]. Tests [443, 103, 948]. Testsuite [239]. TexMo [936]. Text [190, 696, 703, 23, 38, 759, 52, 175, 34]. Texts [506]. Textual [941]. Texture [756, 530]. Their [782, 290, 692, 358, 64, 763, 245, 199, 69, 819]. Them [357, 358]. Theorem [805]. Theoretic [274, 430]. Theory [964, 889, 313, 988, 798, 960, 950]. There [435]. Third [971, 959, 980, 972, 973]. Thread [223]. threaded [609]. Threat [405]. Three [789]. Throttled [498]. TI [765]. Tiles [106]. Time [25, 427, 295, 272, 80, 800, 386, 83, 153, 163, 364, 267, 87, 900, 861, 88, 61, 867, 283, 165, 22, 152, 761, 421, 333, 868, 45, 547]. Time-Evolving [22]. Timed [388, 387]. Timely [153]. TimeSquare [80]. Tips [546]. Tissue [547]. TLA [385]. Tokens [775]. Tomographic [732]. Tomography [882]. Tool [671, 102, 78, 79, 546, 536, 868]. Toolbox [866]. TOOLS [947, 77, 946, 661]. Top [191, 723]. Top- [723]. Top-K [191]. Topological [178]. Topology [952, 884, 175, 394]. Toronto [953]. Total [726]. Touching [125]. Trace [108]. Traceability [139]. Traces [352]. TraceVis [139]. Tracking [475, 834, 262, 547]. Tracts [133]. Trading [481]. Traditional [822]. Traffic [352, 355, 668]. Trail [331]. Training [531, 295, 192, 580, 31, 827, 611, 305]. Trajectories [862]. Trajectory [713]. Transactional [225, 67, 493].

Transactions [225, 260]. Transapical [552]. Transfer [202]. Transfers [498]. Transform [279, 762]. Transform-Domain [279]. Transformation [621, 880, 141, 136, 146, 140, 517, 135, 142, 148]. Transformations [134, 143, 377, 638, 373, 138, 374, 132, 950, 73, 391, 938, 930, 131, 380, 133, 139]. Transformed [863]. Transition [860]. Transitivity [214]. Translated [190]. Translating [385]. Translation [398, 890, 234, 199]. Translational [917, 142]. Translations [190]. Transport [629]. Transportation [861, 863, 864]. Trapdoor [784]. Treat [80]. Treatment [647]. Treaty [201]. Tree [807, 5, 35, 162, 783, 529, 18]. Trees [32, 492, 90, 34]. Trento [960]. TreVisor [777]. Tri [62]. Tri-concepts [62]. Triangle [179]. Triangulated [169, 176, 877]. Triangulations [620]. Trick [809]. Triggers [851]. Trips [351]. TRUST [985, 964, 838, 850, 847, 985]. Trust-Properties [847]. Trusted [833, 786, 837, 834]. Trustworthy [985, 842]. Tumor [735]. Tumours [734]. tuning [609]. Turbines [578]. Turkey [987]. Tutorial [292, 113]. Tutoring [958, 294, 286, 488, 287, 291]. Tweakable [779]. Tweets [189]. Twitter [49]. Two [266, 705, 415, 575, 367, 8, 28, 461, 265]. Two-Class [266]. Two-Layer [575]. Two-Party [415]. Two-Phase [8]. Two-Stage [705]. Two-View [28]. Type [140, 648]. Type-Safe [140]. Typed [805]. Types [256, 728, 250]. Typestate [90]. Typical [426]. Typing [133].

U [696]. U-Compare [696]. UA [928]. Ubiquitous [662]. UCT [838]. UIMA [696]. UK [964, 961]. Ultra [771]. Ultrasound [555, 556, 546, 548]. UML [672, 929, 109, 110, 926, 92, 922, 370, 668, 675, 935, 677]. UML/OCL [926]. Uncertain [712, 714, 57, 711]. Uncertainty [283]. Unclonable [839]. Unconstrained [461]. Uncovering [219]. Undecidability [15]. Understandability [675]. Understandable [519]. Understanding [328, 504, 292, 84, 303, 392, 210, 851]. Unfold [554]. Uni [432]. Uni- [432]. Unified [216, 353, 780]. Uniform [512, 171]. Uniformly [722]. Unify [795]. Unifying [804, 89]. Uniqueness [93]. Unit [163]. Units [74]. University [631]. Unknown [55, 457]. Unmanned [483]. Unpacking [210]. Unpredictable [538]. Unrequited [432]. Unstructured [913]. Unsupervised [736]. Untrusted [844]. Uppaal [123]. Urban [634, 754, 635, 352, 638, 637, 354, 633]. USA [989]. Usability [558]. Usage [363, 663]. Use [558, 297, 637, 160, 247, 748]. User [411, 838, 850, 412, 52, 421, 669, 851]. User-Centered [411]. User-Centered-Trust [838]. User-Defined [421]. Users [629]. Using [531, 442, 443, 496, 188, 672, 101, 752, 655, 706, 116, 705, 616, 352, 53, 621, 582, 434, 880, 545, 120, 629, 524, 758, 739, 643, 456, 475, 704, 163, 74, 550, 401, 233, 367, 46, 708, 770, 473, 49, 478, 890, 912, 57, 90, 323, 771, 71, 827, 238, 301, 815, 450, 678, 172, 623, 733, 763, 392, 433, 462, 15, 131, 107, 124, 232, 757, 679, 668, 834, 95, 872, 440, 122, 539, 476, 133, 408, 171, 318, 205, 477, 219, 289, 776, 725, 615, 677, 892]. UTP [396]. UWB [768].

Vacation [267, 268]. Vadis [212]. Validating [916]. Validation [385, 926, 239, 375]. Value [818, 9, 508, 790]. Values [635, 630, 447, 763]. Valve [552]. VANET [270]. Variability [513, 275]. Variable [585]. Variance [279, 659]. Variates [282]. Variational [476]. Various [401]. VDM [959, 325]. Vector [226, 859, 733]. Vectors [434, 276, 17]. Vehicle [570, 216]. Vehicles [74, 349, 863]. Vendor [85]. Ventilation [661]. Ventricle [733]. Veracity [406]. Verbs [692]. Verification [671, 389, 65, 384, 537, 90, 391, 623, 508, 201, 399, 305]. Verified [102]. Verifying [164, 103, 847, 843]. Version [247]. Versioning [379]. versus [330, 2]. Vertical [903, 855, 938]. Vertices [436]. Very [727]. Vessel [731]. via [788, 180, 916, 670, 452, 856, 195, 342]. Vickers [456]. Video [302, 622, 499, 623, 761]. Video-Based [623]. Video-Game [302]. Vienna [985]. View [550, 28]. Vinci [546]. Virtual [496, 559, 267, 579, 89, 827, 487, 142, 289]. Virtualization [844]. Visible [460, 51]. Vision [535, 763]. Visual [641, 448, 599]. Visualization [555, 291, 139]. VMs [271]. Voice [398]. Volumetric [880]. Voronoi [640]. Voronoi-Based [640]. Voter [424]. Voxel [473]. Voxel-Based [473]. vs [296].

WADGMM [987]. Waiting [267]. Wall [540]. Wall-Following [540]. Warehouse [656, 660]. Watson [913]. Wavelength [460]. Wavelet [454, 576, 476]. Wavelet-Based [576]. WAW [965]. Way [681, 265, 19]. WCCI [954]. Weak [449]. Weakly [905]. Web [965, 676, 307, 383, 670, 312, 707, 846, 596, 487, 592]. Wedding [626]. Weight [572]. Weighted [25, 50, 614, 46]. Weights [46]. WeightTransmitter [46]. Well [354, 370]. Well-Being [354]. WG [964]. Which [487]. while [435]. White [734]. Who [358]. Wide [771, 792]. Wide-Band [771]. WiFi [365]. Wikimantic [701]. Wikipedia [708]. Wild [358]. Wind [578]. Wire [363]. Wireless [775, 384, 715, 393, 483, 394]. WISTP [964]. Within [357, 356, 906, 907, 118, 291]. without [786, 26]. Won’t [720]. Wood [649]. Words [7, 205]. Work [850]. Workflow [895, 893, 894]. Workflows [696, 498]. Workload [294]. Works [635]. Workshop [964, 965, 955, 952, 987]. World [955, 954, 199]. Worn [363]. Wound [742]. WRF [231]. Writing [190, 884]. Written [819]. WSNs [666]. WSTS [889].

X [551]. X-Ray [551]. X10 [252]. XML [932, 135].

Z [959]. Zero [790]. Zero-Value [790]. Zone [455].

References

Balan:2012:PC

[1] M. Sakthi Balan and Helmut Jürgensen. Peptide computers. Lecture Notes in Computer Science, 7300:1–29, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31644-9_1/.

Bockenhauer:2012:PRV

[2] Hans-Joachim Böckenhauer, Juraj Hromkovič, Dennis Komm, and Richard Královič. On the power of randomness versus advice in online computation. Lecture Notes in Computer Science, 7300:30–43, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31644-9_2/.

Ehrenfeucht:2012:RER

[3] Andrzej Ehrenfeucht, Jetty Kleijn, Maciej Koutny, and Grzegorz Rozenberg. Relevance of entities in reaction systems. Lecture Notes in Computer Science, 7300:44–55, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31644-9_3/.

Ewert:2012:GRC

[4] Sigrid Ewert and Max Rabkin. Generalized random context picture grammars: The state of the art. Lecture Notes in Computer Science, 7300: 56–74, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31644-9_4/.

Fernau:2012:CDT

[5] Henning Fernau. Cooperating distributed tree automata. Lecture Notes in Computer Science, 7300:75–85, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31644-9_5/.

Holzer:2012:NCD

[6] Markus Holzer. A note on combined derivation modes for cooperating distributed grammar systems. Lecture Notes in Computer Science, 7300:86–98, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31644-9_6/.

Huova:2012:EPS

[7] Mari Huova and Juhani Karhumäki. Equations in the partial semigroup of words with overlapping products. Lecture Notes in Computer Science, 7300:99–110, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31644-9_7/.

Kutrib:2012:CSS

[8] Martin Kutrib and Friedrich Otto. On CD-systems of stateless deterministic two-phase RR(1)-automata. Lecture Notes in Computer Science, 7300: 111–137, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31644-9_8/.

Lange:2012:BFV

[9] Klaus-Jörn Lange. The Boolean formula value problem as formal language. Lecture Notes in Computer Science, 7300:138–144, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31644-9_9/.

Manea:2012:HLS

[10] Florin Manea, Robert Mercas, and Victor Mitrana. Hairpin lengthening and shortening of regular languages. Lecture Notes in Computer Science, 7300:145–159, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-31644-9_10/.

Meduna:2012:OSR

[11] Alexander Meduna and Petr Zemek. One-sided random context grammars with leftmost derivations. Lecture Notes in Computer Science, 7300: 160–173, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31644-9_11/.

Nishida:2012:EPA

[12] Taishin Y. Nishida. Earley’s parsing algorithm and k-Petri net controlled grammars. Lecture Notes in Computer Science, 7300:174–185, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31644-9_12/.

Okhotin:2012:DCI

[13] Alexander Okhotin, Xiaoxue Piao, and Kai Salomaa. Descriptional complexity of input-driven pushdown automata. Lecture Notes in Computer Science, 7300:186–206, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31644-9_13/.

Paun:2012:TFM

[14] Gheorghe Păun. Towards “Fypercomputations” (in membrane computing). Lecture Notes in Computer Science, 7300:207–220, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31644-9_14/.

Salomaa:2012:USC

[15] Arto Salomaa. Undecidability of state complexities using mirror images. Lecture Notes in Computer Science, 7300:221–235, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31644-9_15/.

Staiger:2012:ASC

[16] Ludwig Staiger. Asymptotic subword complexity. Lecture Notes in Computer Science, 7300:236–245, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31644-9_16/.

Stiebe:2012:GCP

[17] Ralf Stiebe. On grammars controlled by Parikh vectors. Lecture Notes in Computer Science, 7300:246–264, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31644-9_17/.

Vaszil:2012:NCT

[18] György Vaszil. On the nonterminal complexity of tree controlled grammars. Lecture Notes in Computer Science, 7300:265–272, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31644-9_18/.

Zheng:2012:OWF

[19] Shenggen Zheng, Daowen Qiu, Lvzhou Li, and Jozef Gruska. One-way finite automata with quantum and classical states. Lecture Notes in Computer Science, 7300:273–290, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31644-9_19/.

Anonymous:2012:BMa

[20] Anonymous. Back matter. Lecture Notes in Computer Science, 7300: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-31644-9/1.

Anonymous:2012:FMa

[21] Anonymous. Front matter. Lecture Notes in Computer Science, 7300: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31644-9/1.

Rossi:2012:TER

[22] Ryan Rossi and Jennifer Neville. Time-evolving relational classification and ensemble methods. Lecture Notes in Computer Science, 7301: 1–13, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30217-6_1/.

Li:2012:ALH

[23] Xiao Li, Da Kuang, and Charles X. Ling. Active learning for hierarchical text classification. Lecture Notes in Computer Science, 7301: 14–25, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30217-6_2/.

DeLong:2012:TEM

[24] Colin DeLong and Jaideep Srivastava. TeamSkill evolved: Mixed classification schemes for team-based multi-player games. Lecture Notes in Computer Science, 7301:26–37, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30217-6_3/.

Adhikari:2012:NWE

[25] Ratnadip Adhikari and R. K. Agrawal. A novel weighted ensemble technique for time series forecasting. Lecture Notes in Computer Science, 7301:38–49, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30217-6_4/.

Salem:2012:TEL

[26] Houssam Salem, Pramuditha Suraweera, Geoffrey I. Webb, and Janice R. Boughton. Techniques for efficient learning without search. Lecture Notes in Computer Science, 7301:50–61, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30217-6_5/.

Fu:2012:AMB

[27] JuiHsi Fu and SingLing Lee. An aggressive margin-based algorithm for incremental learning. Lecture Notes in Computer Science, 7301: 62–73, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30217-6_6/.

Nguyen:2012:TVO

[28] Tam T. Nguyen, Kuiyu Chang, and Siu Cheung Hui. Two-view online learning. Lecture Notes in Computer Science, 7301:74–85, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30217-6_7/.

Liao:2012:GCE

[29] Zhihua Liao and Zili Zhang. A generic classifier-ensemble approach for biomedical named entity recognition. Lecture Notes in Computer Science, 7301:86–97, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30217-6_8/.

Zighed:2012:NRC

[30] Djamel Abdelkader Zighed, Diala Ezzeddine, and Fabien Rico. Neighborhood random classification. Lecture Notes in Computer Science, 7301: 98–108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30217-6_9/.

Mirylenka:2012:SFS

[31] Katsiaryna Mirylenka, George Giannakopoulos, and Themis Palpanas. SRF: a framework for the study of classifier behavior under training set mislabeling noise. Lecture Notes in Computer Science, 7301:109– 121, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30217-6_10/.

Hoens:2012:BDT

[32] T. Ryan Hoens, Qi Qian, Nitesh V. Chawla, and Zhi-Hua Zhou. Building decision trees for the multi-class imbalance problem. Lecture Notes in Computer Science, 7301:122–134, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30217-6_11/.

Li:2012:SRF

[33] Bingguo Li, Xiaojun Chen, Mark Junjie Li, Joshua Zhexue Huang, and Shengzhong Feng. Scalable random forests for massive data. Lecture Notes in Computer Science, 7301:135–146, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30217-6_12/.

Xu:2012:HRF

[34] Baoxun Xu, Joshua Zhexue Huang, Graham Williams, Mark Junjie Li, and Yunming Ye. Hybrid random forests: Advantages of mixed trees in classifying text data. Lecture Notes in Computer Science, 7301: 147–158, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30217-6_13/.

Fu:2012:LTS

[35] Bin Fu, Zhihai Wang, Rong Pan, Guandong Xu, and Peter Dolog. Learning tree structure of label dependency for multi-label learning. Lecture Notes in Computer Science, 7301:159–170, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30217-6_14/.

Fu:2012:MIL

[36] Zhichun Fu, Jun Zhou, Peter Christen, and Mac Boot. Multiple instance learning for group record linkage. Lecture Notes in Computer Science, 7301:171–182, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-30217-6_15/.

Shirai:2012:ISR

[37] Yasuyuki Shirai, Koji Tsuruma, Yuko Sakurai, Satoshi Oyama, and Shin ichi Minato. Incremental set recommendation based on class differences. Lecture Notes in Computer Science, 7301:183–194, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30217-6_16/.

Liu:2012:ALC

[38] Yue Liu, Lin Dai, Weitao Zhou, and Heyan Huang. Active learning for cross language text categorization. Lecture Notes in Computer Science, 7301:195–206, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-30217-6_17/.

Xiao:2012:EAM

[39] Han Xiao, Thomas Stibor, and Claudia Eckert. Evasion attack of multi-class linear classifiers. Lecture Notes in Computer Science, 7301: 207–218, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30217-6_18/.

Kuang:2012:FMC

[40] Da Kuang, Charles X. Ling, and Jun Du. Foundation of mining classimbalanced data. Lecture Notes in Computer Science, 7301:219–230, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30217-6_19/.

Ni:2012:ALC

[41] Eileen A. Ni and Charles X. Ling. Active learning with c-certainty. Lecture Notes in Computer Science, 7301:231–242, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30217-6_20/.

Anonymous:2012:FMb

[42] Anonymous. Front matter. Lecture Notes in Computer Science, 7301: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-30217-6/1.

Nguyen:2012:HEF

[43] Hai-Long Nguyen, Yew-Kwong Woon, Wee-Keong Ng, and Li Wan. Heterogeneous ensemble for feature drifts in data streams. Lecture Notes in Computer Science, 7302:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30220-6_1/.

Brahmi:2012:OIC

[44] Hanen Brahmi, Imen Brahmi, and Sadok Ben Yahia. OMC-IDS: At the cross-roads of OLAP mining and intrusion detection. Lecture Notes in Computer Science, 7302:13–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30220-6_2/.

Xie:2012:TLT

[45] Jierui Xie and Boleslaw K. Szymanski. Towards linear time overlapping community detection in social networks. Lecture Notes in Computer Science, 7302:25–36, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-30220-6_3/.

Koh:2012:WWA

[46] Yun Sing Koh, Russel Pears, and Gillian Dobbie. WeightTransmitter: Weighted association rule mining using landmark weights. Lecture Notes in Computer Science, 7302:37–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30220-6_4/.

Inaba:2012:COC

[47] Daiki Inaba, Ken ichi Fukui, Kazuhisa Sato, Junichirou Mizusaki, and Masayuki Numao. Co-occurring cluster mining for damage patterns analysis of a fuel cell. Lecture Notes in Computer Science, 7302:49– 60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30220-6_5/.

Bouasker:2012:NEC

[48] Souad Bouasker, Tarek Hamrouni, and Sadok Ben Yahia. New exact concise representation of rare correlated patterns: Application to intrusion detection. Lecture Notes in Computer Science, 7302:61–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30220-6_6/.

Lu:2012:LAM

[49] Rong Lu, Zhiheng Xu, Yang Zhang, and Qing Yang. Life activity modeling of news event on Twitter using energy function. Lecture Notes in Computer Science, 7302:73–84, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30220-6_7/.

Akoglu:2012:QRL

[50] Leman Akoglu, Pedro O. S. Vaz de Melo, and Christos Faloutsos. Quantifying reciprocity in large weighted communication networks. Lecture Notes in Computer Science, 7302:85–96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30220-6_8/.

Yan:2012:HGS

[51] Rui Yan, Zi Yuan, Xiaojun Wan, Yan Zhang, and Xiaoming Li. Hierarchical graph summarization: Leveraging hybrid information through visible and invisible linkage. Lecture Notes in Computer Science, 7302: 97–108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30220-6_9/.

Peng:2012:MMU

[52] Bingyue Peng, Yujing Wang, and Jian-Tao Sun. Mining mobile users’ activities based on search query text and context. Lecture Notes in Computer Science, 7302:109–120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30220-6_10/.

Chaudhury:2012:SIS

[53] Arpan Chaudhury, Partha Basuchowdhuri, and Subhashis Majumder. Spread of information in a social network using influential nodes. Lecture Notes in Computer Science, 7302:121–132, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30220-6_11/.

Srinivas:2012:DCP

[54] P. Gowtham Srinivas, P. Krishna Reddy, S. Bhargav, R. Uday Kiran, and D. Satheesh Kumar. Discovering coverage patterns for banner advertisement placement. Lecture Notes in Computer Science, 7302: 133–144, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30220-6_12/.

Duan:2012:DUI

[55] Juang-Lin Duan, Shashi Prasad, and Jen-Wei Huang. Discovering unknown but interesting items on personal social network. Lecture Notes in Computer Science, 7302:145–156, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30220-6_13/.

Salas:2012:PND

[56] Hugo Alatrista Salas, Sandra Bringay, Frédéric Flouvat, and Nazha Selmaoui-Folcher. The pattern next door: Towards spatio-sequential pattern discovery. Lecture Notes in Computer Science, 7302:157–168, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30220-6_14/.

Matsumoto:2012:AOD

[57] Takazumi Matsumoto and Edward Hung. Accelerating outlier detection with uncertain data using graphics processors. Lecture Notes in Computer Science, 7302:169–180, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30220-6_15/.

Mougel:2012:FCC

[58] Pierre-Nicolas Mougel, Christophe Rigotti, and Olivier Gandrillon. Finding collections of k-clique percolated components in attributed graphs. Lecture Notes in Computer Science, 7302:181–192, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.com/chapter/10.1007/978-3-642-30220-6_16/.

Cai:2012:RHL

[59] Xiongcai Cai, Michael Bain, Alfred Krzywicki, Wayne Wobcke, Yang Sok Kim, and Paul Compton. Reciprocal and heterogeneous link prediction in social networks. Lecture Notes in Computer Science, 7302: 193–204, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30220-6_17/.

Liu:2012:DMS

[60] Kai Liu, William K. Cheung, and Jiming Liu. Detecting multiple stochastic network motifs in network data. Lecture Notes in Computer Science, 7302:205–217, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-30220-6_18/.

Marascu:2012:SSM

[61] Alice Marascu, Suleiman A. Khan, and Themis Palpanas. Scalable similarity matching in streaming time series. Lecture Notes in Computer Science, 7302:218–230, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30220-6_19/.

Trabelsi:2012:SMF

[62] Chiraz Trabelsi, Nader Jelassi, and Sadok Ben Yahia. Scalable mining of frequent tri-concepts from folksonomies. Lecture Notes in Computer Science, 7302:231–242, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30220-6_20/.

Anonymous:2012:FMc

[63] Anonymous. Front matter. Lecture Notes in Computer Science, 7302: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-30220-6/1.

Meyer:2012:PTC

[64] Bertrand Meyer, Alexander Kogtenkov, and Anton Akhi. Processors and their collection. Lecture Notes in Computer Science, 7303:1–15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31202-1_1/.

Briand:2012:TTV

[65] Lionel C. Briand. Tackling the testing and verification of multicore and concurrent software as a search problem. Lecture Notes in Computer Science, 7303:16–17, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/accesspage/
chapter/10.1007/978-3-642-31202-1_2.

Christmann:2012:OCR

[66] Constantin Christmann, Erik Hebisch, and Anette Weisbecker. Oversubscription of computational resources on multicore desktop systems. Lecture Notes in Computer Science, 7303:18–29, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31202-1_3/.

Schindewolf:2012:CTM

[67] Martin Schindewolf and Wolfgang Karl. Capturing transactional memory Application’s behavior — the prerequisite for performance analysis. Lecture Notes in Computer Science, 7303:30–41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.com/chapter/10.1007/978-3-642-31202-1_4/.

Waller:2012:CID

[68] Jan Waller and Wilhelm Hasselbring. A comparison of the influence of different multi-core processors on the runtime overhead for application-level monitoring. Lecture Notes in Computer Science, 7303: 42–53, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31202-1_5/.

Strebelow:2012:AEP

[69] Ronald Strebelow and Christian Prehofer. Analysis of event processing design patterns and their performance dependency on I/O notification mechanisms. Lecture Notes in Computer Science, 7303:54–65, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31202-1_6/.

Schonherr:2012:NIC

[70] Jan H. Schönherr, Bianca Lutz, and Jan Richling. Non-intrusive coscheduling for general purpose operating systems. Lecture Notes in Computer Science, 7303:66–77, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31202-1_7/.

Molitorisz:2012:APU

[71] Korbinian Molitorisz, Jochen Schimmel, and Frank Otto. Automatic parallelization using AutoFutures. Lecture Notes in Computer Science, 7303:78–81, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31202-1_8/.

Hollmann:2012:ICA

[72] Andreas Hollmann and Michael Gerndt. Invasive computing: An application assisted resource management approach. Lecture Notes in Computer Science, 7303:82–85, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31202-1_9/.

Imre:2012:PGT

[73] Gábor Imre and Gergely Mezei. Parallel graph transformations on multicore systems. Lecture Notes in Computer Science, 7303:86–89, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31202-1_10/.

Gut:2012:REC

[74] Georg Gut, Christian Allmann, Markus Schurius, and Karsten Schmidt. Reduction of electronic control units in electric vehicles using multicore technology. Lecture Notes in Computer Science, 7303:90–93, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31202-1_11/.

Anonymous:2012:BMb

[75] Anonymous. Back matter. Lecture Notes in Computer Science, 7303: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-31202-1/1.

Anonymous:2012:FMd

[76] Anonymous. Front matter. Lecture Notes in Computer Science, 7303: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31202-1/1.

Bergmann:2012:IEM

[77] Gábor Bergmann, Ábel Hegedüs, and Ákos Horváth. Integrating efficient model queries in state-of-the-art EMF tools. Lecture Notes in Computer Science, 7304:1–8, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-30561-0_1/.

Catano:2012:PFM

[78] Néstor Cataño, Sorren Hanvey, and Camilo Rueda. Poporo: a formal methods tool for fast-checking of social network privacy policies. Lecture Notes in Computer Science, 7304:9–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30561-0_2/.

Corral:2012:DMT

[79] Luis Corral, Alberto Sillitti, and Giancarlo Succi. DroidSense: a mobile tool to analyze software development processes by measuring team proximity. Lecture Notes in Computer Science, 7304:17–33, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30561-0_3/.

DeAntoni:2012:TTY

[80] Julien DeAntoni and Frédéric Mallet. TimeSquare: Treat your models with logical time. Lecture Notes in Computer Science, 7304:34– 41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30561-0_4/.

Derezinska:2012:QEO

[81] Anna Derezińska and Marcin Rudnik. Quality evaluation of objectoriented and standard mutation operators applied to C# programs. Lecture Notes in Computer Science, 7304:42–57, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-30561-0_5/.

Favre:2012:CPS

[82] Jean-Marie Favre, Ralf Lämmel, and Thomas Schmorleiz. 101companies: a community project on software technologies and software languages. Lecture Notes in Computer Science, 7304:58–74, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30561-0_6/.

Fiamberti:2012:OOA

[83] Francesco Fiamberti and Daniela Micucci. An object-oriented application framework for the development of real-time systems. Lecture Notes in Computer Science, 7304:75–90, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30561-0_7/.

Greiler:2012:MTC

[84] Michaela Greiler, Arie van Deursen, and Andy Zaidman. Measuring test case similarity to support test suite understanding. Lecture Notes in Computer Science, 7304:91–107, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30561-0_8/.

Jezek:2012:EOE

[85] Kamil Ježek, Premek Brada, and Lukáš Holý. Enhancing OSGi with explicit, vendor independent extra-functional properties. Lecture Notes in Computer Science, 7304:108–123, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30561-0_9/.

Vrany:2012:EML

[86] Jan Vraný, Jan Kurš, and Claus Gittinger. Efficient method lookup customization for Smalltalk. Lecture Notes in Computer Science, 7304: 124–139, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30561-0_10/.

Langr:2012:FRT

[87] Daniel Langr, Pavel Tvrdík, and Tomáš Dytrych. Fake run-time selection of template arguments in C++. Lecture Notes in Computer Science, 7304: 140–154, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30561-0_11/.

Lilis:2012:SCT

[88] Yannis Lilis and Anthony Savidis. Supporting compile-time debugging and precise error reporting in meta-programs. Lecture Notes in Computer Science, 7304:155–170, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30561-0_12/.

Marr:2012:IUM

[89] Stefan Marr and Theo D’Hondt. Identifying a unifying mechanism for the implementation of concurrency abstractions on multi-language virtual machines. Lecture Notes in Computer Science, 7304:171–186, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30561-0_13/.

Mehnert:2012:VST

[90] Hannes Mehnert and Jonathan Aldrich. Verification of snapshotable trees using access permissions and typestate. Lecture Notes in Computer Science, 7304:187–201, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30561-0_14/.

Ng:2012:MSC

[91] Nicholas Ng, Nobuko Yoshida, and Kohei Honda. Multiparty session C: Safe parallel programming with message optimisation. Lecture Notes in Computer Science, 7304:202–218, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30561-0_15/.

Ochoa:2012:NIU

[92] Martín Ochoa, Jan Jürjens, and Jorge Cuéllar. Non-interference on UML state-charts. Lecture Notes in Computer Science, 7304:219–235, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30561-0_16/.

Olah:2012:RUC

[93] Mark J. Olah, David Mohr, and Darko Stefanovic. Representing uniqueness constraints in object-relational mapping. Lecture Notes in Computer Science, 7304:236–251, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30561-0_17/.

Olszak:2012:DSM

[94] Andrzej Olszak and Eric Bouwers. Detection of seed methods for quantification of feature confinement. Lecture Notes in Computer Science, 7304: 252–268, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30561-0_18/.

Soeken:2012:ABD

[95] Mathias Soeken, Robert Wille, and Rolf Drechsler. Assisted behavior driven development using natural language processing. Lecture Notes in Computer Science, 7304:269–287, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30561-0_19/.

Sureka:2012:LCB

[96] Ashish Sureka. Learning to classify bug reports into components. Lecture Notes in Computer Science, 7304:288–303, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30561-0_20/.

Anonymous:2012:FMe

[97] Anonymous. Front matter. Lecture Notes in Computer Science, 7304: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-30561-0/1.

Kuehlmann:2012:TPT

[98] Andreas Kuehlmann. The technology and psychology of testing your code as you develop it. Lecture Notes in Computer Science, 7305:1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-30473-6_1.

Pasareanu:2012:CMC

[99] Corina S. Păsăreanu. Combining model checking and symbolic execution for software testing. Lecture Notes in Computer Science, 7305:2, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-30473-6_2.

Armando:2012:MCA

[100] Alessandro Armando, Giancarlo Pellegrino, Roberto Carbone, and Alessio Merlo. From model-checking to automated testing of security protocols: Bridging the gap. Lecture Notes in Computer Science, 7305: 3–18, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30473-6_3/.

Bengolea:2012:UCC

[101] Valeria Bengolea, Nazareno Aguirre, Darko Marinov, and Marcelo F. Frias. Using coverage criteria on RepOK to reduce bounded-exhaustive test suites. Lecture Notes in Computer Science, 7305:19–34, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30473-6_4/.

Carlier:2012:FSD

[102] Matthieu Carlier, Catherine Dubois, and Arnaud Gotlieb. A first step in the design of a formally verified constraint-based testing tool: Focal-Test. Lecture Notes in Computer Science, 7305:35–50, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30473-6_5/.

Kiniry:2012:TLS

[103] Joseph R. Kiniry, Daniel M. Zimmerman, and Ralph Hyland. Testing library specifications by verifying conformance tests. Lecture Notes in Computer Science, 7305:51–66, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30473-6_6/.

Lochau:2012:IMB

[104] Malte Lochau, Ina Schaefer, Jochen Kamischke, and Sascha Lity. Incremental model-based testing of delta-oriented software product lines. Lecture Notes in Computer Science, 7305:67–82, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-30473-6_7/.

deLeon:2012:CRL

[105] Hernán Ponce de León, Stefan Haar, and Delphine Longuet. Conformance relations for labeled event structures. Lecture Notes in Computer Science, 7305:83–98, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30473-6_8/.

Chedor:2012:TGR

[106] Sébastien Chédor, Thierry Jéron, and Christophe Morvan. Test generation from recursive tiles systems. Lecture Notes in Computer Science, 7305: 99–114, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30473-6_9/.

Senni:2012:GTD

[107] Valerio Senni and Fabio Fioravanti. Generation of test data structures using constraint logic programming. Lecture Notes in Computer Science, 7305:115–131, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-30473-6_10/.

Sulzmann:2012:CFT

[108] Martin Sulzmann and Axel Zechner. Constructive finite trace analysis with linear temporal logic. Lecture Notes in Computer Science, 7305: 132–148, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30473-6_11/.

Brosch:2012:TSB

[109] Petra Brosch, Uwe Egly, Sebastian Gabmeyer, Gerti Kappel, Martina Seidl, and Hans Tompits. Towards scenario-based testing of UML diagrams. Lecture Notes in Computer Science, 7305:149–155, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30473-6_12/.

Bruning:2012:EDO

[110] Jens Brüning, Martin Gogolla, Lars Hamann, and Mirco Kuhlmann. Evaluating and debugging OCL expressions in UML models. Lecture Notes in Computer Science, 7305:156–162, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30473-6_13/.

Creignou:2012:FSR

[111] Nadia Creignou, Uwe Egly, and Martina Seidl. A framework for the specification of random SAT and QSAT formulas. Lecture Notes in Computer Science, 7305:163–168, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30473-6_14/.

Kosmatov:2012:LST

[112] Nikolai Kosmatov, Nicky Williams, Bernard Botella, Muriel Roger, and Omar Chebaro. A lesson on structural testing with PathCrawler-online.com. Lecture Notes in Computer Science, 7305:169–175, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.com/chapter/10.1007/978-3-642-30473-6_15/.

Kosmatov:2012:TAS

[113] Nikolai Kosmatov and Nicky Williams. Tutorial on automated structural testing with PathCrawler. Lecture Notes in Computer Science, 7305:176, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-30473-6_16.

Anonymous:2012:BMc

[114] Anonymous. Back matter. Lecture Notes in Computer Science, 7305: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-30473-6/1.

Anonymous:2012:FMf

[115] Anonymous. Front matter. Lecture Notes in Computer Science, 7305: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-30473-6/1.

Bozga:2012:MDA

[116] Marius Bozga, Mohamad Jaber, Nikolaos Maris, and Joseph Sifakis. Modeling dynamic architectures using dy-BIP. Lecture Notes in Computer Science, 7306:1–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30564-1_1/.

Istoan:2012:DCO

[117] Paul Istoan. Defining composition operators for BPMN. Lecture Notes in Computer Science, 7306:17–34, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30564-1_2/.

Lanoix:2012:RBS

[118] Arnaud Lanoix, Olga Kouchnarenko, Samuel Colin, and Vincent Poirriez. Relaxing B sharing restrictions within CSP——B. Lecture Notes in Computer Science, 7306:35–50, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30564-1_3/.

Caporuscio:2012:PDF

[119] Mauro Caporuscio, Marco Funaro, and Carlo Ghezzi. PaCE: a data-flow coordination language for asynchronous network-based applications. Lecture Notes in Computer Science, 7306:51–67, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-30564-1_4/.

Danylenko:2012:ALC

[120] Antonina Danylenko and Welf Löwe. Adaptation of legacy codes to context-aware composition using aspect-oriented programming. Lecture Notes in Computer Science, 7306:68–85, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30564-1_5/.

Ansaloni:2012:CRC

[121] Danilo Ansaloni, Walter Binder, Christoph Bockisch, Eric Bodden, and Kardelen Hatun. Challenges for refinement and composition of instrumentations: Position paper. Lecture Notes in Computer Science, 7306: 86–96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30564-1_6/.

Timbermont:2012:CCI

[122] Stijn Timbermont, Coen De Roover, and Theo D’Hondt. Constructing customized interpreters from reusable evaluators using game. Lecture Notes in Computer Science, 7306:97–113, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30564-1_7/.

Boudjadar:2012:REU

[123] Abdeldjalil Boudjadar, Jean-Paul Bodeveix, and Mamoun Filali. Revising and extending the Uppaal communication mechanism. Lecture Notes in Computer Science, 7306:114–131, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30564-1_8/.

Shah:2012:AMJ

[124] Syed Muhammad Ali Shah, Jens Dietrich, and Catherine McCartin. On the automated modularisation of Java programs using service locators. Lecture Notes in Computer Science, 7306:132–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30564-1_9/.

Hesenius:2012:TFS

[125] Marc Hesenius, Carlos Dario Orozco Medina, and Dominikus Herzberg. Touching factor: Software development on tablets. Lecture Notes in Computer Science, 7306:148–161, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30564-1_10/.

Cazzola:2012:DSL

[126] Walter Cazzola. Domain-specific languages in few steps. Lecture Notes in Computer Science, 7306:162–177, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30564-1_11/.

Boffoli:2012:BPL

[127] Nicola Boffoli, Danilo Caivano, Daniela Castelluccia, and Giuseppe Visaggio. Business process lines and decision tables driving flexibility by selection. Lecture Notes in Computer Science, 7306:178–193, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30564-1_12/.

Anonymous:2012:BMd

[128] Anonymous. Back matter. Lecture Notes in Computer Science, 7306: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-30564-1/1.

Anonymous:2012:FMg

[129] Anonymous. Front matter. Lecture Notes in Computer Science, 7306: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-30564-1/1.

Terwilliger:2012:HCY

[130] James F. Terwilliger, Anthony Cleve, and Carlo A. Curino. How clean is your sandbox? Lecture Notes in Computer Science, 7307:

1–23, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30476-7_1/.

Sen:2012:UMP

[131] Sagar Sen, Jean-Marie Mottu, Massimo Tisi, and Jordi Cabot. Using models of partial knowledge to test model transformations. Lecture Notes in Computer Science, 7307:24–39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30476-7_2/.

Guerra:2012:SDT

[132] Esther Guerra. Specification-driven test generation for model transformations. Lecture Notes in Computer Science, 7307:40–55, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30476-7_3/.

Vallecillo:2012:TMT

[133] Antonio Vallecillo and Martin Gogolla. Typing model transformations using tracts. Lecture Notes in Computer Science, 7307:56–71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30476-7_4/.

Andova:2012:RCE

[134] Suzana Andova, Mark G. J. van den Brand, and Luc Engelen. Reusable and correct endogenous model transformations. Lecture Notes in Computer Science, 7307:72–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30476-7_5/.

Pacheco:2012:MSB

[135] Hugo Pacheco and Alcino Cunha. Multifocal: a strategic bidirectional transformation language for XML schemas. Lecture Notes in Computer Science, 7307:89–104, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30476-7_6/.

Cunha:2012:BTM

[136] Jácome Cunha, João P. Fernandes, Jorge Mendes, Hugo Pacheco, and João Saraiva. Bidirectional transformation of model-driven spreadsheets. Lecture Notes in Computer Science, 7307:105–120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30476-7_7/.

vandenBos:2012:DSO

[137] Jeroen van den Bos and Tijs van der Storm. Domain-specific optimization in digital forensics. Lecture Notes in Computer Science, 7307: 121–136, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30476-7_8/.

Fernandez-Ropero:2012:EAB

[138] María Fernández-Ropero, Ricardo Pérez-Castillo, Barbara Weber, and Mario Piattini. Empirical assessment of business model transformations based on model simulation. Lecture Notes in Computer Science, 7307: 137–151, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30476-7_9/.

vanAmstel:2012:TVM

[139] Marcel F. van Amstel, Mark G. J. van den Brand, and Alexander Serebrenik. Traceability visualization in model transformations with Trace-Vis. Lecture Notes in Computer Science, 7307:152–159, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30476-7_10/.

George:2012:TSM

[140] Lars George, Arif Wider, and Markus Scheidgen. Type-safe model transformation languages as internal DSLs in Scala. Lecture Notes in Computer Science, 7307:160–175, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30476-7_11/.

Cuadrado:2012:TFM

[141] Jesús Sánchez Cuadrado. Towards a family of model transformation languages. Lecture Notes in Computer Science, 7307:176–191, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30476-7_12/.

Wagelaar:2012:TSC

[142] Dennis Wagelaar, Ludovico Iovino, Davide Di Ruscio, and Alfonso Pierantonio. Translational semantics of a co-evolution specific language with the EMF transformation virtual machine. Lecture Notes in Computer Science, 7307:192–207, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30476-7_13/.

Atkinson:2012:TML

[143] Colin Atkinson, Ralph Gerbig, and Christian Tunjic. Towards multi-level aware model transformations. Lecture Notes in Computer Science, 7307: 208–223, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30476-7_14/.

Varro:2012:AGM

[144] Gergely Varró, Frederik Deckwerth, Martin Wieber, and Andy Schürr. An algorithm for generating model-sensitive search plans for EMF models. Lecture Notes in Computer Science, 7307:224–239, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30476-7_15/.

TranconyWidemann:2012:PPM

[145] Baltasar Trancón y Widemann and Markus Lepper. Paisley: Pattern matching à la carte. Lecture Notes in Computer Science, 7307: 240–247, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30476-7_16/.

Demuth:2012:CDM

[146] Andreas Demuth, Roberto E. Lopez-Herrejon, and Alexander Egyed. Constraint-driven modeling through transformation. Lecture Notes in Computer Science, 7307:248–263, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30476-7_17/.

Kosiuczenko:2012:ICM

[147] Piotr Kosiuczenko. The impact of class model redesign on state machines. Lecture Notes in Computer Science, 7307:264–279, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30476-7_18/.

Wimmer:2012:FFR

[148] Manuel Wimmer, Gerti Kappel, Angelika Kusel, Werner Retschitzegger, and Johannes Schönböck. Fact or fiction — reuse in rule-based modelto-model transformation languages. Lecture Notes in Computer Science, 7307:280–295, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-30476-7_19/.

Anonymous:2012:BMe

[149] Anonymous. Back matter. Lecture Notes in Computer Science, 7307: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-30476-7/1.

Anonymous:2012:FMh

[150] Anonymous. Front matter. Lecture Notes in Computer Science, 7307: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-30476-7/1.

Panunzio:2012:ARC

[151] Marco Panunzio and Tullio Vardanega. Ada Ravenscar code archetypes for component-based development. Lecture Notes in Computer Science, 7308:1–17, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30598-6_1/.

Saez:2012:IFM

[152] Sergio Sáez, Jorge Real, and Alfons Crespo. An integrated framework for multiprocessor, multimoded real-time applications. Lecture Notes in Computer Science, 7308:18–34, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30598-6_2/.

Garcia-Valls:2012:IMT

[153] Marisol García-Valls and Felipe Ibáñez-Vázquez. Integrating middleware for timely reconfiguration of distributed soft real-time systems with Ada DSA. Lecture Notes in Computer Science, 7308:35–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30598-6_3/.

Ruiz:2012:SCK

[154] José F. Ruiz, Cyrille Comar, and Yannick Moy. Source code as the key artifact in requirement-based development: The case of Ada 2012. Lecture Notes in Computer Science, 7308:49–59, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30598-6_4/.

Tempelmeier:2012:TCP

[155] Theodor Tempelmeier. Teaching ‘concepts of programming languages’ with Ada. Lecture Notes in Computer Science, 7308:60–74, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30598-6_5/.

Forler:2012:DAC

[156] Christian Forler, Stefan Lucks, and Jakob Wenzel. Designing the API for a cryptographic library. Lecture Notes in Computer Science, 7308: 75–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30598-6_6/.

Martinez:2012:HSR

[157] Patricia López Martínez and Tullio Vardanega. Handling synchronization requirements under separation of concerns in model-driven componentbased development. Lecture Notes in Computer Science, 7308:89–104, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30598-6_7/.

Faria:2012:AMC

[158] José Miguel Faria, João Martins, and Jorge Sousa Pinto. An approach to model checking Ada programs. Lecture Notes in Computer Science, 7308: 105–118, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30598-6_8/.

Edmunds:2012:FMA

[159] Andrew Edmunds, Abdolbaghi Rezazadeh, and Michael Butler. Formal modelling for Ada implementations: Tasking event-B. Lecture Notes in Computer Science, 7308:119–132, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30598-6_9/.

Iliasov:2012:AFD

[160] Alexei Iliasov. Augmenting formal development with use case reasoning. Lecture Notes in Computer Science, 7308:133–146, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30598-6_10/.

Pereverzeva:2012:FGO

[161] Inna Pereverzeva, Elena Troubitsyna, and Linas Laibinis. Formal goaloriented development of resilient MAS in event-B. Lecture Notes in Computer Science, 7308:147–161, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30598-6_11/.

Poon:2012:CCC

[162] Pak-Lok Poon, Tsong Yueh Chen, and T. H. Tse. Choices, choices: Comparing between CHOC’LATE and the classification-tree methodology. Lecture Notes in Computer Science, 7308:162–176, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30598-6_12/.

Gregertsen:2012:IPE

[163] Kristoffer Nyborg Gregertsen and Amund Skavhaug. Improving the performance of execution time control by using a hardware time management unit. Lecture Notes in Computer Science, 7308:177–192, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30598-6_13/.

Fairbairn:2012:IVE

[164] Mark Louis Fairbairn and Alan Burns. Implementing and verifying EDF preemption-level resource control. Lecture Notes in Computer Science, 7308:193–206, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-30598-6_14/.

Polzlbauer:2012:ECH

[165] Florian Pölzlbauer, Iain Bate, and Eugen Brenner. Efficient constraint handling during designing reliable automotive real-time systems. Lecture Notes in Computer Science, 7308:207–220, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30598-6_15/.

Anonymous:2012:BMf

[166] Anonymous. Back matter. Lecture Notes in Computer Science, 7308: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-30598-6/1.

Anonymous:2012:FMi

[167] Anonymous. Front matter. Lecture Notes in Computer Science, 7308: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-30598-6/1.

Mazo:2012:FLI

[168] Loïc Mazo. A framework for label images. Lecture Notes in Computer Science, 7309:1–10, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30238-1_1/.

Ayala:2012:PDM

[169] Rafael Ayala and Desamparados Fernández-Ternero. Perfect discrete Morse functions on triangulated 3-manifolds. Lecture Notes in Computer Science, 7309:11–19, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30238-1_2/.

Damiand:2012:ROD

[170] Guillaume Damiand and Rocio Gonzalez-Diaz. Removal operations in n D generalized maps for efficient homology computation. Lecture Notes in Computer Science, 7309:20–29, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30238-1_3/.

Vucini:2012:ERN

[171] Erald Vuçini. Enhancing the reconstruction from non-uniform point sets using persistence information. Lecture Notes in Computer Science, 7309: 30–38, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30238-1_4/.

Pena-Cantillana:2012:PSD

[172] Francisco Peña-Cantillana and Ainhoa Berciano. Parallel skeletonizing of digital images by using cellular automata. Lecture Notes in Computer Science, 7309:39–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30238-1_5/.

Heras:2012:TCC

[173] Jónathan Heras, Maxime Dénès, and Gadea Mata. Towards a certified computation of homology groups for digital images. Lecture Notes in Computer Science, 7309:49–57, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30238-1_6/.

Pacheco:2012:EAC

[174] Ana Pacheco and Pedro Real. An efficient algorithm to compute subsets of points in Z^{n}. Lecture Notes in Computer Science, 7309:58–67, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30238-1_7/.

Wagner:2012:CTT

[175] Hubert Wagner, Paweł Dłotko, and Marian Mrozek. Computational topology in text mining. Lecture Notes in Computer Science, 7309: 68–78, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30238-1_8/.

Mesmoudi:2012:CCM

[176] Mohammed Mostefa Mesmoudi and Leila De Floriani. Concentrated curvature for mean curvature estimation in triangulated surfaces. Lecture Notes in Computer Science, 7309:79–87, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30238-1_9/.

Escribano:2012:DSP

[177] Carmen Escribano and Antonio Giraldo. Deletion of (26,6)-simple points as multivalued retractions. Lecture Notes in Computer Science, 7309: 88–97, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30238-1_10/.

Comic:2012:TOC

[178] Lidija Čomić and Leila De Floriani. Topological operators on cell complexes in arbitrary dimensions. Lecture Notes in Computer Science, 7309: 98–107, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30238-1_11/.

Carnero:2012:TMC

[179] Javier Carnero, Helena Molina-Abril, and Pedro Real. Triangle mesh compression and homological spanning forests. Lecture Notes in Computer Science, 7309:108–116, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30238-1_12/.

Brendel:2012:HCA

[180] Piotr Brendel, Paweł Dłotko, and Marian Mrozek. Homology computations via acyclic subspace. Lecture Notes in Computer Science, 7309: 117–127, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30238-1_13/.

Cerri:2012:MSA

[181] Andrea Cerri, Barbara Di Fabio, and Filippo Medri. Multi-scale approximation of the matching distance for shape retrieval. Lecture Notes in Computer Science, 7309:128–138, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30238-1_14/.

Gutierrez:2012:PHR

[182] Antonio Gutierrez and David Monaghan. Persistent homology for 3D reconstruction evaluation. Lecture Notes in Computer Science, 7309: 139–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30238-1_15/.

Cagliari:2012:PMS

[183] Francesca Cagliari and Massimo Ferri. Persistence modules, shape description, and completeness. Lecture Notes in Computer Science, 7309: 148–156, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30238-1_16/.

Anonymous:2012:BMg

[184] Anonymous. Back matter. Lecture Notes in Computer Science, 7309: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-30238-1/1.

Anonymous:2012:FMj

[185] Anonymous. Front matter. Lecture Notes in Computer Science, 7309: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-30238-1/1.

Akther:2012:ALS

[186] Aysha Akther, Heung-Nam Kim, Majdi Rawashdeh, and Abdulmotaleb El Saddik. Applying latent semantic analysis to tag-based community recommendations. Lecture Notes in Computer Science, 7310: 1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30353-1_1/.

Alhossaini:2012:MLP

[187] Maher Alhossaini and J. Christopher Beck. Macro learning in planning as parameter configuration. Lecture Notes in Computer Science, 7310: 13–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30353-1_2/.

Bagheri:2012:EPC

[188] Mohammad Ali Bagheri, Qigang Gao, and Sergio Escalera. Efficient pairwise classification using local cross off strategy. Lecture Notes in Computer Science, 7310:25–36, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30353-1_3/.

Bobicev:2012:LST

[189] Victoria Bobicev, Marina Sokolova, Yasser Jafer, and David Schramm. Learning sentiments from tweets with personal health information. Lecture Notes in Computer Science, 7310:37–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30353-1_4/.

Carter:2012:SPQ

[190] Dave Carter and Diana Inkpen. Searching for poor quality machine translated text: Learning the difference between human writing and machine translations. Lecture Notes in Computer Science, 7310:49–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30353-1_5/.

Fournier-Viger:2012:MTK

[191] Philippe Fournier-Viger, Cheng-Wei Wu, and Vincent S. Tseng. Mining top-K association rules. Lecture Notes in Computer Science, 7310: 61–73, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30353-1_6/.

Guo:2012:CSS

[192] Yuanyuan Guo, Harry Zhang, and Bruce Spencer. Cost-sensitive selftraining. Lecture Notes in Computer Science, 7310:74–84, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30353-1_7/.

Heras:2012:ESE

[193] Federico Heras, Antonio Morgado, and Joao Marques-Silva. An empirical study of encodings for group MaxSAT. Lecture Notes in Computer Science, 7310:85–96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30353-1_8/.

Hunter:2012:APL

[194] Aaron Hunter. Actions, preferences, and logic programs. Lecture Notes in Computer Science, 7310:97–108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30353-1_9/.

Juma:2012:PBP

[195] Farah Juma, Eric I. Hsu, and Sheila A. McIlraith. Preference-based planning via MaxSAT. Lecture Notes in Computer Science, 7310:109– 120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30353-1_10/.

Kennedy:2012:GEA

[196] Alistair Kennedy, Anna Kazantseva, Diana Inkpen, and Stan Szpakowicz. Getting emotional about news summarization. Lecture Notes in Computer Science, 7310:121–132, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30353-1_11/.

Mousumi:2012:EPO

[197] Fouzia Mousumi and Kevin Grant. Exploiting the probability of observation for efficient Bayesian network inference. Lecture Notes in Computer Science, 7310:133–144, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30353-1_12/.

Noorian:2012:SRB

[198] Zeinab Noorian, Mahdi Noorian, Michael Fleming, and Stephen Marsh. A strategic reputation-based mechanism for mobile ad hoc networks. Lecture Notes in Computer Science, 7310:145–157, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30353-1_13/.

Sankaran:2012:DAT

[199] Baskaran Sankaran, Majid Razmara, Atefeh Farzindar, Wael Khreich, and Fred Popowich. Domain adaptation techniques for machine translation and their evaluation in a real-world setting. Lecture Notes in Computer Science, 7310:158–169, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30353-1_14/.

Shao:2012:ALA

[200] Hang Shao and Nathalie Japkowicz. Applying least angle regression to ELM. Lecture Notes in Computer Science, 7310:170–180, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30353-1_15/.

Sharma:2012:CBO

[201] Shiven Sharma, Colin Bellinger, and Nathalie Japkowicz. Clustering based one-class classification for compliance verification of the comprehensive nuclear-test-ban treaty. Lecture Notes in Computer Science, 7310: 181–193, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30353-1_16/.

Silver:2012:IMT

[202] Daniel L. Silver and Liangliang Tu. Image morphing: Transfer learning between tasks that have multiple outputs. Lecture Notes in Computer Science, 7310:194–205, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30353-1_17/.

Tan:2012:FSD

[203] Xing Tan. A formal study on the dualities in temporal projection problems. Lecture Notes in Computer Science, 7310:206–217, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30353-1_18/.

Thompson:2012:PGP

[204] Craig D. S. Thompson and Michael C. Horsch. Predicting good propagation methods for constraint satisfaction. Lecture Notes in Computer Science, 7310:218–229, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30353-1_19/.

Wolkowicz:2012:AIF

[205] Jacek Wołkowicz and Vlado Kešelj. Analysis of important factors for measuring similarity of symbolic music using n-gram-based, bag-of-words approach. Lecture Notes in Computer Science, 7310:230–241, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30353-1_20/.

Anonymous:2012:FMk

[206] Anonymous. Front matter. Lecture Notes in Computer Science, 7310: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-30353-1/1.

Bonissone:2012:LML

[207] Piero P. Bonissone. Lazy meta-learning: Creating customized model ensembles on demand. Lecture Notes in Computer Science, 7311:1– 23, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30687-7_1/.

Miikkulainen:2012:MLT

[208] Risto Miikkulainen, Eliana Feasley, and Leif Johnson. Multiagent learning through neuroevolution. Lecture Notes in Computer Science, 7311: 24–46, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30687-7_2/.

Watts:2012:REH

[209] Lloyd Watts. Reverse-engineering the human auditory pathway. Lecture Notes in Computer Science, 7311:47–59, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30687-7_3/.

Yao:2012:UUE

[210] Xin Yao. Unpacking and understanding evolutionary algorithms. Lecture Notes in Computer Science, 7311:60–76, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30687-7_4/.

Ashlock:2012:REC

[211] Daniel Ashlock, Cameron McGuinness, and Wendy Ashlock. Representation in evolutionary computation. Lecture Notes in Computer Science, 7311:77–97, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30687-7_5/.

Michalewicz:2012:QVE

[212] Zbigniew Michalewicz. Quo vadis, evolutionary computation? Lecture Notes in Computer Science, 7311:98–121, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30687-7_6/.

Shim:2012:PGA

[213] Vui Ann Shim and Kay Chen Tan. Probabilistic graphical approaches for learning, modeling, and sampling in evolutionary multiobjective optimization. Lecture Notes in Computer Science, 7311:122– 144, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30687-7_7/.

DeBaets:2012:QTS

[214] Bernard De Baets. The quest for transitivity, a showcase of fuzzy relational calculus. Lecture Notes in Computer Science, 7311:145–165, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30687-7_8/.

Rifqi:2012:CIF

[215] Maria Rifqi. Cognition-inspired fuzzy modelling. Lecture Notes in Computer Science, 7311:166–184, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30687-7_9/.

Tanaka:2012:UFM

[216] Kazuo Tanaka. A unified fuzzy model-based framework for modeling and control of complex systems: From flying vehicle control to brainmachine cooperative control. Lecture Notes in Computer Science, 7311: 185–208, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30687-7_10/.

Cherkassky:2012:PLK

[217] Vladimir Cherkassky. Predictive learning, knowledge discovery and philosophy of science. Lecture Notes in Computer Science, 7311:209–233, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30687-7_11/.

Kasabov:2012:ESN

[218] Nikola Kasabov. Evolving spiking neural networks and neurogenetic systems for spatio- and spectro-temporal data modelling and pattern recognition. Lecture Notes in Computer Science, 7311:234–260, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30687-7_12/.

Yang:2012:UNC

[219] Chenhui Yang, Hongwei Mao, Yuan Yuan, and Bing Cheng. Uncovering the neural code using a rat model during a learning control task. Lecture Notes in Computer Science, 7311:261–279, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30687-7_13/.

Anonymous:2012:BMh

[220] Anonymous. Back matter. Lecture Notes in Computer Science, 7311: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-30687-7/1.

Anonymous:2012:FMl

[221] Anonymous. Front matter. Lecture Notes in Computer Science, 7311: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-30687-7/1.

Mehta:2012:SPE

[222] Kshitij Mehta, Edgar Gabriel, and Barbara Chapman. Specification and performance evaluation of parallel I/O interfaces for OpenMP. Lecture Notes in Computer Science, 7312:1–14, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30961-8_1/.

Eichenberger:2012:DOT

[223] Alexandre E. Eichenberger, Christian Terboven, Michael Wong, and Dieter an Mey. The design of OpenMP thread affinity. Lecture Notes in Computer Science, 7312:15–28, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30961-8_2/.

Royuela:2012:ASO

[224] Sara Royuela, Alejandro Duran, Chunhua Liao, and Daniel J. Quinlan. Auto-scoping for OpenMP tasks. Lecture Notes in Computer Science, 7312:29–43, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30961-8_3/.

Bihari:2012:CIT

[225] Barna L. Bihari, Michael Wong, Amy Wang, Bronis R. de Supinski, and Wang Chen. A case for including transactions in OpenMP II: Hardware transactional memory. Lecture Notes in Computer Science, 7312: 44–58, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30961-8_4/.

Klemm:2012:EOV

[226] Michael Klemm, Alejandro Duran, Xinmin Tian, Hideki Saito, and Diego Caballero. Extending OpenMP* with vector constructs for modern multicore SIMD architectures. Lecture Notes in Computer Science, 7312: 59–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30961-8_5/.

Tahan:2012:ITC

[227] Oussama Tahan, Mats Brorsson, and Mohamed Shawky. Introducing task cancellation to OpenMP. Lecture Notes in Computer Science, 7312: 73–87, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30961-8_6/.

Thoman:2012:AOL

[228] Peter Thoman, Herbert Jordan, Simone Pellegrini, and Thomas Fahringer. Automatic OpenMP loop scheduling: a combined compiler and runtime approach. Lecture Notes in Computer Science, 7312:88–101, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30961-8_7/.

Broquedis:2012:LEO

[229] François Broquedis, Thierry Gautier, and Vincent Danjean. libOMP, an efficient OpenMP runtime system for both fork-join and data flow paradigms. Lecture Notes in Computer Science, 7312:102–115, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30961-8_8/.

Chen:2012:CAR

[230] Li Chen, Baojiang Shou, Xionghui Hou, and Lei Huang. A compilerassisted runtime-prefetching scheme for heterogeneous platforms. Lecture Notes in Computer Science, 7312:116–129, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30961-8_9/.

Meadows:2012:EWI

[231] Larry Meadows. Experiments with WRF on Intel(R) Many Integrated Core (Intel MIC) architecture. Lecture Notes in Computer Science, 7312: 130–139, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30961-8_10/.

Shan:2012:OAA

[232] Hongzhang Shan, Erich Strohmaier, James Amundson, and Eric G. Stern. Optimizing the advanced accelerator simulation framework Synergia using OpenMP. Lecture Notes in Computer Science, 7312:140– 153, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30961-8_11/.

Jin:2012:UCD

[233] Haoqiang Jin, Mark Kellogg, and Piyush Mehrotra. Using compiler directives for accelerating CFD applications on GPUs. Lecture Notes in Computer Science, 7312:154–168, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30961-8_12/.

Sabne:2012:ECO

[234] Amit Sabne, Putt Sakdhnagool, and Rudolf Eigenmann. Effects of compiler optimizations in OpenMP to CUDA translation. Lecture Notes in Computer Science, 7312:169–181, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30961-8_13/.

Terboven:2012:AOT

[235] Christian Terboven, Dirk Schmidl, Tim Cramer, and Dieter an Mey. Assessing OpenMP tasking implementations on NUMA architectures. Lecture Notes in Computer Science, 7312:182–195, 2012. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-30961-8_14/.

Schmidl:2012:PAT

[236] Dirk Schmidl, Peter Philippen, Daniel Lorenz, Christian Rössel, and Markus Geimer. Performance analysis techniques for task-based OpenMP applications. Lecture Notes in Computer Science, 7312:196– 209, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30961-8_15/.

Agathos:2012:TBE

[237] Spiros N. Agathos, Panagiotis E. Hadjidoukas, and Vassilios V. Dimakopoulos. Task-based execution of nested OpenMP loops. Lecture Notes in Computer Science, 7312:210–222, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30961-8_16/.

Muller:2012:SOA

[238] Matthias S. Müller, John Baron, William C. Brantley, Huiyu Feng, and Daniel Hackenberg. SPEC OMP2012— an application benchmark suite for parallel systems using OpenMP. Lecture Notes in Computer Science, 7312:223–236, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-30961-8_17/.

Wang:2012:OVT

[239] Cheng Wang, Sunita Chandrasekaran, and Barbara Chapman. An OpenMP 3.1 validation testsuite. Lecture Notes in Computer Science, 7312:237–249, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-30961-8_18/.

Perla:2012:PAH

[240] Francesca Perla and Paolo Zanetti. Performance analysis of an hybrid MPI/OpenMP ALM software for life insurance policies on multicore architectures. Lecture Notes in Computer Science, 7312:250–253, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30961-8_19/.

Maheo:2012:AOL

[241] Aurèle Mahéo, Souad Koliaï, Patrick Carribault, Marc Pérache, and William Jalby. Adaptive OpenMP for large NUMA nodes. Lecture Notes in Computer Science, 7312:254–257, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30961-8_20/.

Anonymous:2012:FMm

[242] Anonymous. Front matter. Lecture Notes in Computer Science, 7312: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-30961-8/1.

Odersky:2012:WCM

[243] Martin Odersky. When compilers are mirrors. Lecture Notes in Computer Science, 7313:1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/accesspage/
chapter/10.1007/978-3-642-31057-7_1.

Oliveira:2012:EM

[244] Bruno C. d. S. Oliveira and William R. Cook. Extensibility for the masses. Lecture Notes in Computer Science, 7313:2–27, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31057-7_2/.

Robbes:2012:EDS

[245] Romain Robbes, David Röthlisberger, and Éric Tanter. Extensions during software evolution: Do objects meet their promise? Lecture Notes in Computer Science, 7313:28–52, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31057-7_3/.

Reichenbach:2012:PPD

[246] Christoph Reichenbach, Yannis Smaragdakis, and Neil Immerman. PQL: a purely-declarative Java extension for parallel programming. Lecture Notes in Computer Science, 7313:53–78, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31057-7_4/.

Negara:2012:IDU

[247] Stas Negara, Mohsen Vakilian, Nicholas Chen, Ralph E. Johnson, and Danny Dig. Is it dangerous to use version control histories to study source code evolution? Lecture Notes in Computer Science, 7313: 79–103, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31057-7_5/.

Morandat:2012:EDR

[248] Floréal Morandat, Brandon Hill, Leo Osvald, and Jan Vitek. Evaluating the design of the R language. Lecture Notes in Computer Science, 7313: 104–131, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31057-7_6/.

Doherty:2012:MSA

[249] Jesse Doherty and Laurie Hendren. McSAF: a static analysis framework for MATLAB. Lecture Notes in Computer Science, 7313:132–155, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31057-7_7/.

Ostlund:2012:MAE

[250] Johan Östlund and Tobias Wrigstad. Multiple aggregate entry points for ownership types. Lecture Notes in Computer Science, 7313:156– 180, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31057-7_8/.

Huang:2012:ICO

[251] Wei Huang, Werner Dietl, Ana Milanova, and Michael D. Ernst. Inference and checking of object ownership. Lecture Notes in Computer Science, 7313:181–206, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-31057-7_9/.

Zibin:2012:OIX

[252] Yoav Zibin, David Cunningham, Igor Peshansky, and Vijay Saraswat. Object initialization in X10. Lecture Notes in Computer Science, 7313: 207–231, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31057-7_10/.

Wrigstad:2012:SA

[253] Tobias Wrigstad. Structured aliasing. Lecture Notes in Computer Science, 7313:232, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/accesspage/chapter/
10.1007/978-3-642-31057-7_11.

Bierman:2012:PPF

[254] Gavin Bierman, Claudio Russo, Geoffrey Mainland, Erik Meijer, and Mads Torgersen. Pause ’n’ play: Formalizing asynchronous C^{♯}. Lecture Notes in Computer Science, 7313:233–257, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31057-7_12/.

Rytz:2012:LPE

[255] Lukas Rytz, Martin Odersky, and Philipp Haller. Lightweight polymorphic effects. Lecture Notes in Computer Science, 7313:258–282, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31057-7_13/.

Burckhardt:2012:CTE

[256] Sebastian Burckhardt, Manuel Fähndrich, Daan Leijen, and Benjamin P. Wood. Cloud types for eventual consistency. Lecture Notes in Computer Science, 7313:283–307, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31057-7_14/.

Gudka:2012:LIP

[257] Khilan Gudka, Tim Harris, and Susan Eisenbach. Lock inference in the presence of large libraries. Lecture Notes in Computer Science, 7313: 308–332, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31057-7_15/.

Karim:2012:AMJ

[258] Rezwana Karim, Mohan Dhawan, Vinod Ganapathy, and Chung chieh Shan. An analysis of the Mozilla Jetpack extension framework. Lecture Notes in Computer Science, 7313:333–355, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31057-7_16/.

Gil:2012:SFJ

[259] Joseph Gil and Yuval Shimron. Smaller footprint for Java collections. Lecture Notes in Computer Science, 7313:356–382, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31057-7_17/.

Dhawan:2012:EJT

[260] Mohan Dhawan, Chung chieh Shan, and Vinod Ganapathy. Enhancing JavaScript with transactions. Lecture Notes in Computer Science, 7313:

383–408, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31057-7_18/.

Kossakowski:2012:JED

[261] Grzegorz Kossakowski, Nada Amin, Tiark Rompf, and Martin Odersky. JavaScript as an embedded DSL. Lecture Notes in Computer Science, 7313:409–434, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-31057-7_19/.

Sridharan:2012:CTP

[262] Manu Sridharan, Julian Dolby, Satish Chandra, Max Schäfer, and Frank Tip. Correlation tracking for points-to analysis of JavaScript. Lecture Notes in Computer Science, 7313:435–458, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31057-7_20/.

Anonymous:2012:FMn

[263] Anonymous. Front matter. Lecture Notes in Computer Science, 7313: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31057-7/1.

Kim:2012:QSM

[264] Chesoong Kim, Olga Dudina, Alexander Dudin, and Sergey Dudin. Queueing system MAP /M /N as a model of call center with call-back option. Lecture Notes in Computer Science, 7314:1–15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30782-9_1/.

Phung-Duc:2012:TWC

[265] Tuan Phung-Duc and Wouter Rogiest. Two way communication retrial queues with balanced call blending. Lecture Notes in Computer Science, 7314:16–31, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30782-9_2/.

Bruneel:2012:ATC

[266] Herwig Bruneel, Tom Maertens, Bart Steyaert, Dieter Claeys, and Dieter Fiems. Analysis of a two-class FCFS queueing system with interclass correlation. Lecture Notes in Computer Science, 7314:32–46, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30782-9_3/.

Kempa:2012:VWT

[267] Wojciech M. Kempa. The virtual waiting time in a finite-buffer queue with a single vacation policy. Lecture Notes in Computer Science, 7314: 47–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30782-9_4/.

Saffer:2012:APG

[268] Zsolt Saffer, Sergey Andreev, and Yevgeni Koucheryavy. Analysis of periodically gated vacation model and its application to IEEE 802.16 network. Lecture Notes in Computer Science, 7314:61–75, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30782-9_5/.

Aouled:2012:CCF

[269] Idriss-Ismael Aouled and Hind Castel-Taleb. Combined CAC and forced handoff for mobile network performability. Lecture Notes in Computer Science, 7314:76–90, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30782-9_6/.

Giang:2012:MCC

[270] Anh Tuan Giang and Anthony Busson. Modeling CSMA/CA in VANET. Lecture Notes in Computer Science, 7314:91–105, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30782-9_7/.

Gribaudo:2012:CRV

[271] Marco Gribaudo, Pietro Piazzolla, and Giuseppe Serazzi. Consolidation and replication of VMs matching performance objectives. Lecture Notes in Computer Science, 7314:106–120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30782-9_8/.

Bruneel:2012:ADT

[272] Herwig Bruneel, Joris Walraevens, Dieter Claeys, and Sabine Wittevrongel. Analysis of a discrete-time queue with geometrically distributed service capacities. Lecture Notes in Computer Science, 7314:121–135, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30782-9_9/.

Busic:2012:PSN

[273] Ana Bušić, Bruno Gaujal, and Florence Perronnin. Perfect sampling of networks with finite and infinite capacity queues. Lecture Notes in Computer Science, 7314:136–149, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30782-9_10/.

DeCuypere:2012:QTA

[274] Eline De Cuypere, Koen De Turck, and Dieter Fiems. A queueing theoretic approach to decoupling inventory. Lecture Notes in Computer Science, 7314:150–164, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-30782-9_11/.

Tsimashenka:2012:CVS

[275] Iryna Tsimashenka, William Knottenbelt, and Peter Harrison. Controlling variability in split-merge systems. Lecture Notes in Computer Science, 7314:165–177, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-30782-9_12/.

Fourneau:2012:SIC

[276] Jean-Michel Fourneau and Franck Quessette. Some improvements for the computation of the steady-state distribution of a Markov chain by monotone sequences of vectors. Lecture Notes in Computer Science, 7314: 178–192, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30782-9_13/.

Stefanek:2012:MFA

[277] Anton Stefanek, Richard A. Hayden, Mark Mac Gonagle, and Jeremy T. Bradley. Mean-field analysis of Markov models with reward feedback. Lecture Notes in Computer Science, 7314:193–211, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-30782-9_14/.

Balsamo:2012:LRP

[278] Simonetta Balsamo, Gian-Luca Dei Rossi, and Andrea Marin. Lumping and reversed processes in cooperating automata. Lecture Notes in Computer Science, 7314:212–226, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30782-9_15/.

DeTurck:2012:TDS

[279] Koen De Turck, Sofian De Clercq, Sabine Wittevrongel, Herwig Bruneel, and Dieter Fiems. Transform-domain solutions of Poisson’s equation with applications to the asymptotic variance. Lecture Notes in Computer Science, 7314:227–239, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30782-9_16/.

Kaj:2012:ASM

[280] Ingemar Kaj and Victorien Konané. Analytical and stochastic modelling of battery cell dynamics. Lecture Notes in Computer Science, 7314: 240–254, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30782-9_17/.

Altman:2012:BPM

[281] Eitan Altman and Dieter Fiems. Branching processes, the Max–Plus algebra and network calculus. Lecture Notes in Computer Science, 7314: 255–270, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30782-9_18/.

Horvath:2012:EGP

[282] Gábor Horváth, Philipp Reinecke, Miklós Telek, and Katinka Wolter. Efficient generation of PH-distributed random variates. Lecture Notes in Computer Science, 7314:271–285, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30782-9_19/.

Nechval:2012:FPL

[283] Nicholas Nechval, Maris Purgailis, Uldis Rozevskis, Inta Bruna, and Konstantin Nechval. Finding prediction limits for a future number of failures in the prescribed time interval under parametric uncertainty. Lecture Notes in Computer Science, 7314:286–301, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30782-9_20/.

Anonymous:2012:BMi

[284] Anonymous. Back matter. Lecture Notes in Computer Science, 7314: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-30782-9/1.

Anonymous:2012:FMo

[285] Anonymous. Front matter. Lecture Notes in Computer Science, 7314: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-30782-9/1.

Jraidi:2012:ISI

[286] Imène Jraidi, Pierre Chalfoun, and Claude Frasson. Implicit strategies for intelligent tutoring systems. Lecture Notes in Computer Science, 7315:1–10, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30950-2_1/.

Ogan:2012:RRI

[287] Amy Ogan, Samantha Finkelstein, Erin Walker, Ryan Carlson, and Justine Cassell. Rudeness and rapport: Insults and learning gains in peer tutoring. Lecture Notes in Computer Science, 7315:11–21, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30950-2_2/.

Hayashi:2012:PEL

[288] Yugo Hayashi. On pedagogical effects of learner-support agents in collaborative interaction. Lecture Notes in Computer Science, 7315:22– 32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30950-2_3/.

Zhang:2012:EAD

[289] Li Zhang. Exploration of affect detection using semantic cues in virtual improvisation. Lecture Notes in Computer Science, 7315:33–39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30950-2_4/.

Harley:2012:MLC

[290] Jason M. Harley, François Bouchet, and Roger Azevedo. Measuring learners’ co-occurring emotional responses during their interaction with a pedagogical agent in MetaTutor. Lecture Notes in Computer Science, 7315:40–45, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30950-2_5/.

Shanabrook:2012:VSA

[291] David Hilton Shanabrook, Ivon Arroyo, Beverly Park Woolf, and Winslow Burleson. Visualization of student activity patterns within intelligent tutoring systems. Lecture Notes in Computer Science, 7315:46–51, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30950-2_6/.

Grafsgaard:2012:TML

[292] Joseph F. Grafsgaard, Kristy Elizabeth Boyer, and James C. Lester. Toward a machine learning framework for understanding affective tutorial interaction. Lecture Notes in Computer Science, 7315:52–58, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30950-2_7/.

Strain:2012:ERB

[293] Amber Chauncey Strain, Roger Azevedo, and Sidney D’Mello. Exploring relationships between learners’ affective states, metacognitive processes, and learning outcomes. Lecture Notes in Computer Science, 7315: 59–64, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30950-2_8/.

Chaouachi:2012:MWE

[294] Maher Chaouachi and Claude Frasson. Mental workload, engagement and emotions: An exploratory study for intelligent tutoring systems. Lecture Notes in Computer Science, 7315:65–71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30950-2_9/.

Brawner:2012:RTM

[295] Keith W. Brawner and Benjamin S. Goldberg. Real-time monitoring of ECG and GSR signals during computer-based training. Lecture Notes in Computer Science, 7315:72–77, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30950-2_10/.

Hussain:2012:CVD

[296] Md. Sazzad Hussain, Hamed Monkaresi, and Rafael A. Calvo. Categorical vs. dimensional representations in multimodal affect detection during learning. Lecture Notes in Computer Science, 7315:78–83, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30950-2_11/.

Chalfoun:2012:CPA

[297] Pierre Chalfoun and Claude Frasson. Cognitive priming: Assessing the use of non-conscious perception to enhance learner’s reasoning ability. Lecture Notes in Computer Science, 7315:84–89, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30950-2_12/.

Rai:2012:MLE

[298] Dovan Rai and Joseph E. Beck. Math learning environment with gamelike elements: An incremental approach for enhancing student engagement and learning effectiveness. Lecture Notes in Computer Science, 7315: 90–100, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30950-2_13/.

Matsuda:2012:MFL

[299] Noboru Matsuda, Evelyn Yarzebinski, Victoria Keiser, Rohan Raizada, and Gabriel Stylianides. Motivational factors for learning by teaching. Lecture Notes in Computer Science, 7315:101–111, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30950-2_14/.

Muir:2012:AAS

[300] Mary Muir and Cristina Conati. An analysis of attention to studentadaptive hints in an educational game. Lecture Notes in Computer Science, 7315:112–122, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30950-2_15/.

Muratet:2012:SGS

[301] Mathieu Muratet, Elisabeth Delozanne, Patrice Torguet, and Fabienne Viallet. Serious game and students’ learning motivation: Effect of context using Prog&Play. Lecture Notes in Computer Science, 7315: 123–128, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30950-2_16/.

Derbali:2012:EEP

[302] Lotfi Derbali and Claude Frasson. Exploring the effects of prior videogame experience on learner’s motivation during interactions with Heap-Motiv. Lecture Notes in Computer Science, 7315:129–134, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30950-2_17/.

Marne:2012:DPL

[303] Bertrand Marne, John Wisdom, Benjamin Huynh-Kim-Bang, and Jean-Marc Labat. A design pattern library for mutual understanding and cooperation in serious game design. Lecture Notes in Computer Science, 7315: 135–140, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30950-2_18/.

Sabourin:2012:PSS

[304] Jennifer Sabourin, Lucy R. Shores, Bradford W. Mott, and James C. Lester. Predicting student self-regulation strategies in game-based learning environments. Lecture Notes in Computer Science, 7315:141–150, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30950-2_19/.

Wang:2012:TAV

[305] Ning Wang, David V. Pynadath, and Stacy C. Marsella. Toward automatic verification of multiagent systems for training simulations. Lecture Notes in Computer Science, 7315:151–161, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30950-2_20/.

Anonymous:2012:FMp

[306] Anonymous. Front matter. Lecture Notes in Computer Science, 7315: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-30950-2/1.

Borger:2012:CRAa

[307] Egon Börger, Antonio Cisternino, and Vincenzo Gervasi. Contribution to a rigorous analysis of Web application frameworks. Lecture Notes in Computer Science, 7316:1–20, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30885-7_1/.

Hayes:2012:IOS

[308] Ian J. Hayes and Robert J. Colvin. Integrated operational semantics: Small-step, big-step and multi-step. Lecture Notes in Computer Science, 7316:21–35, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30885-7_2/.

Arcaini:2012:TGS

[309] Paolo Arcaini, Francesco Bolis, and Angelo Gargantini. Test generation for sequential nets of abstract state machines. Lecture Notes in Computer Science, 7316:36–50, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30885-7_3/.

Banach:2012:ACS

[310] Richard Banach, Huibiao Zhu, Wen Su, and Xiaofeng Wu. ASM and controller synthesis. Lecture Notes in Computer Science, 7316:51–64, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30885-7_4/.

Banach:2012:CAP

[311] Richard Banach, Huibiao Zhu, Wen Su, and Xiaofeng Wu. Continuous ASM, and a pacemaker sensing fragment. Lecture Notes in Computer Science, 7316:65–78, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30885-7_5/.

Gervasi:2012:AMC

[312] Vincenzo Gervasi. An ASM model of concurrency in a Web browser. Lecture Notes in Computer Science, 7316:79–93, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-30885-7_6/.

Fraikin:2012:MSC

[313] Benoît Fraikin, Marc Frappier, and Richard St-Denis. Modeling the supervisory control theory with Alloy. Lecture Notes in Computer Science, 7316:94–107, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30885-7_7/.

Milicevic:2012:PAO

[314] Aleksandar Milicevic and Daniel Jackson. Preventing arithmetic overflows in Alloy. Lecture Notes in Computer Science, 7316:108–121, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30885-7_8/.

Montaghami:2012:EAP

[315] Vajih Montaghami and Derek Rayside. Extending Alloy with partial instances. Lecture Notes in Computer Science, 7316:122–135, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30885-7_9/.

Nelson:2012:TMC

[316] Timothy Nelson, Daniel J. Dougherty, Kathi Fisler, and Shriram Krishnamurthi. Toward a more complete Alloy. Lecture Notes in Computer Science, 7316:136–149, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30885-7_10/.

Vakili:2012:TLM

[317] Amirhossein Vakili and Nancy A. Day. Temporal logic model checking in Alloy. Lecture Notes in Computer Science, 7316:150–163, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30885-7_11/.

Wang:2012:AAM

[318] Ting Wang and Dongyao Ji. Active attacking multicast key management protocol using Alloy. Lecture Notes in Computer Science, 7316: 164–177, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30885-7_12/.

Abrial:2012:FHS

[319] Jean-Raymond Abrial, Wen Su, and Huibiao Zhu. Formalizing hybrid systems with event-B. Lecture Notes in Computer Science, 7316: 178–193, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30885-7_13/.

Deharbe:2012:SSR

[320] David Déharbe, Pascal Fontaine, Yoann Guyot, and Laurent Voisin. SMT solvers for Rodin. Lecture Notes in Computer Science, 7316: 194–207, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30885-7_14/.

Grov:2012:RPI

[321] Gudmund Grov, Andrew Ireland, and Maria Teresa Llano. Refinement plans for informed formal design. Lecture Notes in Computer Science, 7316:208–222, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-30885-7_15/.

Hallerstede:2012:RII

[322] Stefan Hallerstede and Thai Son Hoang. Refinement by interface instantiation. Lecture Notes in Computer Science, 7316:223–237, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30885-7_16/.

Mentre:2012:DPO

[323] David Mentré, Claude Marché, Jean-Christophe Filliâtre, and Masashi Asuka. Discharging proof obligations from Atelier B using multiple automated provers. Lecture Notes in Computer Science, 7316:238–251, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30885-7_17/.

Jones:2012:SAL

[324] Cliff B. Jones, Matthew J. Lovert, and L. Jason Steggles. A semantic analysis of logics that cope with partial terms. Lecture Notes in Computer Science, 7316:252–265, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30885-7_18/.

Nielsen:2012:CVE

[325] Claus Ballegaard Nielsen, Kenneth Lausdahl, and Peter Gorm Larsen. Combining VDM with executable code. Lecture Notes in Computer Science, 7316:266–279, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30885-7_19/.

Cristia:2012:ETT

[326] Maximiliano Cristiá and Claudia Frydman. Extending the test template framework to deal with axiomatic descriptions, quantifiers and set comprehensions. Lecture Notes in Computer Science, 7316:280–293, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30885-7_20/.

Anonymous:2012:FMq

[327] Anonymous. Front matter. Lecture Notes in Computer Science, 7316: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-30885-7/1.

Bradley:2012:UI

[328] Aaron R. Bradley. Understanding IC3. Lecture Notes in Computer Science, 7317:1–14, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-31612-8_1/.

Knuth:2012:SAC

[329] Donald Knuth. Satisfiability and the art of computer programming. Lecture Notes in Computer Science, 7317:15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/accesspage/chapter/10.1007/978-3-642-31612-8_2.

Balint:2012:CPD

[330] Adrian Balint and Uwe Schöning. Choosing probability distributions for stochastic local search and the role of make versus break. Lecture Notes in Computer Science, 7317:16–29, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31612-8_3/.

Goultiaeva:2012:TRE

[331] Alexandra Goultiaeva and Fahiem Bacchus. Off the trail: Re-examining the CDCL algorithm. Lecture Notes in Computer Science, 7317:30– 43, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31612-8_4/.

Bonet:2012:ISR

[332] Maria Luisa Bonet and Sam Buss. An improved separation of regular resolution from pool resolution and clause learning. Lecture Notes in Computer Science, 7317:44–57, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31612-8_5/.

Slivovsky:2012:CRP

[333] Friedrich Slivovsky and Stefan Szeider. Computing resolution-path dependencies in linear time. Lecture Notes in Computer Science, 7317: 58–71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31612-8_6/.

Gaspers:2012:SBN

[334] Serge Gaspers and Stefan Szeider. Strong backdoors to nested satisfiability. Lecture Notes in Computer Science, 7317:72–85, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31612-8_7/.

VanGelder:2012:EFL

[335] Allen Van Gelder, Samuel B. Wood, and Florian Lonsing. Extended failedliteral preprocessing for quantified Boolean formulas. Lecture Notes in Computer Science, 7317:86–99, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31612-8_8/.

Egly:2012:SSR

[336] Uwe Egly. On sequent systems and resolution for QBFs. Lecture Notes in Computer Science, 7317:100–113, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31612-8_9/.

Janota:2012:SQC

[337] Mikoláš Janota, William Klieber, Joao Marques-Silva, and Edmund Clarke. Solving QBF with counterexample guided refinement. Lecture Notes in Computer Science, 7317:114–128, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31612-8_10/.

Balabanov:2012:HQB

[338] Valeriy Balabanov, Hui-Ju Katherine Chiang, and Jie-Hong Roland Jiang. Henkin quantifiers and Boolean formulae. Lecture Notes in Computer Science, 7317:129–142, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31612-8_11/.

Ganesh:2012:LPS

[339] Vijay Ganesh, Charles W. O’Donnell, Mate Soos, Srinivas Devadas, and Martin C. Rinard. Lynx: a programmatic SAT solver for the RNAfolding problem. Lecture Notes in Computer Science, 7317:143–156, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31612-8_12/.

Hoder:2012:GPD

[340] Kryštof Hoder and Nikolaj Bjørner. Generalized property directed reachability. Lecture Notes in Computer Science, 7317:157–171, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31612-8_13/.

Ermon:2012:SAC

[341] Stefano Ermon, Ronan Le Bras, Carla P. Gomes, Bart Selman, and R. Bruce van Dover. SMT-aided combinatorial materials discovery. Lecture Notes in Computer Science, 7317:172–185, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31612-8_14/.

Zhang:2012:FII

[342] Jian Zhang, Feifei Ma, and Zhiqiang Zhang. Faulty interaction identification via constraint solving and optimization. Lecture Notes in Computer Science, 7317:186–199, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31612-8_15/.

Audemard:2012:RCE

[343] Gilles Audemard, Benoît Hoessen, Saïd Jabbour, Jean-Marie Lagniez, and Cédric Piette. Revisiting clause exchange in parallel SAT solving. Lecture Notes in Computer Science, 7317:200–213, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31612-8_16/.

Hyvarinen:2012:DSP

[344] Antti E. J. Hyvärinen and Norbert Manthey. Designing scalable parallel SAT solvers. Lecture Notes in Computer Science, 7317:214–227, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31612-8_17/.

Xu:2012:ECS

[345] Lin Xu, Frank Hutter, Holger Hoos, and Kevin Leyton-Brown. Evaluating component solver contributions to portfolio-based algorithm selectors. Lecture Notes in Computer Science, 7317:228–241, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31612-8_18/.

Nadel:2012:ESS

[346] Alexander Nadel and Vadim Ryvchin. Efficient SAT solving under assumptions. Lecture Notes in Computer Science, 7317:242–255, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31612-8_19/.

Nadel:2012:PIS

[347] Alexander Nadel, Vadim Ryvchin, and Ofer Strichman. Preprocessing in incremental SAT. Lecture Notes in Computer Science, 7317: 256–269, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31612-8_20/.

Anonymous:2012:FMr

[348] Anonymous. Front matter. Lecture Notes in Computer Science, 7317: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31612-8/1.

Li:2012:PDB

[349] Kun Li, Man Lu, Fenglong Lu, Qin Lv, Li Shang, and Dragan Maksimovic. Personalized driving behavior monitoring and analysis for emerging hybrid vehicles. Lecture Notes in Computer Science, 7319:1–19, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31205-2_1/.

Maekawa:2012:MSB

[350] Takuya Maekawa, Yasue Kishino, Yutaka Yanagisawa, and Yasushi Sakurai. Mimic sensors: Battery-shaped sensor node for detecting electrical events of handheld devices. Lecture Notes in Computer Science, 7319: 20–38, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31205-2_2/.

Hwang:2012:LCB

[351] Inseok Hwang, Hyukjae Jang, Taiwoo Park, Aram Choi, Youngki Lee, and Chanyou Hwang. Leveraging children’s behavioral distribution and singularities in new interactive environments: Study in kindergarten field trips. Lecture Notes in Computer Science, 7319:39–56, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31205-2_3/.

Castro:2012:UTM

[352] Pablo Samuel Castro, Daqing Zhang, and Shijian Li. Urban traffic modelling and prediction using large scale taxi GPS traces. Lecture Notes in Computer Science, 7319:57–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31205-2_4/.

Tominaga:2012:UFM

[353] Shoji Tominaga, Masamichi Shimosaka, Rui Fukui, and Tomomasa Sato. A unified framework for modeling and predicting going-out behavior. Lecture Notes in Computer Science, 7319:73–90, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31205-2_5/.

Lathia:2012:HIC

[354] Neal Lathia, Daniele Quercia, and Jon Crowcroft. The hidden image of the city: Sensing community well-being from urban mobility. Lecture Notes in Computer Science, 7319:91–98, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31205-2_6/.

Gornerup:2012:SMC

[355] Olof Görnerup. Scalable mining of common routes in mobile communication network traffic data. Lecture Notes in Computer Science, 7319: 99–106, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31205-2_7/.

Bates:2012:AER

[356] Oliver Bates, Adrian K. Clear, Adrian Friday, Mike Hazas, and Janine Morley. Accounting for energy-reliant services within everyday life at home. Lecture Notes in Computer Science, 7319:107–124, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31205-2_8/.

Lu:2012:SBA

[357] Jiakang Lu and Kamin Whitehouse. Smart blueprints: Automatically generated maps of homes and the devices within them. Lecture Notes in Computer Science, 7319:125–142, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31205-2_9/.

Mennicken:2012:HNH

[358] Sarah Mennicken and Elaine M. Huang. Hacking the natural habitat: An in-the-wild study of smart homes, their development, and the people who live in them. Lecture Notes in Computer Science, 7319: 143–160, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31205-2_10/.

Li:2012:DSA

[359] Ming Li, Lars Mahnkopf, and Leif Kobbelt. The design of a Segway AR-tactile navigation system. Lecture Notes in Computer Science, 7319: 161–178, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31205-2_11/.

Kim:2012:RGM

[360] SeungJun Kim, Jin-Hyuk Hong, Kevin A. Li, Jodi Forlizzi, and Anind K. Dey. Route guidance modality for elder driver navigation. Lecture Notes in Computer Science, 7319:179–196, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31205-2_12/.

Molyneaux:2012:IEA

[361] David Molyneaux, Shahram Izadi, David Kim, Otmar Hilliges, Steve Hodges, and Xiang Cao. Interactive environment-aware handheld projectors for pervasive computing spaces. Lecture Notes in Computer Science, 7319:197–215, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31205-2_13/.

Villar:2012:NGP

[362] Nicolas Villar, James Scott, Steve Hodges, Kerry Hammil, and Colin Miller. .NET gadgeteer: a platform for custom devices. Lecture Notes in Computer Science, 7319:216–233, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31205-2_14/.

Maekawa:2012:RHE

[363] Takuya Maekawa, Yasue Kishino, Yutaka Yanagisawa, and Yasushi Sakurai. Recognizing handheld electrical device usage with hand-worn coil of wire. Lecture Notes in Computer Science, 7319:234–252, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31205-2_15/.

Hameed:2012:SCR

[364] Bilal Hameed, Farhan Rashid, Frank Dürr, and Kurt Rothermel. Selfcalibration of RFID reader probabilities in a smart real-time factory. Lecture Notes in Computer Science, 7319:253–270, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31205-2_16/.

Pulkkinen:2012:AAD

[365] Teemu Pulkkinen and Petteri Nurmi. AWESOM: Automatic discrete partitioning of indoor spaces for WiFi fingerprinting. Lecture Notes in Computer Science, 7319:271–288, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31205-2_17/.

Schougaard:2012:IPN

[366] Kari Rye Schougaard, Kaj Grønbæk, and Tejs Scharling. Indoor pedestrian navigation based on hybrid route planning and location modeling. Lecture Notes in Computer Science, 7319:289–306, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31205-2_18/.

Kamisaka:2012:EPR

[367] Daisuke Kamisaka, Takafumi Watanabe, Shigeki Muramatsu, and Arei Kobayashi. Estimating position relation between two pedestrians using mobile phones. Lecture Notes in Computer Science, 7319:307–324, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31205-2_19/.

Higuchi:2012:CCC

[368] Takamasa Higuchi, Hirozumi Yamaguchi, and Teruo Higashino. Clearing a crowd: Context-supported neighbor positioning for people-centric navigation. Lecture Notes in Computer Science, 7319:325–342, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31205-2_20/.

Anonymous:2012:FMs

[369] Anonymous. Front matter. Lecture Notes in Computer Science, 7319: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31205-2/1.

Selic:2012:LWK

[370] Bran Selic. The less Well known UML. Lecture Notes in Computer Science, 7320:1–20, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30982-3_1/.

Andova:2012:MBD

[371] Suzana Andova and Mark G. J. van den Brand. MDE basics with a DSL focus. Lecture Notes in Computer Science, 7320:21–57, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30982-3_2/.

Cabot:2012:OCL

[372] Jordi Cabot and Martin Gogolla. Object constraint language (OCL): a definitive guide. Lecture Notes in Computer Science, 7320:58–90, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30982-3_3/.

DiRuscio:2012:MT

[373] Davide Di Ruscio and Romina Eramo. Model transformations. Lecture Notes in Computer Science, 7320:91–136, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30982-3_4/.

Giese:2012:GTM

[374] Holger Giese, Leen Lambers, and Basil Becker. Graph transformations for MDE, adaptation, and models at runtime. Lecture Notes in Computer Science, 7320:137–191, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30982-3_5/.

deCaso:2012:AVA

[375] Guido de Caso and Victor Braberman. Abstractions for validation in action. Lecture Notes in Computer Science, 7320:192–218, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30982-3_6/.

Petriu:2012:SPM

[376] Dorina C. Petriu and Mohammad Alhaj. Software performance modeling. Lecture Notes in Computer Science, 7320:219–262, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30982-3_7/.

Becker:2012:MTN

[377] Steffen Becker. Model transformations in non-functional analysis. Lecture Notes in Computer Science, 7320:263–289, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30982-3_8/.

Cortellessa:2012:SPA

[378] Vittorio Cortellessa and Antinisca Di Marco. Software performance antipatterns: Modeling and analysis. Lecture Notes in Computer Science, 7320:290–335, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-30982-3_9/.

Brosch:2012:IMV

[379] Petra Brosch, Gerti Kappel, and Philip Langer. An introduction to model versioning. Lecture Notes in Computer Science, 7320:336–398, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30982-3_10/.

Vallecillo:2012:FST

[380] Antonio Vallecillo and Martin Gogolla. Formal specification and testing of model transformations. Lecture Notes in Computer Science, 7320: 399–437, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30982-3_11/.

Anonymous:2012:BMj

[381] Anonymous. Back matter. Lecture Notes in Computer Science, 7320: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-30982-3/1.

Anonymous:2012:FMt

[382] Anonymous. Front matter. Lecture Notes in Computer Science, 7320: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-30982-3/1.

Borger:2012:CRAb

[383] Egon Börger, Antonio Cisternino, and Vincenzo Gervasi. Contribution to a rigorous analysis of Web application frameworks. Lecture Notes in Computer Science, 7321:1–20, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30729-4_1/.

Calder:2012:PAE

[384] Muffy Calder and Michele Sevegnani. Process algebra for event-driven runtime verification: a case study of wireless network management. Lecture Notes in Computer Science, 7321:21–23, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30729-4_2/.

Hansen:2012:TTB

[385] Dominik Hansen and Michael Leuschel. Translating TLA^{+} to B for validation with ProB. Lecture Notes in Computer Science, 7321:24– 38, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30729-4_3/.

Dongol:2012:RGR

[386] Brijesh Dongol and Ian J. Hayes. Rely/guarantee reasoning for teleoreactive programs over multiple time bands. Lecture Notes in Computer Science, 7321:39–53, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30729-4_4/.

Isobe:2012:SLC

[387] Yoshinao Isobe, Faron Moller, and Hoang Nga Nguyen. Safety and line capacity in railways — an approach in timed CSP. Lecture Notes in Computer Science, 7321:54–68, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30729-4_5/.

Berthing:2012:RBD

[388] Jesper Berthing, Pontus Boström, and Kaisa Sere. Refinement-based development of timed systems. Lecture Notes in Computer Science, 7321: 69–83, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30729-4_6/.

Blackmore:2012:ACS

[389] Tim Blackmore, David Halliwell, Philip Barker, and Kerstin Eder. Analysing and closing simulation coverage by automatic generation and verification of formal properties from coverage reports. Lecture Notes in Computer Science, 7321:84–98, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30729-4_7/.

Zhang:2012:MCS

[390] Fuyuan Zhang, Flemming Nielson, and Hanne Riis Nielson. Model checking as static analysis: Revisited. Lecture Notes in Computer Science, 7321:99–112, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30729-4_8/.

Ngo:2012:FVC

[391] Van Chan Ngo, Jean-Pierre Talpin, and Thierry Gautier. Formal verification of compiler transformations on polychronous equations. Lecture Notes in Computer Science, 7321:113–127, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30729-4_9/.

Rocha:2012:UPB

[392] Herbert Rocha, Raimundo Barreto, and Lucas Cordeiro. Understanding programming bugs in ANSI-C software using bounded model checking counter-examples. Lecture Notes in Computer Science, 7321:128– 142, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30729-4_10/.

Kazemeyni:2012:MBW

[393] Fatemeh Kazemeyni, Einar Broch Johnsen, and Olaf Owe. MULE-based wireless sensor networks: Probabilistic modeling and quantitative analysis. Lecture Notes in Computer Science, 7321:143–157, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30729-4_11/.

Woehrle:2012:MET

[394] Matthias Woehrle, Rena Bakhshi, and Mohammad Reza Mousavi. Mechanized extraction of topology anti-patterns in wireless networks. Lecture Notes in Computer Science, 7321:158–173, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30729-4_12/.

Lensink:2012:PFC

[395] Leonard Lensink, Sjaak Smetsers, and Marko van Eekelen. A proof framework for concurrent programs. Lecture Notes in Computer Science, 7321: 174–190, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30729-4_13/.

Bresciani:2012:USP

[396] Riccardo Bresciani and Andrew Butterfield. A UTP semantics of pGCL as a homogeneous relation. Lecture Notes in Computer Science, 7321:

191–205, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30729-4_14/.

Tian:2012:BBC

[397] HaiYun Tian, Phillip J. Brooke, and Anne-Gwenn Bosser. Behaviourbased cheat detection in multiplayer games with event-B. Lecture Notes in Computer Science, 7321:206–220, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30729-4_15/.

Diaconescu:2012:RPT

[398] Denisa Diaconescu, Ioana Leustean, Luigia Petre, and Kaisa Sere. Refinement-preserving translation from event-B to register-voice interactive systems. Lecture Notes in Computer Science, 7321:221–236, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30729-4_16/.

Tarasyuk:2012:FMV

[399] Anton Tarasyuk, Elena Troubitsyna, and Linas Laibinis. Formal modelling and verification of service-oriented systems in probabilistic event-B. Lecture Notes in Computer Science, 7321:237–252, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30729-4_17/.

Markovski:2012:PSP

[400] Jasen Markovski, Dirk A. van Beek, and Jos Baeten. Partially-supervised plants: Embedding control requirements in plant components. Lecture Notes in Computer Science, 7321:253–267, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30729-4_18/.

Hooman:2012:EFD

[401] Jozef Hooman, Arjan J. Mooij, and Hans van Wezep. Early fault detection in industry using models at various abstraction levels. Lecture Notes in Computer Science, 7321:268–282, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30729-4_19/.

Ji:2012:PKP

[402] Ran Ji and Richard Bubel. PE-KeY: a partial evaluator for Java programs. Lecture Notes in Computer Science, 7321:283–295, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30729-4_20/.

Anonymous:2012:FMu

[403] Anonymous. Front matter. Lecture Notes in Computer Science, 7321: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-30729-4/1.

Agten:2012:RDL

[404] Pieter Agten, Nick Nikiforakis, Raoul Strackx, Willem De Groef, and Frank Piessens. Recent developments in low-level software security. Lecture Notes in Computer Science, 7322:1–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-30955-7_1/.

Kearney:2012:TCP

[405] Paul Kearney. Towards a C^{2}I platform for combating the cyber-threat. Lecture Notes in Computer Science, 7322:17–19, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-30955-7_2/.

Gollmann:2012:VPR

[406] Dieter Gollmann. Veracity, plausibility, and reputation. Lecture Notes in Computer Science, 7322:20–28, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30955-7_3/.

Hernandez-Castro:2012:AFH

[407] Julio Cesar Hernandez-Castro, Pedro Peris-Lopez, Masoumeh Safkhani, and Nasour Bagheri. Another fallen hash-based RFID authentication protocol. Lecture Notes in Computer Science, 7322:29–37, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30955-7_4/.

Venkatesh:2012:HBD

[408] G. Kirubavathi Venkatesh and R. Anitha Nadarajan. HTTP botnet detection using adaptive learning rate multilayer feed-forward neural network. Lecture Notes in Computer Science, 7322:38–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30955-7_5/.

Liu:2012:HBE

[409] Fanbao Liu and Tao Xie. How to break EAP-MD5. Lecture Notes in Computer Science, 7322:49–57, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30955-7_6/.

Zhou:2012:PPS

[410] Jian Zhou, Jiwu Jing, Ji Xiang, and Lei Wang. Privacy preserving social network publication on bipartite graphs. Lecture Notes in Computer Science, 7322:58–70, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30955-7_7/.

Christin:2012:PBU

[411] Delphine Christin, Pablo Sánchez López, Andreas Reinhardt, and Matthias Hollick. Privacy bubbles: User-centered privacy control for mobile content sharing applications. Lecture Notes in Computer Science, 7322:71–86, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30955-7_8/.

Kiyomoto:2012:PPU

[412] Shinsaku Kiyomoto, Kazuhide Fukushima, and Yutaka Miyake. Privacy preservation of user history graph. Lecture Notes in Computer Science, 7322:87–96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30955-7_9/.

DellAmico:2012:HSP

[413] Matteo Dell’Amico, Gabriel Serme, Muhammad Sabir Idrees, and Anderson Santana de Olivera. HiPoLDS: a security policy language for distributed systems. Lecture Notes in Computer Science, 7322:97–112, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30955-7_10/.

Nassr:2012:RRO

[414] Nezar Nassr and Eric Steegmans. ROAC: a role-oriented access control model. Lecture Notes in Computer Science, 7322:113–127, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30955-7_11/.

Ganesh:2012:OPE

[415] Chaya Ganesh and C. Pandu Rangan. Optimal parameters for efficient two-party computation protocols. Lecture Notes in Computer Science, 7322:128–143, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-30955-7_12/.

Bohli:2012:ASS

[416] Jens-Matthias Bohli, Wenting Li, and Jan Seedorf. Assisting server for secure multi-party computation. Lecture Notes in Computer Science, 7322: 144–159, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30955-7_13/.

ElBansarkhani:2012:ELB

[417] Rachid El Bansarkhani and Mohammed Meziani. An efficient lattice-based secret sharing construction. Lecture Notes in Computer Science, 7322: 160–168, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30955-7_14/.

Souissi:2012:OCP

[418] Youssef Souissi, Nicolas Debande, Sami Mekki, Sylvain Guilley, and Ali Maalaoui. On the optimality of correlation power attack on embedded cryptographic systems. Lecture Notes in Computer Science, 7322: 169–178, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30955-7_15/.

Karakoc:2012:IDC

[419] Ferhat Karakoç, Hüseyin Demirci, and A. Emre Harmancı. Impossible differential cryptanalysis of reduced-round LBlock. Lecture Notes in Computer Science, 7322:179–188, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30955-7_16/.

Grossschadl:2012:EJI

[420] Johann Großschädl, Dan Page, and Stefan Tillich. Efficient Java implementation of elliptic curve cryptography for J2ME-enabled mobile devices. Lecture Notes in Computer Science, 7322:189–207, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30955-7_17/.

Schreckling:2012:KRT

[421] Daniel Schreckling, Joachim Posegga, Johannes Köstler, and Matthias Schaff. Kynoid: Real-time enforcement of fine-grained, user-defined, and data-centric security policies for Android. Lecture Notes in Computer Science, 7322:208–223, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30955-7_18/.

Anonymous:2012:BMk

[422] Anonymous. Back matter. Lecture Notes in Computer Science, 7322: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-30955-7/1.

Anonymous:2012:FMv

[423] Anonymous. Front matter. Lecture Notes in Computer Science, 7322: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-30955-7/1.

Chung:2012:HCG

[424] Fan Chung and Alexander Tsiatas. Hypergraph coloring games and voter models. Lecture Notes in Computer Science, 7323:1–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30541-2_1/.

Alamdari:2012:DPP

[425] Soroush Alamdari and Abbas Mehrabian. On a DAG partitioning problem. Lecture Notes in Computer Science, 7323:17–28, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30541-2_2/.

Cooper:2012:STP

[426] Colin Cooper, Alan Frieze, and Paweł Prałat. Some typical properties of the spatial preferred attachment model. Lecture Notes in Computer Science, 7323:29–40, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30541-2_3/.

Borgs:2012:STA

[427] Christian Borgs and Michael Brautbar. A sublinear time algorithm for PageRank computations. Lecture Notes in Computer Science, 7323:

41–53, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30541-2_4/.

Avrachenkov:2012:QDN

[428] Konstantin Avrachenkov and Nelly Litvak. Quick detection of nodes with large degrees. Lecture Notes in Computer Science, 7323:54–65, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30541-2_5/.

Chung:2012:RSC

[429] Fan Chung and Wenbo Zhao. Ranking and sparsifying a connection graph. Lecture Notes in Computer Science, 7323:66–77, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30541-2_6/.

Goel:2012:GTM

[430] Ashish Goel and Farnaz Ronaghi. A game-theoretic model of attention in social networks. Lecture Notes in Computer Science, 7323: 78–92, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30541-2_7/.

Frieze:2012:CPR

[431] Alan Frieze and Charalampos E. Tsourakakis. On certain properties of random Apollonian networks. Lecture Notes in Computer Science, 7323: 93–112, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30541-2_8/.

Li:2012:MUL

[432] Yanhua Li, Zhi-Li Zhang, and Jie Bao. Mutual or unrequited love: Identifying stable clusters in social networks with uni- and bi-directional links. Lecture Notes in Computer Science, 7323:113–125, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30541-2_9/.

Rossi:2012:DPU

[433] Ryan A. Rossi and David F. Gleich. Dynamic PageRank using evolving teleportation. Lecture Notes in Computer Science, 7323:126–137, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30541-2_10/.

Chung:2012:MCA

[434] Fan Chung, Paul Horn, and Jacob Hughes. Multi-commodity allocation for dynamic demands using PageRank vectors. Lecture Notes in Computer Science, 7323:138–152, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30541-2_11/.

Ray:2012:WTY

[435] Jaideep Ray, Ali Pinar, and C. Seshadhri. Are we there yet? When to stop a Markov chain while generating random graphs. Lecture Notes in Computer Science, 7323:153–164, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30541-2_12/.

Cooper:2012:FAF

[436] Colin Cooper, Tomasz Radzik, and Yiannis Siantos. A fast algorithm to find all high degree vertices in graphs with a power law degree sequence. Lecture Notes in Computer Science, 7323:165–178, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30541-2_13/.

Anonymous:2012:BMl

[437] Anonymous. Back matter. Lecture Notes in Computer Science, 7323: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-30541-2/1.

Anonymous:2012:FMw

[438] Anonymous. Front matter. Lecture Notes in Computer Science, 7323: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-30541-2/1.

Zhang:2012:LGB

[439] Zhihong Zhang and Edwin R. Hancock. Localized graph-based feature selection for clustering. Lecture Notes in Computer Science, 7324: 1–10, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31295-3_1/.

Thomas:2012:OAB

[440] A. Thomas and B. John Oommen. Optimal “anti-Bayesian” parametric pattern classification for the exponential family using order statistics criteria. Lecture Notes in Computer Science, 7324:11–18, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31295-3_2/.

Jarraya:2012:CSD

[441] Salma Kammoun Jarraya, Rania Rebai Boukhriss, Mohamed Hammami, and Hanene Ben-Abdallah. Cast shadow detection based on semisupervised learning. Lecture Notes in Computer Science, 7324:19–26, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31295-3_3/.

Aidos:2012:NNC

[442] Helena Aidos and Ana Fred. k-nearest neighbor classification using dissimilarity increments. Lecture Notes in Computer Science, 7324: 27–33, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31295-3_4/.

Al-Rawi:2012:UPT

[443] Mohammed Sadeq Al-Rawi and João Paulo Silva Cunha. Using permutation tests to study how the dimensionality, the number of classes, and the number of samples affect classification analysis. Lecture Notes in Computer Science, 7324:34–42, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31295-3_5/.

Kolesnikov:2012:DNC

[444] Alexander Kolesnikov and Elena Trichina. Determining the number of clusters with rate-distortion curve modeling. Lecture Notes in Computer Science, 7324:43–50, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31295-3_6/.

Sellaouti:2012:HCB

[445] Aymen Sellaouti, Atef Hamouda, Aline Deruyver, and Cédric Wemmert. Hierarchical classification-based region growing (HCBRG): a collaborative approach for object segmentation and classification. Lecture Notes in Computer Science, 7324:51–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31295-3_7/.

Goncalves:2012:CTS

[446] Nicolau Gonçalves and Ricardo Vigário. Clustering through SOM consistency. Lecture Notes in Computer Science, 7324:61–68, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31295-3_8/.

Glew:2012:MMV

[447] D. Glew and Edward R. Vrscay. Max and Min values of the structural similarity function S(x, a) on the L^{2} sphere S_{R}(a), a ∈ R^{N}. Lecture Notes in Computer Science, 7324:69–78, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31295-3_9/.

Shi:2012:BML

[448] Xun Shi, Neil D. B. Bruce, and John K. Tsotsos. Biologically motivated local contextual modulation improves low-level visual feature representations. Lecture Notes in Computer Science, 7324:79–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31295-3_10/.

Mittal:2012:SOD

[449] Ajay Mittal, Sanjeev Sofat, Edwin Hancock, and Stéphane Mousset. A statistical operator for detecting weak edges in low contrast images. Lecture Notes in Computer Science, 7324:89–96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31295-3_11/.

Oliveira:2012:CCU

[450] Miguel Oliveira, Angel D. Sappa, and Vítor Santos. Color correction using 3D Gaussian mixture models. Lecture Notes in Computer Science, 7324:97–106, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31295-3_12/.

Tomaszewska:2012:BNL

[451] Anna Tomaszewska. Blind noise level detection. Lecture Notes in Computer Science, 7324:107–114, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31295-3_13/.

Guyon:2012:FDR

[452] Charles Guyon, Thierry Bouwmans, and El-Hadi Zahzah. Foreground detection by robust PCA solved via a linearized alternating direction method. Lecture Notes in Computer Science, 7324:115–122, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31295-3_14/.

Peng:2012:MEA

[453] Peng Peng and Ze-Nian Li. A mixture of experts approach to multistrategy image quality assessment. Lecture Notes in Computer Science, 7324:123–130, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-31295-3_15/.

Glew:2012:SSI

[454] D. Glew and Edward R. Vrscay. Self-similarity of images in the wavelet domain in terms of ℓ^{2} and structural similarity (SSIM). Lecture Notes in Computer Science, 7324:131–140, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31295-3_16/.

Ahmad:2012:ABC

[455] Irfan Ahmad and Sabri A. Mahmoud. Arabic bank check analysis and zone extraction. Lecture Notes in Computer Science, 7324:141– 148, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31295-3_17/.

Gadermayr:2012:ISV

[456] Michael Gadermayr and Andreas Uhl. Image segmentation of Vickers indentations using shape from focus. Lecture Notes in Computer Science, 7324:149–157, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-31295-3_18/.

Pratas:2012:DUL

[457] Diogo Pratas and Armando J. Pinho. On the detection of unknown locally repeating patterns in images. Lecture Notes in Computer Science, 7324: 158–165, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31295-3_19/.

Saleem:2012:PES

[458] Sajid Saleem, Abdul Bais, and Robert Sablatnig. A performance evaluation of SIFT and SURF for multispectral image matching. Lecture Notes in Computer Science, 7324:166–173, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31295-3_20/.

Anonymous:2012:FMx

[459] Anonymous. Front matter. Lecture Notes in Computer Science, 7324: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31295-3/1.

Uhl:2012:MSV

[460] Andreas Uhl and Peter Wild. Multi-stage visible wavelength and near infrared Iris segmentation framework. Lecture Notes in Computer Science, 7325:1–10, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31298-4_1/.

Oliveira:2012:TUB

[461] Hélder P. Oliveira and Filipe Magalhães. Two unconstrained biometric databases. Lecture Notes in Computer Science, 7325:11–19, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31298-4_2/.

Roy:2012:MSU

[462] Kaushik Roy and Mohamed S. Kamel. Multibiometric system using level set method and particle swarm optimization. Lecture Notes in Computer Science, 7325:20–29, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31298-4_3/.

Al-Qarni:2012:EII

[463] Garsah Farhan Al-Qarni and Farzin Deravi. Explicit integration of identity information from skin regions to improve face recognition. Lecture Notes in Computer Science, 7325:30–37, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31298-4_4/.

Matos:2012:HGB

[464] Hélder Matos, Hélder P. Oliveira, and Filipe Magalhães. Hand-geometry based recognition system. Lecture Notes in Computer Science, 7325: 38–45, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31298-4_5/.

Zhou:2012:LGI

[465] Lu Bing Zhou and Han Wang. Local gradient increasing pattern (LGIP) for facial representation and gender recognition. Lecture Notes in Computer Science, 7325:46–53, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31298-4_6/.

Mngenge:2012:QBF

[466] Ntethelelo A. Mngenge, Fulufhelo V. Nelwamondo, Tendani Malumedzha, and Ntsika Msimang. Quality-based fingerprint segmentation. Lecture Notes in Computer Science, 7325:54–63, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31298-4_7/.

Niese:2012:NMC

[467] Robert Niese, Ayoub Al-Hamadi, and Bernd Michaelis. A new multicamera based facial expression analysis concept. Lecture Notes in Computer Science, 7325:64–71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31298-4_8/.

DeMarsico:2012:EBF

[468] Maria De Marsico, Michele Nappi, and Daniel Riccio. Entropy in biometric face template analysis. Lecture Notes in Computer Science, 7325: 72–79, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31298-4_9/.

Uhl:2012:CFF

[469] Andreas Uhl and Peter Wild. Combining face with face-part detectors under Gaussian assumption. Lecture Notes in Computer Science, 7325: 80–89, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31298-4_10/.

Dornaika:2012:GLD

[470] F. Dornaika and A. Bosaghzadeh. Generalized local discriminant embedding for face recognition. Lecture Notes in Computer Science, 7325: 90–97, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31298-4_11/.

Borras:2012:DIH

[471] Ricard Borràs, Àgata Lapedriza, and Laura Igual. Depth information in human gait analysis: An experimental study on gender recognition. Lecture Notes in Computer Science, 7325:98–105, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31298-4_12/.

Dubey:2012:DCB

[472] Rachit Dubey, Bingbing Ni, and Pierre Moulin. A depth camera based fall recognition system for the elderly. Lecture Notes in Computer Science, 7325:106–113, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-31298-4_13/.

Lahamy:2012:HVB

[473] Hervé Lahamy and Derek Lichti. Heuristic and voxel-based signature for hand posture recognition using a range camera. Lecture Notes in Computer Science, 7325:114–121, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31298-4_14/.

Diaf:2012:CER

[474] Abdunnaser Diaf, Boubakeur Boufama, and Rachid Benlamri. A compound eigenspace for recognizing directed human activities. Lecture Notes in Computer Science, 7325:122–129, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31298-4_15/.

Gil-Jimenez:2012:HDT

[475] Pedro Gil-Jiménez, Beatriz Losilla-López, Rafael Torres-Cueco, and Aurélio Campilho. Hand detection and tracking using the skeleton of the blob for medical rehabilitation applications. Lecture Notes in Computer Science, 7325:130–137, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31298-4_16/.

Tran:2012:DMI

[476] Minh-Phuong Tran, Renaud Péteri, and Maitine Bergounioux. Denoising 3D medical images using a second order variational model and wavelet shrinkage. Lecture Notes in Computer Science, 7325:138–145, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31298-4_17/.

Xu:2012:RIN

[477] Jinwei Xu and Tuan D. Pham. Robust impulse-noise filtering for biomedical images using numerical interpolation. Lecture Notes in Computer Science, 7325:146–155, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31298-4_18/.

Lyksborg:2012:SMS

[478] Mark Lyksborg, Rasmus Larsen, Per Soelberg Sørensen, Morten Blinkenberg, and Ellen Garde. Segmenting multiple sclerosis lesions using a spatially constrained K-nearest neighbour approach. Lecture Notes in Computer Science, 7325:156–163, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31298-4_19/.

Figueiredo:2012:SMA

[479] Isabel N. Figueiredo, Juan Carlos Moreno, V. B. Surya Prasath, and Pedro N. Figueiredo. A segmentation model and application to endoscopic images. Lecture Notes in Computer Science, 7325:164–171, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31298-4_20/.

Anonymous:2012:FMy

[480] Anonymous. Front matter. Lecture Notes in Computer Science, 7325: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31298-4/1.

Ketter:2012:CSL

[481] Wolfgang Ketter. Competitive simulations: Lessons learned from the trading agent competition. Lecture Notes in Computer Science, 7327:1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-30947-2_1.

Luck:2012:BRN

[482] Michael Luck. Behaviour regulation and normative systems. Lecture Notes in Computer Science, 7327:2, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

accesspage/chapter/10.1007/978-3-642-30947-2_2.

Valavanis:2012:CUS

[483] Kimon Valavanis. Challenges in unmanned systems swarms: a wireless, multi-agent distributed system perspective. Lecture Notes in Computer Science, 7327:3, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/accesspage/
chapter/10.1007/978-3-642-30947-2_3.

Lujak:2012:MTA

[484] Marin Lujak, Holger Billhardt, and Sascha Ossowski. On mobile target allocation with incomplete information in defensive environments. Lecture Notes in Computer Science, 7327:4–13, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30947-2_4/.

Seric:2012:BPF

[485] Ljiljana Šerić, Maja Štula, Darko Stipaničev, and Maja Braović. Bayesian proprioceptor for forest fire observer network. Lecture Notes in Computer Science, 7327:14–23, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30947-2_5/.

Capkovic:2012:SAM

[486] František Čapkovič. Supervision of agents modelling evacuation at crisis situations. Lecture Notes in Computer Science, 7327:24–33, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30947-2_6/.

Torres:2012:WSC

[487] Romina Torres, Denise Rivera, and Hernan Astudillo. Web service compositions which emerge from virtual organizations with fair agreements. Lecture Notes in Computer Science, 7327:34–43, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-30947-2_7/.

Kozierkiewicz-Hetmanska:2012:EIT

[488] Adrianna Kozierkiewicz-Hetmanska. Evaluation of an intelligent tutoring system incorporating learning profile to determine learning scenario. Lecture Notes in Computer Science, 7327:44–53, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-30947-2_8/.

Le:2012:SBL

[489] Nguyen-Thinh Le and Niels Pinkwart. Strategy-based learning through communication with humans. Lecture Notes in Computer Science, 7327: 54–64, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30947-2_9/.

Choinski:2012:OBK

[490] Dariusz Choinski and Michal Senik. Ontology based knowledge management and learning in multi-agent system. Lecture Notes in Computer Science, 7327:65–74, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30947-2_10/.

Ogiela:2012:DAS

[491] Lidia Ogiela and Marek R. Ogiela. Data analysis systems and cognitive modelling processes. Lecture Notes in Computer Science, 7327: 75–83, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30947-2_11/.

Maleszka:2012:POI

[492] Marcin Maleszka and Ngoc Thanh Nguyen. Path-oriented integration method for complex trees. Lecture Notes in Computer Science, 7327: 84–93, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30947-2_12/.

Vranic:2012:IQA

[493] Mihaela Vranić, Damir Pintar, and Zoran Skočir. Integrating quantitative attributes in hierarchical clustering of transactional data. Lecture Notes in Computer Science, 7327:94–103, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30947-2_13/.

Skocir:2012:MMA

[494] Pavle Skocir, Luka Marusic, Marinko Marusic, and Ana Petric. The MARS — a multi-agent recommendation system for games on mobile phones. Lecture Notes in Computer Science, 7327:104–113, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30947-2_14/.

Juric:2012:IIN

[495] Damir Jurić, Marko Banek, and Šandor Dembitz. Informativeness of inflective noun bigrams in croatian. Lecture Notes in Computer Science, 7327:114–123, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-30947-2_15/.

Anderson:2012:MAN

[496] Paul Anderson, Shahriar Bijani, and Alexandros Vichos. Multi-agent negotiation of virtual machine migration using the lightweight coordination calculus. Lecture Notes in Computer Science, 7327:124–133, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30947-2_16/.

Yang:2012:LCM

[497] Cheng-Lin Yang and Yun-Heh Chen-Burger. On-line communities making scense: a hybrid micro-blogging platform community analysis framework. Lecture Notes in Computer Science, 7327:134–143, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30947-2_17/.

Rodriguez:2012:MET

[498] Ricardo J. Rodríguez, Rafael Tolosana-Calasanz, and Omer F. Rana. Measuring the effectiveness of throttled data transfers on data-intensive workflows. Lecture Notes in Computer Science, 7327:144–153, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30947-2_18/.

Nadarajan:2012:GVD

[499] Gayathri Nadarajan and Yun-Heh Chen-Burger. Goal, video description and capability ontologies for Fish4Knowledge domain. Lecture Notes in Computer Science, 7327:154–163, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30947-2_19/.

Voras:2012:ECC

[500] Ivan Voras, Marin Orlić, and Branko Mihaljević. An early comparison of commercial and open-source cloud platforms for scientific environments. Lecture Notes in Computer Science, 7327:164–173, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30947-2_20/.

Anonymous:2012:FMz

[501] Anonymous. Front matter. Lecture Notes in Computer Science, 7327: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-30947-2/1.

Missikoff:2012:FES

[502] Michele Missikoff. The future of enterprise systems in a fully networked society. Lecture Notes in Computer Science, 7328:1–18, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31095-9_1/.

Kurowski:2012:CFP

[503] Krzysztof Kurowski. Challenges for future platforms, services and networked applications. Lecture Notes in Computer Science, 7328:19–30, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31095-9_2/.

Dumas:2012:UBP

[504] Marlon Dumas, Marcello La Rosa, Jan Mendling, Raul Mäesalu, and Hajo A. Reijers. Understanding business process models: The costs and benefits of structuredness. Lecture Notes in Computer Science, 7328: 31–46, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31095-9_3/.

Petrusel:2012:AIM

[505] Razvan Petrusel. Aggregating individual models of decision-making processes. Lecture Notes in Computer Science, 7328:47–63, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31095-9_4/.

Leopold:2012:GNL

[506] Henrik Leopold, Jan Mendling, and Artem Polyvyanyy. Generating natural language texts from business process models. Lecture Notes in Computer Science, 7328:64–79, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31095-9_5/.

Schmeling:2012:TCF

[507] Benjamin Schmeling, Anis Charfi, Marko Martin, and Mira Mezini. Towards conflict-free composition of non-functional concerns. Lecture Notes in Computer Science, 7328:80–94, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31095-9_6/.

Razo-Zapata:2012:FVS

[508] Iván S. Razo-Zapata, Pieter De Leenheer, Jaap Gordijn, and Hans Akkermans. Fuzzy verification of service value networks. Lecture Notes in Computer Science, 7328:95–110, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31095-9_7/.

Mehandjiev:2012:CSC

[509] Nikolay Mehandjiev, Freddy Lécué, Martin Carpenter, and Fethi A. Rabhi. Cooperative service composition. Lecture Notes in Computer Science, 7328:111–126, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31095-9_8/.

deLara:2012:AML

[510] Juan de Lara, Esther Guerra, and Jesús Sánchez-Cuadrado. Abstracting modelling languages: a reutilization approach. Lecture Notes in Computer Science, 7328:127–143, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31095-9_9/.

Kopke:2012:LIS

[511] Julius Köpke and Johann Eder. Logical invalidations of semantic annotations. Lecture Notes in Computer Science, 7328:144–159, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31095-9_10/.

Atzeni:2012:UAN

[512] Paolo Atzeni, Francesca Bugiotti, and Luca Rossi. Uniform access to non-relational database systems: The SOS platform. Lecture Notes in Computer Science, 7328:160–174, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31095-9_11/.

Ghaddar:2012:VSO

[513] Ali Ghaddar, Dalila Tamzalit, Ali Assaf, and Abdalla Bitar. Variability as a service: Outsourcing variability management in multi-tenant SaaS applications. Lecture Notes in Computer Science, 7328:175–189, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31095-9_12/.

Lonn:2012:CPM

[514] Carl-Mikael Lönn, Elin Uppström, Petia Wohed, and Gustaf Juell-Skielse. Configurable process models for the Swedish public sector. Lecture Notes in Computer Science, 7328:190–205, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31095-9_13/.

Dalpiaz:2012:ASC

[515] Fabiano Dalpiaz, Raian Ali, and Paolo Giorgini. Aligning software configuration with business and IT context. Lecture Notes in Computer Science, 7328:206–221, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31095-9_14/.

Engel:2012:MIO

[516] Robert Engel, Wil M. P. van der Aalst, Marco Zapletal, and Christian Pichler. Mining inter-organizational business process models from EDI messages: a case study from the automotive sector. Lecture Notes in Computer Science, 7328:222–237, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31095-9_15/.

Ly:2012:DTS

[517] Linh Thao Ly, Conrad Indiono, Jürgen Mangler, and Stefanie Rinderle-Ma. Data transformation and semantic log purging for process mining. Lecture Notes in Computer Science, 7328:238–253, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31095-9_16/.

vandenBroucke:2012:IAN

[518] Seppe K. L. M. vanden Broucke, Jochen De Weerdt, Bart Baesens, and Jan Vanthienen. Improved artificial negative event generation to enhance process event logs. Lecture Notes in Computer Science, 7328: 254–269, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31095-9_17/.

Maggi:2012:EDU

[519] Fabrizio M. Maggi, R. P. Jagadeesh Chandra Bose, and Wil M. P. van der Aalst. Efficient discovery of understandable declarative process models from event logs. Lecture Notes in Computer Science, 7328: 270–285, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31095-9_18/.

Daskalaki:2012:OMS

[520] Evangelia Daskalaki and Dimitris Plexousakis. OtO matching system: a multi-strategy approach to instance matching. Lecture Notes in Computer Science, 7328:286–300, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31095-9_19/.

Kabir:2012:SSC

[521] Muhammad Ashad Kabir, Jun Han, Jian Yu, and Alan Colman. SCIMS: a social context information management system for socially-aware applications. Lecture Notes in Computer Science, 7328:301–317, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31095-9_20/.

Anonymous:2012:FMba

[522] Anonymous. Front matter. Lecture Notes in Computer Science, 7328: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31095-9/1.

diBaja:2012:LRO

[523] Gabriella Sanniti di Baja, L. Serino, and Carlo Arcelli. From linear representations to object parts. Lecture Notes in Computer Science, 7329:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31149-9_1/.

Dickinson:2012:PGU

[524] Sven J. Dickinson, Alex Levinshtein, and Cristian Sminchisescu. Perceptual grouping using superpixels. Lecture Notes in Computer Science, 7329:13–22, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31149-9_2/.

Sossa:2012:ADA

[525] Humberto Sossa, Beatriz A. Garro, Juan Villegas, Carlos Avilés, and Gustavo Olague. Automatic design of artificial neural networks and associative memories for pattern classification and pattern restoration. Lecture Notes in Computer Science, 7329:23–34, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31149-9_3/.

Frucci:2012:AIS

[526] Maria Frucci, Carlo Arcelli, and Gabriella Sanniti di Baja. An automatic image scaling up algorithm. Lecture Notes in Computer Science, 7329:35–44, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31149-9_4/.

Madrid:2012:STM

[527] Humberto Madrid, Valia Guerra, and Marielba Rojas. Sampling techniques for Monte Carlo matrix multiplication with applications to image processing. Lecture Notes in Computer Science, 7329:45–54, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31149-9_5/.

Hernandez-Rodriguez:2012:EPS

[528] Felipe Hernández-Rodríguez and Mario Castelán. Extended photometric sampling for surface shape recovery. Lecture Notes in Computer Science, 7329:55–64, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31149-9_6/.

Silvan-Cardenas:2012:SMT

[529] José Luis Silván-Cárdenas. A segmentation method for tree crown detection and modelling from LiDAR measurements. Lecture Notes in Computer Science, 7329:65–74, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31149-9_7/.

Kawulok:2012:TAS

[530] Michal Kawulok. Texture analysis for skin probability maps refinement. Lecture Notes in Computer Science, 7329:75–84, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31149-9_8/.

Aguilar-Gonzalez:2012:CCF

[531] Pablo M. Aguilar-González and Vitaly Kober. Composite correlation filters for detection of geometrically distorted objects using noisy training images. Lecture Notes in Computer Science, 7329:85–93, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31149-9_9/.

Ramirez-Acosta:2012:ASC

[532] Alejandro A. Ramírez-Acosta, Mireya S. García-Vázquez, and Mariko Nakano. Adaptive spatial concealment of damaged coded images. Lecture Notes in Computer Science, 7329:94–106, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31149-9_10/.

Saldivar-Pinon:2012:HSR

[533] Leonardo Saldivar-Piñon, Mario I. Chacon-Murguia, and Rafael Sandoval-Rodriguez. Human sign recognition for robot manipulation. Lecture Notes in Computer Science, 7329:107–116, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31149-9_11/.

Sokolova:2012:FSH

[534] Marina V. Sokolova and Antonio Fernández-Caballero. Fuzzy sets for human fall pattern recognition. Lecture Notes in Computer Science, 7329: 117–126, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31149-9_12/.

Espino:2012:VSR

[535] José Guadalupe Rico Espino, José-Joel Gonzalez-Barbosa, and Roberto Augusto Gómez Loenzo. Vision system for 3D reconstruction with telecentric lens. Lecture Notes in Computer Science, 7329:127–136, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31149-9_13/.

Soria:2012:THS

[536] David J. Rios Soria and Satu Elisa Schaeffer. A tool for hand-sign recognition. Lecture Notes in Computer Science, 7329:137–146, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31149-9_14/.

Medina-Perez:2012:IMA

[537] Miguel Angel Medina-Pérez, Milton García-Borroto, and Andres Eduardo Gutierrez-Rodríguez. Improving the multiple alignments strategy for fingerprint verification. Lecture Notes in Computer Science, 7329: 147–154, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31149-9_15/.

Cruz-Perez:2012:BRU

[538] Claudia Cruz-Perez, Oleg Starostenko, Fernando Uceda-Ponga, and Vicente Alarcon-Aquino. Breaking reCAPTCHAs with unpredictable collapse: Heuristic character segmentation and recognition. Lecture Notes in Computer Science, 7329:155–165, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31149-9_16/.

Toca:2012:USR

[539] Cosme E. Santiesteban Toca, Milton García-Borroto, and Jesus S. Aguilar Ruiz. Using short-range interactions and simulated genetic strategy to improve the protein contact map prediction. Lecture Notes in Computer Science, 7329:166–175, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31149-9_17/.

Navarro:2012:AMS

[540] Rodolfo Navarro, Elena Acevedo, Antonio Acevedo, and Fabiola Martínez. Associative model for solving the wall-following problem. Lecture Notes in Computer Science, 7329:176–186, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31149-9_18/.

Tellez:2012:LCR

[541] Eric Sadit Tellez and Edgar Chávez. The list of clusters revisited. Lecture Notes in Computer Science, 7329:187–196, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31149-9_19/.

Parvin:2012:HPD

[542] Hamid Parvin, Sajad Parvin, Zahra Rezaei, and Moslem Mohamadi. A heuristically perturbation of dataset to achieve a diverse ensemble of classifiers. Lecture Notes in Computer Science, 7329:197–206, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31149-9_20/.

Anonymous:2012:FMbb

[543] Anonymous. Front matter. Lecture Notes in Computer Science, 7329: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31149-9/1.

Malti:2012:TBC

[544] Abed Malti, Adrien Bartoli, and Toby Collins. Template-based conformal shape-from-motion-and-shading for laparoscopy. Lecture Notes in Computer Science, 7330:1–10, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30618-1_1/.

Collins:2012:TLM

[545] Toby Collins and Adrien Bartoli. Towards live monocular 3D laparoscopy using shading and specularity information. Lecture Notes in Computer Science, 7330:11–21, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30618-1_2/.

Mohareri:2012:ADL

[546] Omid Mohareri, Mahdi Ramezani, Troy Adebar, Purang Abolmaesumi, and Septimiu Salcudean. Automatic detection and localization of da Vinci tool tips in 3D ultrasound. Lecture Notes in Computer Science, 7330:22–32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30618-1_3/.

Yip:2012:RTM

[547] Michael C. Yip, David G. Lowe, Septimiu E. Salcudean, and Robert N. Rohling. Real-time methods for long-term tissue feature tracking in endoscopic scenes. Lecture Notes in Computer Science, 7330:33–43, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30618-1_4/.

Najafi:2012:CFD

[548] Mohammad Najafi, Narges Afsham, Purang Abolmaesumi, and Robert Rohling. A closed-form differential formulation for ultrasound spatial calibration. Lecture Notes in Computer Science, 7330:44–53, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30618-1_5/.

Arnold:2012:MBR

[549] Patrik Arnold, Frank Preiswerk, Beat Fasel, Rares Salomir, and Klaus Scheffler. Model-based respiratory motion compensation in MRgHIFU.

Lecture Notes in Computer Science, 7330:54–63, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-30618-1_6/.

Hacihaliloglu:2012:NIM

[550] Ilker Hacihaliloglu, David R. Wilson, Michael Gilbart, Michael Hunt, and Purang Abolmaesumi. Non-iterative multi-modal partial view to full view image registration using local phase-based image projections. Lecture Notes in Computer Science, 7330:64–73, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30618-1_7/.

Schumann:2012:HSR

[551] Steffen Schumann, Moritz Tannast, Mathias Bergmann, Michael Thali, and Lutz-P. Nolte. A hierarchical strategy for reconstruction of 3D acetabular surface models from 2D calibrated X-ray images. Lecture Notes in Computer Science, 7330:74–83, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30618-1_8/.

Moore:2012:NPG

[552] John Moore, Chris Wedlake, Daniel Bainbridge, Gerard Guiraudon, Michael Chu, and Bob Kiaii. A navigation platform for guidance of beating heart transapical mitral valve repair. Lecture Notes in Computer Science, 7330:84–93, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30618-1_9/.

Kaeppler:2012:MEM

[553] Sebastian Kaeppler, Alexander Brost, Martin Koch, Wen Wu, Felix Bourier, and Terrence Chen. Motion estimation model for cardiac and respiratory motion compensation. Lecture Notes in Computer Science, 7330:94–103, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30618-1_10/.

Ma:2012:CUN

[554] YingLiang Ma, Rashed Karim, R. James Housden, Geert Gijsbers, and Roland Bullens. Cardiac unfold: a novel technique for image-guided cardiac catheterization procedures. Lecture Notes in Computer Science, 7330: 104–114, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30618-1_11/.

Brattain:2012:EUP

[555] Laura J. Brattain, Nikolay V. Vasilyev, and Robert D. Howe. Enabling 3D ultrasound procedure guidance through enhanced visualization. Lecture Notes in Computer Science, 7330:115–124, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30618-1_12/.

DeNigris:2012:FRR

[556] Dante De Nigris, D. Louis Collins, and Tal Arbel. Fast and robust registration based on gradient orientations: Case study matching intraoperative ultrasound to pre-operative MRI in neurosurgery. Lecture Notes in Computer Science, 7330:125–134, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30618-1_13/.

Xiao:2012:ABS

[557] Yiming Xiao, Lara Bailey, M. Mallar Chakravarty, Silvain Beriault, and Abbas F. Sadikot. Atlas-based segmentation of the subthalamic nucleus, red nucleus, and substantia nigra for deep brain stimulation by incorporating multiple MRI contrasts. Lecture Notes in Computer Science, 7330: 135–145, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30618-1_14/.

Bigdelou:2012:TSU

[558] Ali Bigdelou, Aslı Okur, Max-Emanuel Hoffmann, Bamshad Azizi, and Nassir Navab. Towards systematic usability evaluations for the OR: An introduction to OR-use framework. Lecture Notes in Computer Science, 7330:146–156, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-30618-1_15/.

Hernansanz:2012:IDS

[559] Albert Hernansanz, Davide Zerbato, Lorenza Gasperotti, Michele Scandola, and Paolo Fiorini. Improving the development of surgical skills with virtual fixtures in simulation. Lecture Notes in Computer Science, 7330: 157–166, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30618-1_16/.

Tao:2012:SHM

[560] Lingling Tao, Ehsan Elhamifar, Sanjeev Khudanpur, Gregory D. Hager, and René Vidal. Sparse hidden Markov models for surgical gesture classification and skill evaluation. Lecture Notes in Computer Science, 7330: 167–177, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30618-1_17/.

Anonymous:2012:BMm

[561] Anonymous. Back matter. Lecture Notes in Computer Science, 7330: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-30618-1/1.

Anonymous:2012:FMbc

[562] Anonymous. Front matter. Lecture Notes in Computer Science, 7330: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-30618-1/1.

Konigsberg:2012:BIS

[563] Zvi Retchkiman Konigsberg. The biological interaction stability problem. Lecture Notes in Computer Science, 7331:1–10, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.com/chapter/10.1007/978-3-642-30976-2_1/.

Folly:2012:PBI

[564] Komla A. Folly. Population-based incremental with adaptive learning rate strategy. Lecture Notes in Computer Science, 7331:11–20, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30976-2_2/.

Yu:2012:SBA

[565] Ling Yu, Peng Xu, and Xi Chen. A SI-based algorithm for structural damage detection. Lecture Notes in Computer Science, 7331: 21–28, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30976-2_3/.

Cao:2012:QIB

[566] Jinlong Cao and Hongyuan Gao. A quantum-inspired bacterial swarming optimization algorithm for discrete optimization problems. Lecture Notes in Computer Science, 7331:29–36, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30976-2_4/.

Kundu:2012:SIC

[567] Anirban Kundu and Chunlin Ji. Swarm intelligence in cloud environment. Lecture Notes in Computer Science, 7331:37–44, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30976-2_5/.

Xing:2012:SIS

[568] Bo Xing, Wen-Jing Gao, Fulufhelo V. Nelwamondo, Kimberly Battle, and Tshilidzi Marwala. Swarm intelligence supported e-remanufacturing. Lecture Notes in Computer Science, 7331:45–52, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-30976-2_6/.

Yeh:2012:GBP

[569] Ming-Feng Yeh, Cheng Wen, and Min-Shyang Leu. Grey-based particle swarm optimization algorithm. Lecture Notes in Computer Science, 7331:53–62, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30976-2_7/.

Li:2012:QBP

[570] Ya Li, Dan Li, and Dong Wang. Quantum-behaved particle swarm optimization algorithm based on border mutation and chaos for vehicle routing problem. Lecture Notes in Computer Science, 7331:63–73, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30976-2_8/.

Li:2012:IMC

[571] Wei xing Li, Qian Zhou, Yu Zhu, and Feng Pan. An improved MOPSO with a crowding distance based external archive maintenance strategy. Lecture Notes in Computer Science, 7331:74–82, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-30976-2_9/.

Ting:2012:EIW

[572] T. O. Ting, Yuhui Shi, Shi Cheng, and Sanghyuk Lee. Exponential inertia weight for particle swarm optimization. Lecture Notes in Computer Science, 7331:83–90, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30976-2_10/.

Zhou:2012:CMP

[573] Jiarui Zhou, Zhen Ji, Zexuan Zhu, and Siping Chen. A coevolutionary memetic particle swarm optimizer. Lecture Notes in Computer Science, 7331:91–100, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30976-2_11/.

Wei:2012:HMO

[574] Wei Wei, Weihui Zhang, Yuan Jiang, and Hao Li. Handling multioptimization with gender-hierarchy based particle swarm optimizer. Lecture Notes in Computer Science, 7331:101–108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-30976-2_12/.

Huang:2012:CSD

[575] Guoliang Huang, Xinling Shi, and Zhenzhou An. The comparative study of different number of particles in clustering based on two-layer particle swarm optimization. Lecture Notes in Computer Science, 7331: 109–115, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30976-2_13/.

Tian:2012:IPS

[576] Yubo Tian, Donghui Gao, and Xiaolong Li. Improved particle swarm optimization with wavelet-based mutation operation. Lecture Notes in Computer Science, 7331:116–124, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30976-2_14/.

Chi:2012:EBP

[577] Yuhong Chi, Fuchun Sun, Langfan Jiang, Chunming Yu, and Ping Zhang. Elastic boundary for particle swarm optimization. Lecture Notes in Computer Science, 7331:125–132, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30976-2_15/.

Tsai:2012:OLW

[578] Ming-Tang Tsai and Szu-Wzi Wu. Optimization locations of wind turbines with the particle swarm optimization. Lecture Notes in Computer Science, 7331:133–141, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-30976-2_16/.

Liu:2012:PBA

[579] Zhanghui Liu and Xiaoli Wang. A PSO-based algorithm for load balancing in virtual machines of cloud computing environment. Lecture Notes in Computer Science, 7331:142–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30976-2_17/.

Lin:2012:TAP

[580] Xiufang Lin, Jun Sun, Vasile Palade, Wei Fang, Xiaojun Wu, and Wenbo Xu. Training ANFIS parameters with a quantum-behaved particle swarm optimization algorithm. Lecture Notes in Computer Science, 7331: 148–155, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30976-2_18/.

Sun:2012:RIM

[581] Ying Sun and Yue lin Gao. Research on improved model of loans portfolio optimization based on adaptive particle swarm optimization algorithm. Lecture Notes in Computer Science, 7331:156–163, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30976-2_19/.

Chou:2012:HDO

[582] Penchen Chou. High-dimension optimization problems using specified particle swarm optimization. Lecture Notes in Computer Science, 7331: 164–172, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30976-2_20/.

Anonymous:2012:FMbd

[583] Anonymous. Front matter. Lecture Notes in Computer Science, 7331: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-30976-2/1.

Razavi:2012:ALS

[584] Seyed Naser Razavi, Nicolas Gaud, Abderrafiâa Koukam, and Naser Mozayani. An automatic learning system to derive multipole and local expansions for the Fast Multipole Method. Lecture Notes in Computer Science, 7332:1–10, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31020-1_1/.

Liu:2012:IRA

[585] Cheng Liu, Yong Liang, Xin-Ze Luan, Kwong-Sak Leung, Tak-Ming Chan, and Zong-Ben Xu. Iterative L_{1/2} regularization algorithm for variable selection in the Cox proportional hazards model. Lecture Notes in Computer Science, 7332:11–17, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31020-1_2/.

Zhang:2012:ASE

[586] Junbo Zhang, Fuping Pan, and Yongyong Yan. Automatic scoring on English passage reading quality. Lecture Notes in Computer Science, 7332:18–25, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31020-1_3/.

Song:2012:LSB

[587] Bo Song and Miaoyan Li. An e-learning system based on GWT and Berkeley DB. Lecture Notes in Computer Science, 7332:26–32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31020-1_4/.

Xiao:2012:ERS

[588] Jie Xiao and Liang He. An expandable recommendation system on IPTV. Lecture Notes in Computer Science, 7332:33–40, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31020-1_5/.

Cheng:2012:IPD

[589] Wei-Chen Cheng. Intrinsic protein distribution on manifolds embedded in low-dimensional space. Lecture Notes in Computer Science, 7332: 41–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31020-1_6/.

You:2012:NAM

[590] Zhuhong You, Yingke Lei, Zhen Ji, and Zexuan Zhu. A novel approach to modelling protein-protein interaction networks. Lecture Notes in Computer Science, 7332:49–57, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31020-1_7/.

Zhang:2012:AOP

[591] Jiuling Zhang, Beixing Deng, and Xing Li. Additive order preserving encryption based encrypted documents ranking in secure cloud storage. Lecture Notes in Computer Science, 7332:58–65, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31020-1_8/.

Wu:2012:RWI

[592] Jian Wu and Siyong Xiong. Research of Web image retrieval technology based on hu invariant moments. Lecture Notes in Computer Science, 7332:66–73, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31020-1_9/.

Huang:2012:CBM

[593] Xi Huang, Ying Tan, and Xingui He. A classifier based on minimum circum circle. Lecture Notes in Computer Science, 7332:74–83, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31020-1_10/.

Yang:2012:RDS

[594] Xiquan Yang, Meijia Wang, Lin Fang, Lin Yue, and Yinghua Lv. Research on domain-specific features clustering based spectral clustering. Lecture Notes in Computer Science, 7332:84–92, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31020-1_11/.

Wei:2012:IAK

[595] Yang Wei. An iterative approach to keywords extraction. Lecture Notes in Computer Science, 7332:93–99, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31020-1_12/.

Liang:2012:KAF

[596] Rupeng Liang, Hongwei Li, Jian Chen, Leilei Ma, and Hu Chen. Knowledge annotation framework oriented geospatial semantic Web service management. Lecture Notes in Computer Science, 7332:100–107, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31020-1_13/.

Zang:2012:OSS

[597] Wenjuan Zang, Yankui Liu, and Zhenhong Li. Optimizing supplier selection with disruptions by chance-constrained programming. Lecture Notes in Computer Science, 7332:108–116, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31020-1_14/.

Bellaachia:2012:FLN

[598] Abdelghani Bellaachia and Anasse Bari. Flock by leader: a novel machine learning biologically inspired clustering algorithm. Lecture Notes in Computer Science, 7332:117–126, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31020-1_15/.

Touzi:2012:CVC

[599] Amel Grissa Touzi, Amira Aloui, and Rim Mahouachi. Cluster KDD: a visual clustering and knowledge discovery platform based on concept lattice. Lecture Notes in Computer Science, 7332:127–136, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31020-1_16/.

Qu:2012:DII

[600] Zhe Qu and Qin Wang. Design and implementation of an intelligent automatic question answering system based on data mining. Lecture Notes in Computer Science, 7332:137–146, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31020-1_17/.

Zeng:2012:CEC

[601] Huanglin Zeng and Xuefei Tang. Comprehensive evaluation of Chinese liquor quality based on improved gray-clustering analysis. Lecture Notes in Computer Science, 7332:147–154, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31020-1_18/.

Yang:2012:OBH

[602] Xiquan Yang, Rui Gao, Zhengfu Han, and Xin Sui. Ontology-based hazard information extraction from Chinese food complaint documents. Lecture Notes in Computer Science, 7332:155–163, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31020-1_19/.

Liu:2012:NCF

[603] Qun Liu, Yi Gao, and Zhiming Peng. A novel collaborative filtering algorithm based on social network. Lecture Notes in Computer Science, 7332: 164–174, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31020-1_20/.

Anonymous:2012:FMbe

[604] Anonymous. Front matter. Lecture Notes in Computer Science, 7332: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31020-1/1.

Versaci:2012:PAO

[605] Francesco Versaci and Keshav Pingali. Processor allocation for optimistic parallelization of irregular programs. Lecture Notes in Computer Science, 7333:1–14, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31125-3_1/.

Timm:2012:FBG

[606] Constantin Timm, Markus Görlich, Frank Weichert, Peter Marwedel, and Heinrich Müller. Feedback-based global instruction scheduling for GPGPU applications. Lecture Notes in Computer Science, 7333:15– 28, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31125-3_2/.

Boratto:2012:PAL

[607] Murilo Boratto, Pedro Alonso, Carla Ramiro, Marcos Barreto, and Leandro Coelho. Parallel algorithm for landform attributes representation on multicore and multi-GPU systems. Lecture Notes in Computer Science, 7333:29–43, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31125-3_3/.

Epicoco:2012:PME

[608] Italo Epicoco and Silvia Mocavero. The performance model of an enhanced parallel algorithm for the SOR method. Lecture Notes in Computer Science, 7333:44–56, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31125-3_4/.

Sanjuan-Estrada:2012:PDC

[609] Juan Francisco Sanjuan-Estrada, Leocadio Gonzalez Casado, and Immaculada García. Performance driven cooperation between kernel and auto-tuning multi-threaded interval B&B applications. Lecture Notes in Computer Science, 7333:57–70, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31125-3_5/.

Arefin:2012:NBG

[610] Ahmed Shamsul Arefin, Carlos Riveros, Regina Berretta, and Pablo Moscato. k NN-Borůvka–GPU: a fast and scalable MST construction from k NN graphs on GPU. Lecture Notes in Computer Science, 7333: 71–86, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31125-3_6/.

Shah:2012:GHA

[611] Habib Shah, Rozaida Ghazali, Nazri Mohd Nawi, and Mustafa Mat Deris. Global hybrid ant bee colony algorithm for training artificial neural networks. Lecture Notes in Computer Science, 7333:87–100, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31125-3_7/.

Alves:2012:EIE

[612] Daniel S. F. Alves, Felipe M. G. França, Luiza de Macedo Mourelle, and Nadia Nedjah. The effect of intelligent escape on distributed SER-based search. Lecture Notes in Computer Science, 7333:101–112, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31125-3_8/.

Silva:2012:ABS

[613] Luneque Silva Jr., Nadia Nedjah, Luiza de Macedo Mourelle, and Fábio Gonçalves Pessanha. ACO-based static routing for network-on-chips. Lecture Notes in Computer Science, 7333:113–124, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31125-3_9/.

Fonseca:2012:GAA

[614] Leonardo G. Fonseca, Heder S. Bernardino, and Helio J. C. Barbosa. A genetic algorithm assisted by a locally weighted regression surrogate model. Lecture Notes in Computer Science, 7333:125–135, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31125-3_10/.

deMendonca:2012:SRQ

[615] Rafael Mathias de Mendonça, Nadia Nedjah, and Luiza de Macedo Mourelle. Swarm robots with queue organization using infrared communication. Lecture Notes in Computer Science, 7333:136–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31125-3_11/.

Calazan:2012:SGP

[616] Rogério M. Calazan, Nadia Nedjah, and Luiza de Macedo Mourelle. Swarm grid: a proposal for high performance of parallel particle swarm optimization using GPGPU. Lecture Notes in Computer Science, 7333: 148–160, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31125-3_12/.

Elsayed:2012:AIS

[617] Samir A. Mohamed Elsayed, Sanguthevar Rajasekaran, and Reda A. Ammar. An artificial immune system approach to associative classification. Lecture Notes in Computer Science, 7333:161–171, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31125-3_13/.

deOliveira:2012:RDR

[618] Sanderson L. Gonzaga de Oliveira. A review on Delaunay refinement techniques. Lecture Notes in Computer Science, 7333:172–187, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31125-3_14/.

Ahmadian:2012:APD

[619] Kushan Ahmadian and Marina Gavrilova. Axis-parallel dimension reduction for biometric research. Lecture Notes in Computer Science, 7333: 188–197, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31125-3_15/.

deOliveira:2012:OPR

[620] Sanderson L. Gonzaga de Oliveira. An overview of procedures for refining triangulations. Lecture Notes in Computer Science, 7333:198– 213, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31125-3_16/.

Choi:2012:ICU

[621] Joonsoo Choi, Jaewee Heo, Kwang-Soo Hahn, and Junho Kim. DEM interpolation from contours using medial axis transformation. Lecture Notes in Computer Science, 7333:214–227, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31125-3_17/.

Filho:2012:AHD

[622] José Luiz de Souza Filho, Roger Correia Silva, Dhiego Oliveira Sad, and Renan Dembogurski. Analysis of a high definition camera-projector video system for geometry reconstruction. Lecture Notes in Computer Science, 7333:228–239, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-31125-3_18/.

Pereira:2012:VBF

[623] Tiago F. Pereira, Marcus A. Angeloni, Flávio O. Simões, and José Eduardo C. Silva. Video-based face verification with local binary patterns and SVM using GMM supervectors. Lecture Notes in Computer Science, 7333:240–252, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-31125-3_19/.

Fort:2012:GBI

[624] Marta Fort and J. Antoni Sellarès. GPU-based influence regions optimization. Lecture Notes in Computer Science, 7333:253–266, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31125-3_20/.

Anonymous:2012:FMbf

[625] Anonymous. Front matter. Lecture Notes in Computer Science, 7333: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31125-3/1.

deFelice:2012:KIM

[626] Annunziata de Felice, Isabella Martucci, and Dario Antonio Schirone. Knowledge and innovation in manufacturing sector: The case of wedding dresses in Southern Italy. Lecture Notes in Computer Science, 7334:1–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31075-1_1/.

Schirone:2012:MSS

[627] Dario Antonio Schirone and Germano Torkan. Marketing strategies: Support and enhancement of core business. Lecture Notes in Computer Science, 7334:17–26, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-31075-1_2/.

DeMare:2012:RQS

[628] Gianluigi De Mare, Antonio Nesticò, and Francesco Tajani. The rational quantification of social housing. Lecture Notes in Computer Science, 7334:27–43, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31075-1_3/.

DellOrco:2012:SUD

[629] Mauro Dell’Orco and Michele Ottomanelli. Simulation of users decision in transport mode choice using neuro-fuzzy approach. Lecture Notes in Computer Science, 7334:44–53, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31075-1_4/.

Cerreta:2012:MSD

[630] Maria Cerreta, Simona Panaro, and Daniele Cannatella. Multidimensional spatial decision-making process: Local shared values in action. Lecture Notes in Computer Science, 7334:54–70, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31075-1_5/.

Campobasso:2012:PSF

[631] Francesco Campobasso and Annarita Fanizzi. A proposal for a stepwise fuzzy regression: An application to the Italian University system. Lecture Notes in Computer Science, 7334:71–87, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31075-1_6/.

Perchinunno:2012:CAS

[632] Paola Perchinunno and Dario Antonio Schirone. Cluster analysis for strategic management: a case study of IKEA. Lecture Notes in Computer Science, 7334:88–101, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31075-1_7/.

Montrone:2012:CLD

[633] Silvestro Montrone and Paola Perchinunno. Clustering for the localization of degraded urban areas. Lecture Notes in Computer Science, 7334: 102–115, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31075-1_8/.

Balena:2012:BAE

[634] Pasquale Balena, Giovanna Mangialardi, and Carmelo Maria Torre. A BEP analysis of energy supply for sustainable urban microgrids. Lecture Notes in Computer Science, 7334:116–127, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31075-1_9/.

Carbonara:2012:EIW

[635] Sebastiano Carbonara. The effect of infrastructural works on urban property values: The asse attrezzato in Pescara, Italy. Lecture Notes in Computer Science, 7334:128–143, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31075-1_10/.

Selicato:2012:PIM

[636] Marco Selicato, Carmelo Maria Torre, and Giovanni La Trofa. Prospect of integrate monitoring: a multidimensional approach. Lecture Notes in Computer Science, 7334:144–156, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31075-1_11/.

Girard:2012:UAM

[637] Luigi Fusco Girard and Carmelo Maria Torre. The use of Ahp in a multiactor evaluation for urban development programs: a case study. Lecture Notes in Computer Science, 7334:157–167, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31075-1_12/.

Cerreta:2012:AUT

[638] Maria Cerreta and Pasquale De Toro. Assessing urban transformations: a SDSS for the master plan of Castel Capuano, Naples. Lecture Notes in Computer Science, 7334:168–180, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31075-1_13/.

Carvalho:2012:CCP

[639] Luis Paulo da Silva Carvalho and Paulo Caetano da Silva. Computational context to promote geographic information systems toward human-centric perspectives. Lecture Notes in Computer Science, 7334: 181–193, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31075-1_14/.

Ghandehari:2012:VBC

[640] Mehran Ghandehari and Farid Karimipour. Voronoi-based curve reconstruction: Issues and solutions. Lecture Notes in Computer Science, 7334: 194–207, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31075-1_15/.

Gonschorek:2012:GGC

[641] Julia Gonschorek and Lucia Tyrallová. Geovisualization and geostatistics: a concept for the numerical and visual analysis of geographic mass data. Lecture Notes in Computer Science, 7334:208–219, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31075-1_16/.

Tyrallova:2012:SEA

[642] Lucia Tyrallová and Julia Gonschorek. Spatio-explorative analysis and its benefits for a GIS-integrated automated feature identification. Lecture Notes in Computer Science, 7334:220–233, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31075-1_17/.

Fiorese:2012:PSP

[643] Adriano Fiorese, Paulo Simões, and Fernando Boavida. Peer selection in P2P service overlays using geographical location criteria. Lecture Notes in Computer Science, 7334:234–248, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31075-1_18/.

OKelly:2012:MSI

[644] Morton E. O’Kelly. Models for spatial interaction data: Computation and interpretation of accessibility. Lecture Notes in Computer Science, 7334: 249–262, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31075-1_19/.

Lederer:2012:SMH

[645] Daniel Lederer. Am I safe in my home? Fear of crime analyzed with spatial statistics methods in a central European city. Lecture Notes in Computer Science, 7334:263–274, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31075-1_20/.

Anonymous:2012:FMbg

[646] Anonymous. Front matter. Lecture Notes in Computer Science, 7334: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31075-1/1.

Rocha:2012:IRB

[647] Humberto Rocha, Joana M. Dias, Brigida C. Ferreira, and Maria do Carmo Lopes. Incorporating radial basis functions in pattern search methods: Application to beam angle optimization in radiotherapy treatment planning. Lecture Notes in Computer Science, 7335: 1–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31137-6_1/.

Teixeira:2012:CMT

[648] Ana Paula Teixeira and Regina Almeida. On the complexity of a Mehrotra-type predictor-corrector algorithm. Lecture Notes in Computer Science, 7335:17–29, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31137-6_2/.

CostaGomes:2012:DWB

[649] Tiago Costa Gomes, Filipe Pereira e Alvelos, and Maria Sameiro Carvalho. Design of wood biomass supply chains. Lecture Notes in Computer Science, 7335:30–44, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31137-6_3/.

Hendrix:2012:SSP

[650] Eligius M. T. Hendrix, Rene Haijema, Roberto Rossi, and Karin G. J. Pauls-Worm. On solving a stochastic programming model for perishable inventory control. Lecture Notes in Computer Science, 7335:45– 56, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31137-6_4/.

Rocha:2012:AFS

[651] Ana Maria A. C. Rocha, M. Fernanda P. Costa, and Edite M. G. P. Fernandes. An artificial fish swarm filter-based method for constrained global optimization. Lecture Notes in Computer Science, 7335:57–71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31137-6_5/.

Azad:2012:SMK

[652] Md. Abul Kalam Azad, Ana Maria A. C. Rocha, and Edite M. G. P. Fernandes. Solving multidimensional 0–1 knapsack problem with an artificial fish swarm algorithm. Lecture Notes in Computer Science, 7335: 72–86, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31137-6_6/.

Gupta:2012:OMC

[653] Pankaj Gupta, Shilpi Verma, and Mukesh Kumar Mehlawat. Optimization model of COTS selection based on cohesion and coupling for modular software systems under multiple applications environment. Lecture Notes in Computer Science, 7335:87–102, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31137-6_7/.

Fernandes:2012:DFF

[654] Florbela P. Fernandes, M. Fernanda P. Costa, and Edite M. G. P. Fernandes. A derivative-free filter driven multistart technique for global optimization. Lecture Notes in Computer Science, 7335:103–118, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31137-6_8/.

Berenguel:2012:LBU

[655] José L. Berenguel, Leocadio G. Casado, I. García, Eligius M. T. Hendrix, and F. Messine. On lower bounds using additively separable terms in interval B&B. Lecture Notes in Computer Science, 7335:119–132, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31137-6_9/.

Figueiredo:2012:GAJ

[656] José Figueiredo, José A. Oliveira, Luis Dias, and Guilherme A. B. Pereira. A genetic algorithm for the job shop on an ASRS warehouse. Lecture Notes in Computer Science, 7335:133–146, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31137-6_10/.

Ferreira:2012:SPM

[657] Ana C. M. Ferreira, Ana Maria A. C. Rocha, and Senhorinha F. C. F. Teixeira. On solving the profit maximization of small cogeneration systems. Lecture Notes in Computer Science, 7335:147–158, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31137-6_11/.

Hendrix:2012:GOS

[658] Eligius M. T. Hendrix, Leocadio G. Casado, and Paula Amaral. Global optimization simplex bisection revisited based on considerations by Reiner Horst. Lecture Notes in Computer Science, 7335:159–173, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31137-6_12/.

Ferreira:2012:AVA

[659] Manuel Ferreira and José Carlos Teixeira. Application of variance analysis to the combustion of residual oils. Lecture Notes in Computer Science, 7335:174–186, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-31137-6_13/.

Geraldes:2012:WDP

[660] Carla A. S. Geraldes, Maria Sameiro Carvalho, and Guilherme A. B. Pereira. Warehouse design and planning: a mathematical programming approach. Lecture Notes in Computer Science, 7335:187–201, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31137-6_14/.

Teixeira:2012:ACT

[661] José Carlos Teixeira, Ricardo Lomba, Senhorinha F. C. F. Teixeira, and Pedro Lobarinhas. Application of CFD tools to optimize natural building ventilation design. Lecture Notes in Computer Science, 7335: 202–216, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31137-6_15/.

Lee:2012:MIU

[662] Junghoon Lee, Gyung-Leen Park, Min-Jae Kang, Ho-Young Kwak, Sang Joon Lee, and Jikwang Han. Middleware integration for ubiquitous sensor networks in agriculture. Lecture Notes in Computer Science, 7335: 217–226, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31137-6_16/.

Song:2012:UPB

[663] Hokwon Song, Changwoo Min, Jeehong Kim, and Young Ik Eom. Usage pattern-based prefetching: Quick application launch on mobile devices. Lecture Notes in Computer Science, 7335:227–237, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31137-6_17/.

Bae:2012:EEI

[664] Sunwook Bae, Hokwon Song, Changwoo Min, Jeehong Kim, and Young Ik Eom. EIMOS: Enhancing interactivity in mobile operating systems. Lecture Notes in Computer Science, 7335:238–247, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31137-6_18/.

Kim:2012:DMH

[665] Young-Hyuk Kim, Il-Kown Lim, Jae-Pil Lee, Jae-Gwang Lee, and Jae-Kwang Lee. Development of mobile hybrid MedIntegraWeb app for interoperation between u-RPMS and HIS. Lecture Notes in Computer Science, 7335:248–258, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31137-6_19/.

Le:2012:DLM

[666] Duc Tai Le, Thang Le Duc, and Hyunseung Choo. A distributed lifetimemaximizing scheme for connected target coverage in WSNs. Lecture Notes in Computer Science, 7335:259–271, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31137-6_20/.

Anonymous:2012:FMbh

[667] Anonymous. Front matter. Lecture Notes in Computer Science, 7335: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31137-6/1.

Silvestre:2012:MRT

[668] Eduardo Augusto Silvestre and Michel dos Santos Soares. Modeling road traffic signals control using UML and the MARTE profile. Lecture Notes in Computer Science, 7336:1–15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31128-4_1/.

Soares:2012:ATD

[669] Michel dos Santos Soares and Daniel Souza Cioquetta. Analysis of techniques for documenting user requirements. Lecture Notes in Computer Science, 7336:16–28, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31128-4_2/.

Coscia:2012:PWS

[670] José Luis Ordiales Coscia, Marco Crasso, Cristian Mateos, and Alejandro Zunino. Predicting Web service maintainability via object-oriented metrics: a statistics-based approach. Lecture Notes in Computer Science, 7336:29–39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31128-4_3/.

Biehl:2012:EAV

[671] Matthias Biehl. Early automated verification of tool chain design. Lecture Notes in Computer Science, 7336:40–50, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31128-4_4/.

Batista:2012:UUS

[672] Vitor A. Batista, Daniela C. C. Peixoto, Wilson Pádua, and Clarindo Isaías P. S. Pádua. Using UML stereotypes to support the requirement engineering: a case study. Lecture Notes in Computer Science, 7336: 51–66, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31128-4_5/.

doNascimento:2012:IBR

[673] Gleison S. do Nascimento, Cirano Iochpe, Lucinéia Thom, and André C. Kalsing. Identifying business rules to legacy systems reengineering based on BPM and SOA. Lecture Notes in Computer Science, 7336:67–82, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31128-4_6/.

El-Zawawy:2012:AAC

[674] Mohamed A. El-Zawawy. Abstraction analysis and certified flow and context sensitive points-to relation for distributed programs. Lecture Notes in Computer Science, 7336:83–99, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31128-4_7/.

Zhang:2012:AMU

[675] Yan Zhang, Yi Liu, Zhiyi Ma, Xuying Zhao, Xiaokun Zhang, and Tian Zhang. An approach to measure understandability of extended UML based on metamodel. Lecture Notes in Computer Science, 7336: 100–115, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31128-4_8/.

Aguilar:2012:DDA

[676] José Alfonso Aguilar, Irene Garrigós, Jose-Norberto Mazón, and Anibal Zaldívar. Dealing with dependencies among functional and non-functional requirements for impact analysis in Web engineering. Lecture Notes in Computer Science, 7336:116–131, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31128-4_9/.

dosAnjos:2012:AMM

[677] Eudisley Gomes dos Anjos, Ruan Delgado Gomes, and Mário Zenha-Rela. Assessing maintainability metrics in software architectures using COSMIC and UML. Lecture Notes in Computer Science, 7336: 132–146, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31128-4_10/.

Pandey:2012:PDS

[678] Kusum Lata Pandey, Suneeta Agarwal, Sanjay Misra, and Rajesh Prasad. Plagiarism detection in software using efficient string matching. Lecture Notes in Computer Science, 7336:147–156, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31128-4_11/.

Shukla:2012:DSM

[679] Ruchi Shukla, Mukul Shukla, A. K. Misra, T. Marwala, and W. A. Clarke. Dynamic software maintenance effort estimation modeling using neural network, rule engine and multi-regression approach. Lecture Notes in Computer Science, 7336:157–169, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31128-4_12/.

Decker:2012:NMM

[680] Hendrik Decker. New measures for maintaining the quality of databases. Lecture Notes in Computer Science, 7336:170–185, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.com/chapter/10.1007/978-3-642-31128-4_13/.

Aydin:2012:NWD

[681] Ali Orhan Aydin. A new way to determine external quality of ERP software. Lecture Notes in Computer Science, 7336:186–201, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31128-4_14/.

Cunha:2012:TCS

[682] Jácome Cunha, João P. Fernandes, Hugo Ribeiro, and João Saraiva. Towards a catalog of spreadsheet smells. Lecture Notes in Computer Science, 7336:202–216, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31128-4_15/.

Martins:2012:PAM

[683] Pedro Martins, Paulo Lopes, João P. Fernandes, João Saraiva, and João M. P. Cardoso. Program and aspect metrics for MATLAB. Lecture Notes in Computer Science, 7336:217–233, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31128-4_16/.

Misra:2012:SCC

[684] Sanjay Misra, Murat Koyuncu, Marco Crasso, Cristian Mateos, and Alejandro Zunino. A suite of cognitive complexity metrics. Lecture Notes in Computer Science, 7336:234–247, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31128-4_17/.

Adewumi:2012:CMC

[685] Adewole Adewumi, Sanjay Misra, and Nicholas Ikhu-Omoregbe. Complexity metrics for cascading style sheets. Lecture Notes in Computer Science, 7336:248–257, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31128-4_18/.

Khan:2012:SRI

[686] Yasser A. Khan, Mahmoud O. Elish, and Mohamed El-Attar. A systematic review on the impact of CK metrics on the functional correctness of object-oriented classes. Lecture Notes in Computer Science, 7336: 258–273, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31128-4_19/.

Gregio:2012:PMA

[687] André Ricardo Abed Grégio, Vitor Monte Afonso, and Dario Simões Fernandes Filho. Pinpointing malicious activities through network and system-level malware execution behavior. Lecture Notes in Computer Science, 7336:274–285, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31128-4_20/.

Anonymous:2012:FMbi

[688] Anonymous. Front matter. Lecture Notes in Computer Science, 7336: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31128-4/1.

Thalheim:2012:SSP

[689] Bernhard Thalheim. Syntax, semantics and pragmatics of conceptual modelling. Lecture Notes in Computer Science, 7337:1–10, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31178-9_1/.

Stuckenschmidt:2012:MDA

[690] Heiner Stuckenschmidt and Cäcilia Zirn. Multi-dimensional analysis of political documents. Lecture Notes in Computer Science, 7337:11–22, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31178-9_2/.

Lappas:2012:FRM

[691] Theodoros Lappas. Fake reviews: The malicious perspective. Lecture Notes in Computer Science, 7337:23–34, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31178-9_3/.

Klenner:2012:PPV

[692] Manfred Klenner and Stefanos Petrakis. Polarity preference of verbs: What could verbs reveal about the polarity of their objects? Lecture Notes in Computer Science, 7337:35–46, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31178-9_4/.

Schlaf:2012:LQP

[693] Antje Schlaf, Amit Kirschenbaum, Robert Remus, and Thomas Efer. Labeling queries for a people search engine. Lecture Notes in Computer Science, 7337:47–57, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31178-9_5/.

Dwarakanath:2012:LGT

[694] Anurag Dwarakanath and Shubhashis Sengupta. Litmus: Generation of test cases from functional requirements in natural language. Lecture Notes in Computer Science, 7337:58–69, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31178-9_6/.

Silveira:2012:EMD

[695] Sara Botelho Silveira and António Branco. Extracting multi-document summaries with a double clustering approach. Lecture Notes in Computer Science, 7337:70–81, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31178-9_7/.

Kontonasios:2012:DMT

[696] Georgios Kontonasios, Ioannis Korkontzelos, and Sophia Ananiadou. Developing multilingual text mining workflows in UIMA and U-compare. Lecture Notes in Computer Science, 7337:82–93, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31178-9_8/.

Perea-Ortega:2012:GEQ

[697] José M. Perea-Ortega and L. Alfonso Ureña-López. Geographic expansion of queries to improve the geographic information retrieval task. Lecture Notes in Computer Science, 7337:94–103, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31178-9_9/.

Chali:2012:LGD

[698] Yllias Chali, Sadid A. Hasan, and Kaisar Imam. Learning good decompositions of complex questions. Lecture Notes in Computer Science, 7337: 104–115, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31178-9_10/.

Nguyen:2012:SSL

[699] Le Minh Nguyen and Akira Shimazu. A semi supervised learning model for mapping sentences to logical form with ambiguous supervision. Lecture Notes in Computer Science, 7337:116–127, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31178-9_11/.

Leveling:2012:ESR

[700] Johannes Leveling. On the effect of stopword removal for SMS-based FAQ retrieval. Lecture Notes in Computer Science, 7337:128–139, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31178-9_12/.

Boston:2012:WDS

[701] Christopher Boston, Sandra Carberry, and Hui Fang. Wikimantic: Disambiguation for short queries. Lecture Notes in Computer Science, 7337: 140–151, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31178-9_13/.

Ogrodniczuk:2012:PLP

[702] Maciej Ogrodniczuk and Adam Przepiórkowski. Polish language processing chains for multilingual information systems. Lecture Notes in Computer Science, 7337:152–157, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31178-9_14/.

Kysenko:2012:GAN

[703] Volodymyr Kysenko, Karl Rupp, Oleksandr Marchenko, and Siegfried Selberherr. GPU-accelerated non-negative matrix factorization for text mining. Lecture Notes in Computer Science, 7337:158–163, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31178-9_15/.

Giordani:2012:GSQ

[704] Alessandra Giordani and Alessandro Moschitti. Generating SQL queries using natural language syntactic dependencies and metadata. Lecture Notes in Computer Science, 7337:164–170, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31178-9_16/.

Buitinck:2012:TSN

[705] Lars Buitinck and Maarten Marx. Two-stage named-entity recognition using averaged perceptrons. Lecture Notes in Computer Science, 7337: 171–176, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31178-9_17/.

Bloom:2012:UNL

[706] Niels Bloom. Using natural language processing to improve document categorization with associative networks. Lecture Notes in Computer Science, 7337:177–182, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31178-9_18/.

Jebari:2012:MML

[707] Chaker Jebari. MLICC: a multi-label and incremental centroid-based classification of Web pages by genre. Lecture Notes in Computer Science, 7337: 183–190, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31178-9_19/.

Kramer:2012:CIG

[708] Gerwin Kramer, Gosse Bouma, Dennis Hendriksen, and Mathijs Homminga. Classifying image galleries into a taxonomy using metadata and Wikipedia. Lecture Notes in Computer Science, 7337:191–196, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31178-9_20/.

Anonymous:2012:FMbj

[709] Anonymous. Front matter. Lecture Notes in Computer Science, 7337: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31178-9/1.

Maier:2012:NOD

[710] David Maier, V. M. Megler, António M. Baptista, Alex Jaramillo, and Charles Seaton. Navigating oceans of data. Lecture Notes in Computer Science, 7338:1–19, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31235-9_1/.

Patroumpas:2012:PRM

[711] Kostas Patroumpas, Marios Papamichalis, and Timos Sellis. Probabilistic range monitoring of streaming uncertain positions in GeoSocial networks. Lecture Notes in Computer Science, 7338:20–37, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31235-9_2/.

Bernecker:2012:PFP

[712] Thomas Bernecker, Hans-Peter Kriegel, Matthias Renz, Florian Verhein, and Andreas Züfle. Probabilistic frequent pattern growth for itemset mining in uncertain databases. Lecture Notes in Computer Science, 7338: 38–55, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31235-9_3/.

Xie:2012:ETQ

[713] Xike Xie, Reynold Cheng, and Man Lung Yiu. Evaluating trajectory queries over imprecise location data. Lecture Notes in Computer Science, 7338:56–74, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31235-9_4/.

Dai:2012:ERQ

[714] Dongbo Dai, Jiang Xie, Huiran Zhang, and Jiaqi Dong. Efficient range queries over uncertain strings. Lecture Notes in Computer Science, 7338: 75–95, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31235-9_5/.

Hubig:2012:CPS

[715] Nina Hubig, Andreas Züfle, Tobias Emrich, Mario A. Nascimento, and Matthias Renz. Continuous probabilistic sum queries in wireless sensor networks with ranges. Lecture Notes in Computer Science, 7338: 96–105, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31235-9_6/.

Dannecker:2012:PMC

[716] Lars Dannecker, Matthias Böehm, Wolfgang Lehner, and Gregor Hackenbroich. Partitioning and multi-core parallelization of multi-equation forecast models. Lecture Notes in Computer Science, 7338:106–123, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31235-9_7/.

Lu:2012:IGA

[717] Mian Lu, Yuwei Tan, Jiuxin Zhao, Ge Bai, and Qiong Luo. Integrating GPU-accelerated sequence alignment and SNP detection for genome resequencing analysis. Lecture Notes in Computer Science, 7338:124– 140, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31235-9_8/.

Vlachou:2012:DRS

[718] Akrivi Vlachou, Christos Doulkeridis, and Maria Halkidi. Discovering representative skyline points over distributed data. Lecture Notes in Computer Science, 7338:141–158, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31235-9_9/.

Dobos:2012:SIP

[719] László Dobos, Tamás Budavári, Nolan Li, Alexander S. Szalay, and István Csabai. SkyQuery: An implementation of a parallel probabilistic join engine for cross-identification of multiple astronomical databases. Lecture Notes in Computer Science, 7338:159–167, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31235-9_10/.

Dahimene:2012:EFM

[720] Ryadh Dahimene, Cedric Du Mouza, and Michel Scholl. Efficient filtering in micro-blogging systems: We won’t get flooded again. Lecture Notes in Computer Science, 7338:168–176, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31235-9_11/.

Koschmieder:2012:RPQ

[721] André Koschmieder and Ulf Leser. Regular path queries on large graphs. Lecture Notes in Computer Science, 7338:177–194, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31235-9_12/.

Lu:2012:SCI

[722] Xuesong Lu and Stéphane Bressan. Sampling connected induced subgraphs uniformly at random. Lecture Notes in Computer Science, 7338:

195–212, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31235-9_13/.

Valari:2012:DTD

[723] Elena Valari, Maria Kontaki, and Apostolos N. Papadopoulos. Discovery of top-k dense subgraphs in dynamic graph collections. Lecture Notes in Computer Science, 7338:213–230, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31235-9_14/.

Yu:2012:EEP

[724] Weiren Yu, Jiajin Le, Xuemin Lin, and Wenjie Zhang. On the efficiency of estimating penetrating rank on large graphs. Lecture Notes in Computer Science, 7338:231–249, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31235-9_15/.

Zhang:2012:TEJ

[725] Xiaofei Zhang, Lei Chen, and Min Wang. Towards efficient join processing over large RDF graph using MapReduce. Lecture Notes in Computer Science, 7338:250–259, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31235-9_16/.

Ioannidis:2012:PDI

[726] Yannis Ioannidis. Panel on “data infrastructures and data management research: Close relatives or total strangers?”. Lecture Notes in Computer Science, 7338:260–261, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/accesspage/
chapter/10.1007/978-3-642-31235-9_17.

Fenz:2012:ESS

[727] Dandy Fenz, Dustin Lange, Astrid Rheinländer, Felix Naumann, and Ulf Leser. Efficient similarity search in very large string sets. Lecture Notes in Computer Science, 7338:262–279, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31235-9_18/.

Gunnemann:2012:SCN

[728] Stephan Günnemann, Brigitte Boden, and Thomas Seidl. Substructure clustering: a novel mining paradigm for arbitrary data types. Lecture Notes in Computer Science, 7338:280–297, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31235-9_19/.

Kranen:2012:BAA

[729] Philipp Kranen, Marwan Hassani, and Thomas Seidl. BT* — an advanced algorithm for anytime classification. Lecture Notes in Computer Science, 7338:298–315, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-31235-9_20/.

Anonymous:2012:FMbk

[730] Anonymous. Front matter. Lecture Notes in Computer Science, 7338: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31235-9/1.

Grzegorzek:2012:BVS

[731] Marcin Grzegorzek and Paula Lubina. Blood vessel segmentation in HRT images for glaucoma early detection. Lecture Notes in Computer Science, 7339:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31196-3_1/.

Koprowski:2012:EQL

[732] Robert Koprowski, Slawomir Teper, Edward Wylegala, and Zygmunt Wróbel. Enhancing the quality of layer detection in tomographic images of the eye. Lecture Notes in Computer Science, 7339:13–23, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31196-3_2/.

Pieciak:2012:SLV

[733] Tomasz Pieciak. Segmentation of the left ventricle using active contour method with gradient vector flow forces in short-axis MRI. Lecture Notes in Computer Science, 7339:24–35, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31196-3_3/.

Szwarc:2012:WMS

[734] Paweł Szwarc, Jacek Kawa, and Ewa Pietka. White matter segmentation from MR images in subjects with brain tumours. Lecture Notes in Computer Science, 7339:36–46, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31196-3_4/.

Badura:2012:FLT

[735] Paweł Badura and Ewa Pietka. 3D fuzzy liver tumor segmentation. Lecture Notes in Computer Science, 7339:47–57, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31196-3_5/.

Czajkowska:2012:KFC

[736] Joanna Czajkowska, Monika Bugdol, and Ewa Pietka. Kernelized fuzzy C-means method and Gaussian mixture model in unsupervised cascade clustering. Lecture Notes in Computer Science, 7339:58–66, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31196-3_6/.

Jaworek-Korjakowska:2012:ADM

[737] Joanna Jaworek-Korjakowska. Automatic detection of melanomas: An application based on the ABCD criteria. Lecture Notes in Computer Science, 7339:67–76, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31196-3_7/.

Jasionowska:2012:SDM

[738] Magdalena Jasionowska and Artur Przelaskowski. Subtle directional mammographic findings in multiscale domain. Lecture Notes in Computer Science, 7339:77–84, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-31196-3_8/.

Filipczuk:2012:AND

[739] Paweł Filipczuk, Weronika Wojtak, and Andrzej Obuchowicz. Automatic nuclei detection on cytological images using the firefly optimization algorithm. Lecture Notes in Computer Science, 7339:85–92, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31196-3_9/.

Zarychta:2012:CAL

[740] Piotr Zarychta, Henryk Konik, and Anna Zarychta-Bargieła. Computer assisted location of the lower limb mechanical axis. Lecture Notes in Computer Science, 7339:93–100, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31196-3_10/.

Ostrek:2012:AES

[741] Grzegorz Ostrek and Artur Przelaskowski. Automatic early stroke recognition algorithm in CT images. Lecture Notes in Computer Science, 7339: 101–109, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31196-3_11/.

Majchrzak:2012:DSA

[742] Ewa Majchrzak, Mirosław Dziewoński, Mariusz Nowak, Marek Kawecki, and Michał Bachorz. The design of a system for assisting burn and chronic wound diagnosis. Lecture Notes in Computer Science, 7339: 110–117, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31196-3_12/.

Popielski:2012:FDH

[743] Paweł Popielski and Zygmunt Wróbel. The feature detection on the homogeneous surfaces with projected pattern. Lecture Notes in Computer Science, 7339:118–128, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31196-3_13/.

Przelaskowski:2012:SCC

[744] Artur Przelaskowski and Rafal Jozwiak. Sensed compression with cosine and noiselet measurements for medical imaging. Lecture Notes in Computer Science, 7339:129–142, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31196-3_14/.

Rudzki:2012:IPA

[745] Marcin Rudzki, Monika Bugdol, and Tomasz Ponikiewski. An image processing approach to determination of steel fibers orientation in reinforced concrete. Lecture Notes in Computer Science, 7339:143–150, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31196-3_15/.

Kulikowski:2012:MBS

[746] Juliusz L. Kulikowski. Multichannel biomedical signals analysis based on a split-and-collect approach. Lecture Notes in Computer Science, 7339: 151–160, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31196-3_16/.

Babusiak:2012:DAE

[747] Branko Babusiak and Michal Gala. Detection of abnormalities in ECG. Lecture Notes in Computer Science, 7339:161–171, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31196-3_17/.

Przywara-Chowaniec:2012:RPU

[748] Brygida Przywara-Chowaniec, Lech Poloński, Maciej Gawlikowski, and Tadeusz Pustelny. Research into the possibility to use impedance rheocardiography in a non-invasive assessment of haemodynamic condition of patients with heart diseases. Lecture Notes in Computer Science, 7339: 172–182, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31196-3_18/.

Podsiadly-Marczykowska:2012:OEM

[749] Teresa Podsiadly-Marczykowska and Hanna Goszczynska. Ontology of EEG mapping — preliminary research. Lecture Notes in Computer Science, 7339:183–198, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31196-3_19/.

Kurzynski:2012:CSB

[750] Marek Kurzynski and Andrzej Wolczowski. Control system of bioprosthetic hand based on advanced analysis of biosignals and feedback from the prosthesis sensors. Lecture Notes in Computer Science, 7339: 199–208, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31196-3_20/.

Anonymous:2012:FMbl

[751] Anonymous. Front matter. Lecture Notes in Computer Science, 7339: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31196-3/1.

Bentabet:2012:BIM

[752] Layachi Bentabet and Hui Zhang. Bayesian image matting using infrared and color cues. Lecture Notes in Computer Science, 7340:1–8, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31254-0_1/.

LeMoan:2012:SPD

[753] Steven Le Moan, Ferdinand Deger, Alamin Mansouri, Yvon Voisin, and Jon Y. Hardeberg. Salient pixels and dimensionality reduction for display of multi/hyperspectral images. Lecture Notes in Computer Science, 7340:9–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31254-0_2/.

Bekkari:2012:SHF

[754] Aissam Bekkari, Soufiane Idbraim, Azeddine Elhassouny, Driss Mammass, and Mostafa El yassa. SVM and Haralick features for classification of high resolution satellite images from urban areas. Lecture Notes in Computer Science, 7340:17–26, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31254-0_3/.

Maczkowski:2012:DAE

[755] Grzegorz Maczkowski, Robert Sitnik, and Jakub Krzesłowski. Data acquisition enhancement in shape and multispectral color measurements of 3D objects. Lecture Notes in Computer Science, 7340:27–35, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31254-0_4/.

ElMaliani:2012:MMA

[756] Ahmed Drissi El Maliani, Mohammed El Hassouni, Yannick Berthoumieu, and Driss Aboutajdine. Multi-model approach for multicomponent texture classification. Lecture Notes in Computer Science, 7340:36–44, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31254-0_5/.

Shrestha:2012:SMI

[757] Raju Shrestha and Jon Yngve Hardeberg. Simultaneous multispectral imaging and illuminant estimation using a stereo camera. Lecture Notes in Computer Science, 7340:45–55, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31254-0_6/.

Elhassouny:2012:MFC

[758] Azeddine Elhassouny, Soufiane Idbraim, Aissam Bekkarri, Driss Mammass, and Danielle Ducrot. Multisource fusion/classification using ICM and DSmT with new decision rule. Lecture Notes in Computer Science, 7340:56–64, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31254-0_7/.

Mahani:2012:TEP

[759] Zouhir Mahani, Jalal Zahid, Sahar Saoud, Mohammed El Rhabi, and Abdelilah Hakim. Text enhancement by PDE’s based methods. Lecture Notes in Computer Science, 7340:65–76, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31254-0_8/.

Badri:2012:KBL

[760] Hicham Badri, Mohammed El Hassouni, and Driss Aboutajdine. Kernelbased Laplacian smoothing method for 3D mesh denoising. Lecture Notes in Computer Science, 7340:77–84, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31254-0_9/.

Said:2012:ERT

[761] Yahia Said, Taoufik Saidani, Fethi Smach, Mohamed Atri, and Hichem Snoussi. Embedded real-time video processing system on FPGA. Lecture Notes in Computer Science, 7340:85–92, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31254-0_10/.

Khare:2012:EPI

[762] Ashish Khare, Richa Srivastava, and Rajiv Singh. Edge preserving image fusion based on contourlet transform. Lecture Notes in Computer Science, 7340:93–102, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31254-0_11/.

Qaffou:2012:SVO

[763] Issam Qaffou, Mohamed Sadgal, and Aziz Elfazziki. Selecting vision operators and fixing their optimal parameters values using reinforcement learning. Lecture Notes in Computer Science, 7340:103–112, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31254-0_12/.

Nafchi:2012:PCB

[764] Hossein Ziaei Nafchi and Hamidreza Rashidy Kanan. A phase congruency based document binarization. Lecture Notes in Computer Science, 7340: 113–121, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31254-0_13/.

Skoudarli:2012:PHA

[765] Abdellah Skoudarli, Mokhtar Nibouche, and Amina Serir. Porting a H264/ AVC adaptive in loop deblocking filter to a TI DM6437EVM DSP. Lecture Notes in Computer Science, 7340:122–130, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31254-0_14/.

Falek:2012:MAC

[766] Leila Falek, Hocine Teffahi, and Amar Djeradi. Methodology for acoustic characterization of a labial constraint in speech production. Lecture Notes in Computer Science, 7340:131–141, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31254-0_15/.

Tayebi:2012:POR

[767] Mohamed Tayebi and Mrahi Bouziani. Performance of OFDM in radio mobile channel. Lecture Notes in Computer Science, 7340:142–148, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31254-0_16/.

Chaibi:2012:SCC

[768] H. Chaibi, R. Saadane, My A. Faqihi, and M. Belkasmi. Spatial correlation characterization for UWB indoor channel based on measurements. Lecture Notes in Computer Science, 7340:149–156, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31254-0_17/.

Hosseini:2012:NBS

[769] Shahram Hosseini, Yannick Deville, Sonia El Amine, and Hicham Saylani. Nonlinear blind source separation applied to a simple bijective model. Lecture Notes in Computer Science, 7340:157–165, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31254-0_18/.

Laasri:2012:SSD

[770] El Hassan Ait Laasri, Es-Saïd Akhouayri, Dris Agliz, and Abderrahman Atmani. Seismic signal discrimination between earthquakes and quarry blasts using fuzzy logic approach. Lecture Notes in Computer Science, 7340:166–174, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-31254-0_19/.

Mohammadi:2012:UWB

[771] Zakaria Mohammadi, Rachid Saadane, and Driss Aboutajdine. Ultra wide-band channel characterization using generalized Gamma distributions. Lecture Notes in Computer Science, 7340:175–182, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31254-0_20/.

Anonymous:2012:FMbm

[772] Anonymous. Front matter. Lecture Notes in Computer Science, 7340: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31254-0/1.

Hao:2012:SAM

[773] Feng Hao and Dylan Clarke. Security analysis of a multi-factor authenticated key exchange protocol. Lecture Notes in Computer Science, 7341:1–11, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31284-7_1/.

Nguyen:2012:BAC

[774] Vu Duc Nguyen, Yang-Wai Chow, and Willy Susilo. Breaking an animated CAPTCHA scheme. Lecture Notes in Computer Science, 7341: 12–29, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31284-7_2/.

Ben-David:2012:COM

[775] Assaf Ben-David, Omer Berkman, Yossi Matias, Sarvar Patel, Cem Paya, and Moti Yung. Contextual OTP: Mitigating emerging man-in-the-middle attacks with wireless hardware tokens. Lecture Notes in Computer Science, 7341:30–47, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-31284-7_3/.

Zhang:2012:RUR

[776] Nan Zhang, Jingqiang Lin, Jiwu Jing, and Neng Gao. RIKE: Using revocable identities to support key escrow in PKIs. Lecture Notes in Computer Science, 7341:48–65, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31284-7_4/.

Muller:2012:T

[777] Tilo Müller, Benjamin Taubmann, and Felix C. Freiling. TreVisor. Lecture Notes in Computer Science, 7341:66–83, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31284-7_5/.

Alomair:2012:AEH

[778] Basel Alomair. Authenticated encryption: How reordering can impact performance. Lecture Notes in Computer Science, 7341:84–99, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31284-7_6/.

Zhang:2012:LDC

[779] Haibin Zhang. Length-doubling ciphers and tweakable ciphers. Lecture Notes in Computer Science, 7341:100–116, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31284-7_7/.

Zhang:2012:EHO

[780] Wentao Zhang, Bozhan Su, Wenling Wu, Dengguo Feng, and Chuankun Wu. Extending higher-order integral: An efficient unified algorithm of constructing integral distinguishers for block ciphers. Lecture Notes in Computer Science, 7341:117–134, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31284-7_8/.

Fujioka:2012:SEP

[781] Atsushi Fujioka, Taiichi Saito, and Keita Xagawa. Security enhancements by OR-proof in identity-based identification. Lecture Notes in Computer Science, 7341:135–152, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31284-7_9/.

Chen:2012:IBE

[782] Yu Chen, Zongyang Zhang, Dongdai Lin, and Zhenfu Cao. Identity-based extractable hash proofs and their applications. Lecture Notes in Computer Science, 7341:153–170, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31284-7_10/.

Samelin:2012:SST

[783] Kai Samelin, Henrich C. Pöhls, Arne Bilzhause, Joachim Posegga, and Hermann de Meer. On structural signatures for tree data structures. Lecture Notes in Computer Science, 7341:171–187, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31284-7_11/.

Xie:2012:IPL

[784] Xiang Xie, Rui Xue, and Rui Zhang. Inner-product lossy trapdoor functions and applications. Lecture Notes in Computer Science, 7341: 188–205, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31284-7_12/.

Arriaga:2012:JSS

[785] Afonso Arriaga, Manuel Barbosa, and Pooya Farshim. On the joint security of signature and encryption schemes under randomness reuse: Efficiency and security amplification. Lecture Notes in Computer Science, 7341:206–223, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-31284-7_13/.

Lipmaa:2012:SAE

[786] Helger Lipmaa. Secure accumulators from Euclidean rings without trusted setup. Lecture Notes in Computer Science, 7341:224–240, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31284-7_14/.

Liu:2012:LFA

[787] Zhiqiang Liu, Dawu Gu, Ya Liu, and Wei Li. Linear fault analysis of block ciphers. Lecture Notes in Computer Science, 7341:241–256, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31284-7_15/.

Asano:2012:CBK

[788] Yuki Asano, Shingo Yanagihara, and Tetsu Iwata. Cryptanalysis of 256bit key HyRAL via equivalent keys. Lecture Notes in Computer Science, 7341:257–274, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-31284-7_16/.

Sasaki:2012:DBT

[789] Yu Sasaki and Lei Wang. Distinguishers beyond three rounds of the RIPEMD-128/-160 compression functions. Lecture Notes in Computer Science, 7341:275–292, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31284-7_17/.

Zhang:2012:ZVP

[790] Fangguo Zhang, Qiping Lin, and Shengli Liu. Zero-value point attacks on Kummer-based cryptosystem. Lecture Notes in Computer Science, 7341: 293–310, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31284-7_18/.

Piret:2012:PBC

[791] Gilles Piret, Thomas Roche, and Claude Carlet. PICARO — a block cipher allowing efficient higher-order side-channel resistance. Lecture Notes in Computer Science, 7341:311–328, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31284-7_19/.

Ye:2012:WCP

[792] Xin Ye and Thomas Eisenbarth. Wide collisions in practice. Lecture Notes in Computer Science, 7341:329–343, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31284-7_20/.

Anonymous:2012:FMbn

[793] Anonymous. Front matter. Lecture Notes in Computer Science, 7341: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31284-7/1.

Barthe:2012:PRH

[794] Gilles Barthe, Benjamin Grégoire, and Santiago Zanella Béguelin. Probabilistic relational Hoare logics for computer-aided security proofs. Lecture Notes in Computer Science, 7342:1–6, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31113-0_1/.

Hoare:2012:LPU

[795] Tony Hoare and Stephan van Staden. The laws of programming unify process calculi. Lecture Notes in Computer Science, 7342:7–22, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31113-0_2/.

Ghica:2012:GS

[796] Dan R. Ghica. The geometry of synthesis. Lecture Notes in Computer Science, 7342:23–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/accesspage/
chapter/10.1007/978-3-642-31113-0_3.

Lux:2012:SID

[797] Alexander Lux, Heiko Mantel, and Matthias Perner. Schedulerindependent declassification. Lecture Notes in Computer Science, 7342: 25–47, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31113-0_4/.

Morgan:2012:EPT

[798] Carroll Morgan. Elementary probability theory in the Eindhoven style. Lecture Notes in Computer Science, 7342:48–73, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31113-0_5/.

Mandel:2012:SBS

[799] Louis Mandel and Florence Plateau. Scheduling and buffer sizing of n-synchronous systems. Lecture Notes in Computer Science, 7342:74– 101, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31113-0_6/.

Dongol:2012:DRT

[800] Brijesh Dongol and Ian J. Hayes. Deriving real-time action systems controllers from multiscale system specifications. Lecture Notes in Computer Science, 7342:102–131, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31113-0_7/.

Sergey:2012:CGA

[801] Ilya Sergey, Jan Midtgaard, and Dave Clarke. Calculating graph algorithms for dominance and shortest path. Lecture Notes in Computer Science, 7342:132–156, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31113-0_8/.

Backhouse:2012:FPP

[802] Roland Backhouse. First-past-the-post games. Lecture Notes in Computer Science, 7342:157–176, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31113-0_9/.

Dang:2012:REC

[803] Han-Hing Dang and Bernhard Möller. Reverse exchange for concurrency and local reasoning. Lecture Notes in Computer Science, 7342: 177–197, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31113-0_10/.

Guttmann:2012:UCS

[804] Walter Guttmann. Unifying correctness statements. Lecture Notes in Computer Science, 7342:198–219, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31113-0_11/.

Armstrong:2012:DTP

[805] Alasdair Armstrong, Simon Foster, and Georg Struth. Dependently typed programming based on automated theorem proving. Lecture Notes in Computer Science, 7342:220–240, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31113-0_12/.

Moller:2012:ACD

[806] Bernhard Möller, Patrick Roocks, and Markus Endres. An algebraic calculus of database preferences. Lecture Notes in Computer Science, 7342: 241–262, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31113-0_13/.

Bahr:2012:MTA

[807] Patrick Bahr. Modular tree automata. Lecture Notes in Computer Science, 7342:263–299, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-31113-0_14/.

Paterson:2012:CAF

[808] Ross Paterson. Constructing applicative functors. Lecture Notes in Computer Science, 7342:300–323, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31113-0_15/.

Hinze:2012:KEP

[809] Ralf Hinze. Kan extensions for program optimisation or: Art and Dan explain an old trick. Lecture Notes in Computer Science, 7342: 324–362, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31113-0_16/.

Anonymous:2012:BMn

[810] Anonymous. Back matter. Lecture Notes in Computer Science, 7342: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-31113-0/1.

Anonymous:2012:FMbo

[811] Anonymous. Front matter. Lecture Notes in Computer Science, 7342: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31113-0/1.

Houdek:2012:IRE

[812] Frank Houdek. Improving requirements engineering processes. Lecture Notes in Computer Science, 7343:1–2, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/accesspage/chapter/10.1007/978-3-642-31063-8_1.

Raninen:2012:DDA

[813] Anu Raninen, Tanja Toroi, Hannu Vainio, and Jarmo J. Ahonen. Defect data analysis as input for software process improvement. Lecture Notes in Computer Science, 7343:3–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31063-8_2/.

Heiskanen:2012:TPI

[814] Henri Heiskanen, Mika Maunumaa, and Mika Katara. A test process improvement model for automated test generation. Lecture Notes in Computer Science, 7343:17–31, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31063-8_3/.

Naupac:2012:SPI

[815] Verónica Ñaupac, Robert Arisaca, and Abraham Dávila. Software process improvement and certification of a small company using the NTP 291 100 (MoProSoft). Lecture Notes in Computer Science, 7343:32– 43, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31063-8_4/.

Ferreira:2012:DPO

[816] Nuno Ferreira, Nuno Santos, Ricardo J. Machado, and Dragan Gašević. Derivation of process-oriented logical architectures: An elicitation approach for cloud design. Lecture Notes in Computer Science, 7343: 44–58, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31063-8_5/.

delBianco:2012:PSM

[817] Vieri del Bianco, Luigi Lavazza, and Sandro Morasca. A proposal for simplified model-based cost estimation models. Lecture Notes in Computer Science, 7343:59–73, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31063-8_6/.

Castro:2012:ESP

[818] Oscar Castro, Angelina Espinoza, and Alfonso Martínez-Martínez. Estimating the software product value during the development process. Lecture Notes in Computer Science, 7343:74–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31063-8_7/.

Washizaki:2012:RMP

[819] Hironori Washizaki, Toshikazu Koike, Rieko Namiki, and Hiroyuki Tanabe. Reusability metrics for program source code written in C language and their evaluation. Lecture Notes in Computer Science, 7343: 89–103, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31063-8_8/.

Host:2012:MEP

[820] Martin Höst. Modeling the effects of project management strategies on long-term product knowledge. Lecture Notes in Computer Science, 7343: 104–115, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31063-8_9/.

Vlaanderen:2012:GAP

[821] Kevin Vlaanderen, Peter van Stijn, Sjaak Brinkkemper, and Inge van de Weerd. Growing into agility: Process implementation paths for scrum. Lecture Notes in Computer Science, 7343:116–130, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.com/chapter/10.1007/978-3-642-31063-8_10/.

Llanos:2012:DBT

[822] John Wilmar Castro Llanos and Silvia Teresita Acuña Castillo. Differences between traditional and open source development activities. Lecture Notes in Computer Science, 7343:131–144, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31063-8_11/.

Rodriguez:2012:ADC

[823] Pilar Rodríguez, Jouni Markkula, Markku Oivo, and Juan Garbajosa. Analyzing the drivers of the combination of lean and agile in software development companies. Lecture Notes in Computer Science, 7343: 145–159, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31063-8_12/.

Moe:2012:FSI

[824] Nils Brede Moe, Sebastian Barney, Aybüke Aurum, Mahvish Khurum, and Claes Wohlin. Fostering and sustaining innovation in a fast growing agile company. Lecture Notes in Computer Science, 7343:160–174, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31063-8_13/.

Svensson:2012:SAM

[825] Richard Berntsson Svensson, Aybüke Aurum, Barbara Paech, Tony Gorschek, and Devesh Sharma. Software architecture as a means of communication in a globally distributed software development context. Lecture Notes in Computer Science, 7343:175–189, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31063-8_14/.

Smite:2012:STC

[826] Darja Šmite and Zane Galviņa. Socio-technical congruence sabotaged by a hidden onshore outsourcing relationship: Lessons learned from an empirical study. Lecture Notes in Computer Science, 7343:190–202, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31063-8_15/.

Monasor:2012:PTG

[827] Miguel J. Monasor, Aurora Vizcaíno, and Mario Piattini. Providing training in GSD by using a virtual environment. Lecture Notes in Computer Science, 7343:203–217, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31063-8_16/.

Jantti:2012:IIS

[828] Marko Jäntti. Improving IT service desk and service management processes in Finnish tax administration: a case study on service engineering. Lecture Notes in Computer Science, 7343:218–232, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31063-8_17/.

Meyer:2012:EEK

[829] Sebastian Meyer, Anna Averbakh, Torsten Ronneberger, and Kurt Schneider. Experiences from establishing knowledge management in a joint research project. Lecture Notes in Computer Science, 7343:233–247, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31063-8_18/.

Halstead:2012:ILD

[830] Susanne Halstead, Rosario Ortiz, Mario Córdova, and Miguel Seguí. The impact of lack in domain or technology experience on the accuracy of expert effort estimates in software projects. Lecture Notes in Computer Science, 7343:248–259, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31063-8_19/.

Hanakawa:2012:MMQ

[831] Noriko Hanakawa and Masaki Obana. A metrics for meeting quality on a software requirement acquisition phase. Lecture Notes in Computer Science, 7343:260–274, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31063-8_20/.

Anonymous:2012:FMbp

[832] Anonymous. Front matter. Lecture Notes in Computer Science, 7343: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31063-8/1.

Ekberg:2012:AEP

[833] Jan-Erik Ekberg, Alexandra Afanasyeva, and N. Asokan. Authenticated encryption primitives for size-constrained trusted computing. Lecture Notes in Computer Science, 7344:1–18, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30921-2_1/.

Smart:2012:AET

[834] Matt Smart and Eike Ritter. Auditable envelopes: Tracking anonymity revocation using trusted computing. Lecture Notes in Computer Science, 7344:19–33, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30921-2_2/.

Vasudevan:2012:LTS

[835] Amit Vasudevan, Bryan Parno, Ning Qu, Virgil D. Gligor, and Adrian Perrig. Lockdown: Towards a safe and practical architecture for security applications on commodity platforms. Lecture Notes in Computer Science, 7344:34–54, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30921-2_3/.

DeCristofaro:2012:EFP

[836] Emiliano De Cristofaro and Gene Tsudik. Experimenting with fast private set intersection. Lecture Notes in Computer Science, 7344:55–73, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30921-2_4/.

Sangorrin:2012:RDS

[837] Daniel Sangorrín, Shinya Honda, and Hiroaki Takada. Reliable device sharing mechanisms for dual-OS embedded trusted computing. Lecture Notes in Computer Science, 7344:74–91, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30921-2_5/.

Hasan:2012:MUC

[838] Zahid Hasan, Alina Krischkowsky, and Manfred Tscheligi. Modelling user-centered-trust (UCT) in software systems: Interplay of trust, affect and acceptance model. Lecture Notes in Computer Science, 7344: 92–109, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30921-2_6/.

Murphy:2012:CPU

[839] Julian Murphy. Clockless physical unclonable functions. Lecture Notes in Computer Science, 7344:110–121, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30921-2_7/.

Pirker:2012:LDH

[840] Martin Pirker, Johannes Winter, and Ronald Toegl. Lightweight distributed heterogeneous attested Android clouds. Lecture Notes in Computer Science, 7344:122–141, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30921-2_8/.

Kocabas:2012:CPB

[841] Ünal Kocabaş, Andreas Peter, Stefan Katzenbeisser, and Ahmad-Reza Sadeghi. Converse PUF-based authentication. Lecture Notes in Computer Science, 7344:142–158, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30921-2_9/.

Vasudevan:2012:TEM

[842] Amit Vasudevan, Emmanuel Owusu, Zongwei Zhou, James Newsome, and Jonathan M. McCune. Trustworthy execution on mobile devices: What security properties can my mobile platform give me? Lecture Notes in Computer Science, 7344:159–178, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30921-2_10/.

Schiffman:2012:VSI

[843] Joshua Schiffman, Hayawardh Vijayakumar, and Trent Jaeger. Verifying system integrity by proxy. Lecture Notes in Computer Science, 7344: 179–200, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30921-2_11/.

Cheng:2012:VBP

[844] Yueqiang Cheng and Xuhua Ding. Virtualization based password protection against malware in untrusted operating systems. Lecture Notes in Computer Science, 7344:201–218, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30921-2_12/.

Dmitrienko:2012:SDA

[845] Alexandra Dmitrienko, Ahmad-Reza Sadeghi, Sandeep Tamrakar, and Christian Wachsmann. SmartTokens: Delegable access control with NFC-enabled Smartphones. Lecture Notes in Computer Science, 7344: 219–238, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30921-2_13/.

Kumar:2012:BLA

[846] Apurva Kumar. A belief logic for analyzing security of Web protocols. Lecture Notes in Computer Science, 7344:239–254, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30921-2_14/.

Namiluko:2012:PBM

[847] Cornelius Namiluko and Andrew Martin. Provenance-based model for verifying trust-properties. Lecture Notes in Computer Science, 7344: 255–272, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30921-2_15/.

Cai:2012:PMB

[848] Liang Cai and Hao Chen. On the practicality of motion based keystroke inference attack. Lecture Notes in Computer Science, 7344:273–290, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30921-2_16/.

Gibler:2012:AAD

[849] Clint Gibler, Jonathan Crussell, Jeremy Erickson, and Hao Chen. AndroidLeaks: Automatically detecting potential privacy leaks in Android applications on a large scale. Lecture Notes in Computer Science, 7344: 291–307, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30921-2_17/.

Kirlappos:2012:WTS

[850] Iacovos Kirlappos, M. Angela Sasse, and Nigel Harvey. Why trust seals don’t work: a study of user perceptions and behavior. Lecture Notes in Computer Science, 7344:308–324, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30921-2_18/.

Zheng:2012:LNP

[851] Saijing Zheng, Pan Shi, Heng Xu, and Cheng Zhang. Launching the new profile on facebook: Understanding the triggers and outcomes of users’ privacy concerns. Lecture Notes in Computer Science, 7344: 325–339, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30921-2_19/.

Anonymous:2012:BMo

[852] Anonymous. Back matter. Lecture Notes in Computer Science, 7344: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-30921-2/1.

Anonymous:2012:FMbq

[853] Anonymous. Front matter. Lecture Notes in Computer Science, 7344: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-30921-2/1.

He:2012:ADL

[854] Qing He, Ye Duan, and Danyang Zhang. Automatic detailed localization of facial features. Lecture Notes in Computer Science, 7345:1– 9, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31087-4_1/.

Ozdikis:2012:CBI

[855] Ozer Ozdikis, Pinar Senkul, and Siyamed Sinir. Confidence-based incremental classification for objects with limited attributes in vertical search. Lecture Notes in Computer Science, 7345:10–19, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31087-4_2/.

Huang:2012:ADO

[856] Huaming Huang, Kishan Mehrotra, and Chilukuri K. Mohan. Algorithms for detecting outliers via clustering and ranks. Lecture Notes in Computer Science, 7345:20–29, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31087-4_3/.

Yoshimura:2012:GCO

[857] Takuya Yoshimura, Yutaro Fujii, and Takayuki Ito. Grouping cooccurrence filtering based on Bayesian filtering. Lecture Notes in Computer Science, 7345:30–39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31087-4_4/.

Yang:2012:MFB

[858] Chunsheng Yang, Yubin Yang, and Jie Liu. Model fusion-based batch learning with application to oil spills detection. Lecture Notes in Computer Science, 7345:40–47, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31087-4_5/.

Kryszkiewicz:2012:EDB

[859] Marzena Kryszkiewicz. Efficient determination of binary non-negative vector neighbors with regard to cosine similarity. Lecture Notes in Computer Science, 7345:48–57, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31087-4_6/.

Itoyama:2012:ACR

[860] Katsutoshi Itoyama, Tetsuya Ogata, and Hiroshi G. Okuno. Automatic Chord recognition based on probabilistic integration of acoustic features, Bass sounds, and Chord transition. Lecture Notes in Computer Science, 7345:58–67, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31087-4_7/.

Li:2012:RTT

[861] Haiguang Li, Zhao Li, Robert T. White, and Xindong Wu. A real-time transportation prediction system. Lecture Notes in Computer Science, 7345:68–77, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31087-4_8/.

Lo:2012:SRM

[862] Henry Z. Lo, Dan A. Simovici, and Wei Ding. Several remarks on mining frequent trajectories in graphs. Lecture Notes in Computer Science, 7345:78–87, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31087-4_9/.

Ni:2012:CPI

[863] Daiheng Ni, Hong Liu, Wei Ding, Yuanchang Xie, Honggang Wang, and Hossein Pishro-Nik. Cyber-physical integration to connect vehicles for transformed transportation safety and efficiency. Lecture Notes in Computer Science, 7345:88–94, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31087-4_10/.

Yu:2012:MMR

[864] Tzu-Yang Yu, Christopher Niezrecki, and Farhad Ansari. Multi-modal remote sensing system for transportation infrastructure inspection and monitoring. Lecture Notes in Computer Science, 7345:95–103, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31087-4_11/.

Tidemann:2012:OSF

[865] Axel Tidemann, Finn Olav Bjørnson, and Agnar Aamodt. Operational support in fish farming through case-based reasoning. Lecture Notes in Computer Science, 7345:104–113, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31087-4_12/.

Mohammad:2012:CMT

[866] Yasser Mohammad, Yoshimasa Ohmoto, and Toyoaki Nishida. CPMD: a Matlab toolbox for change point and constrained motif discovery. Lecture Notes in Computer Science, 7345:114–123, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31087-4_13/.

Nair:2012:TIG

[867] Binu M. Nair and Vijayan K. Asari. Time invariant gesture recognition by modelling body posture space. Lecture Notes in Computer Science, 7345: 124–133, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31087-4_14/.

Tan:2012:GBS

[868] Xing Tan. Go beyond the SCOPE: a temporal situation calculus-based software tool for time Petri nets. Lecture Notes in Computer Science, 7345:134–143, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-31087-4_15/.

Lai:2012:ACA

[869] Xiaochen Lai, Quanli Liu, Wei Wang, Likun Li, Simin Lu, and Di Wu. An algorithm of channel assignment of MAC layer in ad hoc network based on dynamic game with perfect and complete information. Lecture Notes in Computer Science, 7345:144–155, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31087-4_16/.

Nguyen:2012:SAQ

[870] Dai Quoc Nguyen, Dat Quoc Nguyen, and Son Bao Pham. A semantic approach for question analysis. Lecture Notes in Computer Science, 7345: 156–165, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31087-4_17/.

Grossmann:2012:SPE

[871] Peter Großmann, Steffen Hölldobler, Norbert Manthey, Karl Nachtigall, and Jens Opitz. Solving periodic event scheduling problems with SAT. Lecture Notes in Computer Science, 7345:166–175, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31087-4_18/.

Tamir:2012:IPC

[872] Dan E. Tamir, Charles R. King, and Mark McKenney. Improving the performance of constructive multi-start search using record-keeping. Lecture Notes in Computer Science, 7345:176–186, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31087-4_19/.

Ma:2012:MPS

[873] Li Ma and Babak Forouraghi. A modified particle swarm optimizer for engineering design. Lecture Notes in Computer Science, 7345:187– 196, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31087-4_20/.

Anonymous:2012:FMbr

[874] Anonymous. Front matter. Lecture Notes in Computer Science, 7345: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31087-4/1.

Stelldinger:2012:CDR

[875] Peer Stelldinger. Connect the dots: The reconstruction of region boundaries from contour sampling points. Lecture Notes in Computer Science, 7346:1–13, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32313-3_1/.

Lachaud:2012:DSA

[876] Jacques-Olivier Lachaud. Digital shape analysis with maximal segments. Lecture Notes in Computer Science, 7346:14–27, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32313-3_2/.

Mesmoudi:2012:DCE

[877] Mohammed Mostefa Mesmoudi and Leila De Floriani. Discrete curvature estimation methods for triangulated surfaces. Lecture Notes in Computer Science, 7346:28–42, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32313-3_3/.

Soille:2012:MHR

[878] Pierre Soille and Laurent Najman. On morphological hierarchical representations for image processing and spatial data clustering. Lecture Notes in Computer Science, 7346:43–67, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32313-3_4/.

Kiwanuka:2012:RMI

[879] Fred N. Kiwanuka and Michael H. F. Wilkinson. Radial moment invariants for attribute filtering in 3D. Lecture Notes in Computer Science, 7346:68–81, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32313-3_5/.

Coeurjolly:2012:VAD

[880] David Coeurjolly. Volumetric analysis of digital objects using distance transformation: Performance issues and extensions. Lecture Notes in Computer Science, 7346:82–92, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32313-3_6/.

Kothe:2012:GAE

[881] Ullrich Köthe, Björn Andres, and Thorben Kröger. Geometric analysis of 3D electron microscopy data. Lecture Notes in Computer Science, 7346: 93–108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32313-3_7/.

Gara:2012:MLP

[882] Mihály Gara, Tamás Sámuel Tasi, and Péter Balázs. Machine learning as a preprocessing phase in discrete tomography. Lecture Notes in Computer Science, 7346:109–124, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32313-3_8/.

Heyvaert:2012:FPE

[883] Michaël Heyvaert and Peter Veelaert. Fast planarity estimation and region growing on GPU. Lecture Notes in Computer Science, 7346: 125–139, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32313-3_9/.

Levillain:2012:WRD

[884] Roland Levillain and Thierry Géraud. Writing reusable digital topology algorithms in a generic image processing framework. Lecture Notes in Computer Science, 7346:140–153, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32313-3_10/.

Gurevich:2012:NIM

[885] Igor Gurevich, Artem Myagkov, and Vera Yashina. A new image-mining technique for automation of Parkinson’s disease research. Lecture Notes in Computer Science, 7346:154–167, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32313-3_11/.

Anonymous:2012:BMp

[886] Anonymous. Back matter. Lecture Notes in Computer Science, 7346: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-32313-3/1.

Anonymous:2012:FMbs

[887] Anonymous. Front matter. Lecture Notes in Computer Science, 7346: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32313-3/1.

Hoare:2012:NMC

[888] Tony Hoare. Net models for concurrent object behaviour. Lecture Notes in Computer Science, 7347:1–2, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

accesspage/chapter/10.1007/978-3-642-31131-4_1.

Finkel:2012:TWC

[889] Alain Finkel and Jean Goubault-Larrecq. The theory of WSTS: The case of complete WSTS. Lecture Notes in Computer Science, 7347: 3–31, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31131-4_2/.

Mailund:2012:UCP

[890] Thomas Mailund, Anders E. Halager, and Michael Westergaard. Using colored Petri nets to construct coalescent hidden Markov models: Automatic translation from demographic specifications to efficient inference methods. Lecture Notes in Computer Science, 7347:32–50, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31131-4_3/.

Sole:2012:SBD

[891] Marc Solé and Josep Carmona. An SMT-based discovery algorithm for C-nets. Lecture Notes in Computer Science, 7347:51–71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31131-4_4/.

vanderAalst:2012:DPM

[892] Wil M. P. van der Aalst. Decomposing process mining problems using passages. Lecture Notes in Computer Science, 7347:72–91, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31131-4_5/.

Liu:2012:CSP

[893] Guan Jun Liu, Jun Sun, Yang Liu, and Jin Song Dong. Complexity of the soundness problem of bounded workflow nets. Lecture Notes in Computer Science, 7347:92–107, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31131-4_6/.

Martos-Salgado:2012:CSP

[894] María Martos-Salgado and Fernando Rosa-Velardo. Cost soundness for priced resource-constrained workflow nets. Lecture Notes in Computer Science, 7347:108–127, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31131-4_7/.

Badouel:2012:RWN

[895] Eric Badouel. On the α-reconstructibility of workflow nets. Lecture Notes in Computer Science, 7347:128–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31131-4_8/.

Weidlich:2012:PFR

[896] Matthias Weidlich and Jan Martijn van der Werf. On profiles and footprints — relational semantics for Petri nets. Lecture Notes in Computer Science, 7347:148–167, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31131-4_9/.

Fahland:2012:DAS

[897] Dirk Fahland and Robert Prüfer. Data and abstraction for scenario-based modeling with Petri nets. Lecture Notes in Computer Science, 7347: 168–187, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31131-4_10/.

Wang:2012:MCP

[898] Xu Wang. Maximal confluent processes. Lecture Notes in Computer Science, 7347:188–207, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31131-4_11/.

Valmari:2012:ONA

[899] Antti Valmari and Henri Hansen. Old and new algorithms for minimal coverability sets. Lecture Notes in Computer Science, 7347:208– 227, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31131-4_12/.

Lehmann:2012:SSS

[900] Andreas Lehmann, Niels Lohmann, and Karsten Wolf. Stubborn sets for simple linear time properties. Lecture Notes in Computer Science, 7347: 228–247, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31131-4_13/.

Evangelista:2012:HFL

[901] Sami Evangelista and Lars Michael Kristensen. Hybrid on-the-fly LTL model checking with the sweep-line method. Lecture Notes in Computer Science, 7347:248–267, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31131-4_14/.

Rakow:2012:SSP

[902] Astrid Rakow. Safety slicing Petri nets. Lecture Notes in Computer Science, 7347:268–287, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31131-4_15/.

Gil-Costa:2012:CPV

[903] Veronica Gil-Costa, Jair Lobos, Alonso Inostrosa-Psijas, and Mauricio Marin. Capacity planning for vertical search engines: An approach based on coloured Petri nets. Lecture Notes in Computer Science, 7347: 288–307, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31131-4_16/.

Nakatumba:2012:ICE

[904] Joyce Nakatumba, Michael Westergaard, and Wil M. P. van der Aalst. An infrastructure for cost-effective testing of operational support algorithms based on colored Petri nets. Lecture Notes in Computer Science, 7347: 308–327, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31131-4_17/.

Bera:2012:DWT

[905] Debjyoti Bera, Kees M. van Hee, and Jan Martijn van der Werf. Designing weakly terminating ROS systems. Lecture Notes in Computer Science, 7347:328–347, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-31131-4_18/.

Cristini:2012:NWN

[906] Frédéric Cristini and Catherine Tessier. Nets-within-nets to model innovative space system architectures. Lecture Notes in Computer Science, 7347:348–367, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-31131-4_19/.

Heitmann:2012:SNW

[907] Frank Heitmann and Michael Köhler-Bußmeier. P - and T -systems in the nets-within-nets-formalism. Lecture Notes in Computer Science, 7347: 368–387, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31131-4_20/.

Anonymous:2012:FMbt

[908] Anonymous. Front matter. Lecture Notes in Computer Science, 7347: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31131-4/1.

Ostell:2012:ENS

[909] James M. Ostell. Entrez: The NCBI search and discovery engine. Lecture Notes in Computer Science, 7348:1–4, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31040-9_1/.

Trissl:2012:IRI

[910] Silke Trißl, Philipp Hussels, and Ulf Leser. InterOnto — ranking inter-ontology links. Lecture Notes in Computer Science, 7348:5–20, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31040-9_2/.

Benik:2012:FCG

[911] Joseph Benik, Caren Chang, Louiqa Raschid, Maria-Esther Vidal, and Guillermo Palma. Finding cross genome patterns in annotation graphs. Lecture Notes in Computer Science, 7348:21–36, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31040-9_3/.

Maskat:2012:PYG

[912] Ruhaila Maskat, Norman W. Paton, and Suzanne M. Embury. Pay-as-you-go ranking of schema mappings using query logs. Lecture Notes in Computer Science, 7348:37–52, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31040-9_4/.

Barker:2012:CSU

[913] Ken Barker. Combining structured and unstructured knowledge sources for question answering in Watson. Lecture Notes in Computer Science, 7348:53–55, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31040-9_5/.

Shironoshita:2012:CDI

[914] E. Patrick Shironoshita, Yves R. Jean-Mary, Ray M. Bradley, and Patricia Buendia. Cancer data integration and querying with GeneTegra. Lecture Notes in Computer Science, 7348:56–70, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31040-9_6/.

Ogbuji:2012:ILD

[915] Chimezie Ogbuji and Rong Xu. Integrating large, disparate biomedical ontologies to boost organ development network connectivity. Lecture Notes in Computer Science, 7348:71–82, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31040-9_7/.

Burger:2012:VCG

[916] John D. Burger, Emily Doughty, Sam Bayer, David Tresner-Kirsch, and Ben Wellner. Validating candidate gene-mutation relations in MEDLINE abstracts via crowdsourcing. Lecture Notes in Computer Science, 7348: 83–91, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31040-9_8/.

Cimino:2012:BTR

[917] James J. Cimino. The biomedical translational research information system: Clinical data integration at the national institutes of health. Lecture Notes in Computer Science, 7348:92, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/accesspage/chapter/10.1007/978-3-642-31040-9_9.

Segagni:2012:OII

[918] Daniele Segagni, Matteo Gabetta, Valentina Tibollo, Alberto Zambelli, and Silvia G. Priori. ONCO-i2b2: Improve patients selection through case-based information retrieval techniques. Lecture Notes in Computer Science, 7348:93–99, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31040-9_10/.

Demner-Fushman:2012:SSF

[919] Dina Demner-Fushman and Swapna Abhyankar. Syntactic-semantic frames for clinical cohort identification queries. Lecture Notes in Computer Science, 7348:100–112, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31040-9_11/.

Anonymous:2012:BMq

[920] Anonymous. Back matter. Lecture Notes in Computer Science, 7348: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-31040-9/1.

Anonymous:2012:FMbu

[921] Anonymous. Front matter. Lecture Notes in Computer Science, 7348: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31040-9/1.

Seidewitz:2012:EUM

[922] Ed Seidewitz. Executable UML: From multi-domain to multi-core. Lecture Notes in Computer Science, 7349:1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

accesspage/chapter/10.1007/978-3-642-31491-9_1.

Lonn:2012:MMA

[923] Henrik Lönn. Models meeting automotive design challenges. Lecture Notes in Computer Science, 7349:2–3, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

accesspage/chapter/10.1007/978-3-642-31491-9_2.

Mosser:2012:CMC

[924] Sébastien Mosser and Mireille Blay-Fornarino. A commutative model composition operator to support software adaptation. Lecture Notes in Computer Science, 7349:4–19, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31491-9_3/.

Votintseva:2012:CSM

[925] Anjelika Votintseva and Petra Witschel. Comparative study of modelbased and multi-domain system engineering approaches for industrial settings. Lecture Notes in Computer Science, 7349:20–31, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31491-9_4/.

Kuhlmann:2012:SSB

[926] Mirco Kuhlmann and Martin Gogolla. Strengthening SAT-based validation of UML/OCL models by representing collections as relations. Lecture Notes in Computer Science, 7349:32–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31491-9_5/.

Elaasar:2012:MIT

[927] Maged Elaasar and Yvan Labiche. Model interchange testing: a process and a case study. Lecture Notes in Computer Science, 7349:49– 61, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31491-9_6/.

Goldschmidt:2012:IDS

[928] Thomas Goldschmidt and Wolfgang Mahnke. An internal domain-specific language for constructing OPC UA queries and event filters. Lecture Notes in Computer Science, 7349:62–73, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31491-9_7/.

Briand:2012:CUS

[929] Lionel Briand, Yvan Labiche, and Yanhua Liu. Combining UML sequence and state machine diagrams for data-flow based integration testing. Lecture Notes in Computer Science, 7349:74–89, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.com/chapter/10.1007/978-3-642-31491-9_8/.

Selim:2012:MTM

[930] Gehan M. K. Selim, Shige Wang, and James R. Cordy. Model transformations for migrating legacy models: An industrial case study. Lecture Notes in Computer Science, 7349:90–101, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31491-9_9/.

Rath:2012:DFE

[931] István Ráth, Ábel Hegedüs, and Dániel Varró. Derived features for EMF by integrating advanced model queries. Lecture Notes in Computer Science, 7349:102–117, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31491-9_10/.

Kolovos:2012:LAM

[932] Dimitrios S. Kolovos and Louis M. Rose. A lightweight approach for managing XML documents with MDE languages. Lecture Notes in Computer Science, 7349:118–132, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31491-9_11/.

Yue:2012:BGB

[933] Tao Yue and Shaukat Ali. Bridging the gap between requirements and aspect state machines to support non-functional testing: Industrial case studies. Lecture Notes in Computer Science, 7349:133–145, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31491-9_12/.

Puissant:2012:BRP

[934] Jorge Pinna Puissant and Ragnhild Van Der Straeten. Badger: a regression planner to resolve design model inconsistencies. Lecture Notes in Computer Science, 7349:146–161, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31491-9_13/.

Zhang:2012:AOM

[935] Gefei Zhang. Aspect-oriented modeling of mutual exclusion in UML state machines. Lecture Notes in Computer Science, 7349:162–177, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31491-9_14/.

Pfeiffer:2012:TML

[936] Rolf-Helge Pfeiffer and Andrzej Wasowski. TexMo: a multi-language development environment. Lecture Notes in Computer Science, 7349: 178–193, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31491-9_15/.

Atkinson:2012:FEM

[937] Colin Atkinson, Ralph Gerbig, and Bastian Kennel. On-the-fly emendation of multi-level models. Lecture Notes in Computer Science, 7349: 194–209, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31491-9_16/.

Rieke:2012:SRR

[938] Jan Rieke and Oliver Sudmann. Specifying refinement relations in vertical model transformations. Lecture Notes in Computer Science, 7349: 210–225, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31491-9_17/.

Behjati:2012:MBA

[939] Razieh Behjati, Shiva Nejati, and Tao Yue. Model-based automated and guided configuration of embedded software systems. Lecture Notes in Computer Science, 7349:226–243, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31491-9_18/.

Buttner:2012:LSR

[940] Fabian Büttner and Jordi Cabot. Lightweight string reasoning for OCL. Lecture Notes in Computer Science, 7349:244–258, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31491-9_19/.

deLara:2012:DST

[941] Juan de Lara and Esther Guerra. Domain-specific textual meta-modelling languages for model driven engineering. Lecture Notes in Computer Science, 7349:259–274, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31491-9_20/.

Anonymous:2012:FMbv

[942] Anonymous. Front matter. Lecture Notes in Computer Science, 7349: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31491-9/1.

Bordihn:2012:LAE

[943] Henning Bordihn, Martin Kutrib, and Bianca Truthe, editors. Languages Alive: Essays Dedicated to Jürgen Dassow on the Occasion of His 65th Birthday, volume 7300 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31643-3 (print), 3-642-31644-1 (e-book). ISSN 0302-9743 (print), 1611-

3349 (electronic). LCCN ???? URL http://www.springerlink.com/
content/978-3-642-31644-9.

Tan:2012:AKDa

[944] Pang-Ning Tan, Sanjay Chawla, Chin Kuan Ho, and James Bailey, editors. Advances in Knowledge Discovery and Data Mining: 16th Pacific– Asia Conference, PAKDD 2012, Kuala Lumpur, Malaysia, May 29–June 1, 2012, Proceedings, Part I, volume 7301 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-30216-5 (print), 3-642-30217-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.

springerlink.com/content/978-3-642-30217-6.

Tan:2012:AKDb

[945] Pang-Ning Tan, Sanjay Chawla, Chin Kuan Ho, and James Bailey, editors. Advances in Knowledge Discovery and Data Mining: 16th Pacific– Asia Conference, PAKDD 2012, Kuala Lumpur, Malaysia, May 29–June 1, 2012, Proceedings, Part II, volume 7302 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30219-X (print), 3-642-30220-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.

springerlink.com/content/978-3-642-30220-6.

Pankratius:2012:MSE

[946] Victor Pankratius and Michael Philippsen, editors. Multicore Software Engineering, Performance, and Tools: International Conference, MSEPT 2012, Prague, Czech Republic, May 31–June 1, 2012. Proceedings, volume 7303 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31201-2 (print), 3-642-31202-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-642-
31202-1.

Furia:2012:OMC

[947] Carlo A. Furia and Sebastian Nanz, editors. Objects, Models, Components, Patterns: 50th International Conference, TOOLS 2012, Prague, Czech Republic, May 29–31, 2012. Proceedings, volume 7304 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30560-1 (print), 3-642-30561-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-30561-0.

Brucker:2012:TPI

[948] Achim D. Brucker and Jacques Julliand, editors. Tests and Proofs: 6th International Conference, TAP 2012, Prague, Czech Republic, May 31–June 1, 2012. Proceedings, volume 7305 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30472-9 (print), 3-642-30473-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-30473-6.

Gschwind:2012:SCI

[949] Thomas Gschwind, Flavio De Paoli, Volker Gruhn, and Matthias Book, editors. Software Composition: 11th International Conference, SC 2012, Prague, Czech Republic, May 31–June 1, 2012. Proceedings, volume 7306 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30563-6 (print), 3-642-30564-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-642-
30564-1.

Hu:2012:TPM

[950] Zhenjiang Hu and Juan de Lara, editors. Theory and Practice of Model Transformations: 5th International Conference, ICMT 2012, Prague, Czech Republic, May 28–29, 2012. Proceedings, volume 7307 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30475-3 (print), 3-642-30476-1 (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL

http://www.springerlink.com/content/978-3-642-30476-7.

Brorsson:2012:RST

[951] Mats Brorsson and Luís Miguel Pinho, editors. Reliable Software Technologies — Ada–Europe 2012: 17th Ada–Europe International Conference on Reliable Software Technologies, Stockholm, Sweden, June 11–15, 2012. Proceedings, volume 7308 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30597-0 (print), 3-642-30598-9 (e-book). ISSN 0302-9743 (print), 1611-

3349 (electronic). LCCN ???? URL http://www.springerlink.com/
content/978-3-642-30598-6.

Ferri:2012:CTI

[952] Massimo Ferri, Patrizio Frosini, Claudia Landi, Andrea Cerri, and Barbara Di Fabio, editors. Computational Topology in Image Context: 4th International Workshop, CTIC 2012, Bertinoro, Italy, May 28–30, 2012.

Proceedings, volume 7309 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30237-8 (print), 3-642-30238-6 (e-book). ISSN 0302-9743 (print), 1611-

3349 (electronic). LCCN ???? URL http://www.springerlink.com/
content/978-3-642-30238-1.

Kosseim:2012:AAI

[953] Leila Kosseim and Diana Inkpen, editors. Advances in Artificial Intelligence: 25th Canadian Conference on Artificial Intelligence, Canadian AI 2012, Toronto, ON, Canada, May 28–30, 2012. Proceedings, volume 7310 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30352-8 (print), 3-642-30353-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-642-
30353-1.

Liu:2012:ACI

[954] Jing Liu, Cesare Alippi, Bernadette Bouchon-Meunier, Garrison W. Greenwood, and Hussein A. Abbass, editors. Advances in Computational Intelligence: IEEE World Congress on Computational Intelligence, WCCI 2012, Brisbane, Australia, June 10–15, 2012. Plenary/Invited Lectures, volume 7311 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30686-1 (print), 3-642-30687-X (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-30687-7.

Chapman:2012:OHW

[955] Barbara M. Chapman, Federico Massaioli, Matthias S. Müller, and Marco Rorro, editors. OpenMP in a Heterogeneous World: 8th International Workshop on OpenMP, IWOMP 2012, Rome, Italy, June 11–13, 2012. Proceedings, volume 7312 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30960-7 (print), 3-642-30961-5 (e-book). ISSN 0302-9743 (print), 1611-

3349 (electronic). LCCN ???? URL http://www.springerlink.com/
content/978-3-642-30961-8.

Noble:2012:EOO

[956] James Noble, editor. ECOOP 2012 — Object-Oriented Programming: 26th European Conference, Beijing, China, June 11–16, 2012. Proceedings, volume 7313 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31056-7 (print), 3-642-31057-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-31057-7.

Al-Begain:2012:ASM

[957] Khalid Al-Begain, Dieter Fiems, and Jean-Marc Vincent, editors. Analytical and Stochastic Modeling Techniques and Applications: 19th International Conference, ASMTA 2012, Grenoble, France, June 4–6, 2012. Proceedings, volume 7314 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30781-7 (print), 3-642-30782-5 (e-book). ISSN 0302-9743 (print), 1611-

3349 (electronic). LCCN ???? URL http://www.springerlink.com/
content/978-3-642-30782-9.

Cerri:2012:ITS

[958] Stefano A. Cerri, William J. Clancey, Giorgos Papadourakis, and Kitty Panourgia, editors. Intelligent Tutoring Systems: 11th International Conference, ITS 2012, Chania, Crete, Greece, June 14–18, 2012. Proceedings, volume 7315 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30949-6 (print), 3-642-30950-X (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-30950-2.

Derrick:2012:ASM

[959] John Derrick, John Fitzgerald, Stefania Gnesi, Sarfraz Khurshid, Michael Leuschel, Steve Reeves, and Elvinia Riccobene, editors. Abstract State Machines, Alloy, B, VDM, and Z: Third International Conference, ABZ 2012, Pisa, Italy, June 18–21, 2012. Proceedings, volume 7316 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30884-8 (print), 3-642-30885-6 (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL

http://www.springerlink.com/content/978-3-642-30885-7.

Cimatti:2012:TAS

[960] Alessandro Cimatti and Roberto Sebastiani, editors. Theory and Applications of Satisfiability Testing — SAT 2012: 15th International Conference, Trento, Italy, June 17–20, 2012. Proceedings, volume 7317 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31611-5 (print), 3-642-31612-3 (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL

http://www.springerlink.com/content/978-3-642-31612-8.

Kay:2012:PCI

[961] Judy Kay, Paul Lukowicz, Hideyuki Tokuda, Patrick Olivier, and Antonio Krüger, editors. Pervasive Computing: 10th International Conference, Pervasive 2012, Newcastle, UK, June 18–22, 2012. Proceedings, volume 7319 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31204-7 (print), 3-642-31205-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-642-
31205-2.

Bernardo:2012:FMM

[962] Marco Bernardo, Vittorio Cortellessa, and Alfonso Pierantonio, editors. Formal Methods for Model-Driven Engineering: 12th International School on Formal Methods for the Design of Computer, Communication, and Software Systems, SFM 2012, Bertinoro, Italy, June 18–23, 2012. Advanced Lectures, volume 7320 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30981-X (print), 3-642-30982-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-30982-3.

Derrick:2012:IFM

[963] John Derrick, Stefania Gnesi, Diego Latella, and Helen Treharne, editors. Integrated Formal Methods: 9th International Conference, IFM 2012, Pisa, Italy, June 18–21, 2012. Proceedings, volume 7321 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30728-0 (print), 3-642-30729-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-30729-4.

Askoxylakis:2012:IST

[964] Ioannis Askoxylakis, Henrich C. Pöhls, and Joachim Posegga, editors. Information Security Theory and Practice. Security, Privacy and Trust in Computing Systems and Ambient Intelligent Ecosystems: 6th IFIP WG 11.2 International Workshop, WISTP 2012, Egham, UK, June 20–22, 2012. Proceedings, volume 7322 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30954-2 (print), 3-642-30955-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-30955-7.

Bonato:2012:AMW

[965] Anthony Bonato and Jeannette Janssen, editors. Algorithms and Models for the Web Graph: 9th International Workshop, WAW 2012, Halifax, NS, Canada, June 22–23, 2012. Proceedings, volume 7323 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30540-7 (print), 3-642-30541-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-30541-2.

Campilho:2012:IARa

[966] Aurélio Campilho and Mohamed Kamel, editors. Image Analysis and Recognition: 9th International Conference, ICIAR 2012, Aveiro, Portugal, June 25–27, 2012. Proceedings, Part I, volume 7324 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31294-2 (print), 3-642-31295-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-31295-3.

Campilho:2012:IARb

[967] Aurélio Campilho and Mohamed Kamel, editors. Image Analysis and Recognition: 9th International Conference, ICIAR 2012, Aveiro, Portugal, June 25–27, 2012. Proceedings, Part II, volume 7325 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31297-7 (print), 3-642-31298-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-31298-4.

Jezic:2012:AMA

[968] Gordan Jezic, Mario Kusek, Ngoc-Thanh Nguyen, Robert J. Howlett, and Lakhmi C. Jain, editors. Agent and Multi-Agent Systems. Technologies and Applications: 6th KES International Conference, KES-AMSTA 2012,Dubrovnik, Croatia, June 25–27, 2012. Proceedings, volume 7327 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30946-1 (print), 3-642-30947-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-30947-2.

Ralyte:2012:AIS

[969] Jolita Ralyté, Xavier Franch, Sjaak Brinkkemper, and Stanislaw Wrycza, editors. Advanced Information Systems Engineering: 24th International Conference, CAiSE 2012, Gdansk, Poland, June 25–29, 2012. Proceedings, volume 7328 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31094-X (print), 3-642-31095-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-31095-9.

Carrasco-Ochoa:2012:PRM

[970] Jesús Ariel Carrasco-Ochoa, José Francisco Martínez-Trinidad, José Arturo Olvera López, and Kim L. Boyer, editors. Pattern Recognition: 4th Mexican Conference, MCPR 2012, Huatulco, Mexico, June 27–30, 2012. Proceedings, volume 7329 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31148-2 (print), 3-642-31149-0 (e-book). ISSN 0302-9743 (print), 1611-

3349 (electronic). LCCN ???? URL http://www.springerlink.com/
content/978-3-642-31149-9.

Abolmaesumi:2012:IPC

[971] Purang Abolmaesumi, Leo Joskowicz, Nassir Navab, and Pierre Jannin, editors. Information Processing in Computer-Assisted Interventions: Third International Conference, IPCAI 2012, Pisa, Italy, June 27, 2012. Proceedings, volume 7330 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30617-9 (print), 3-642-30618-7 (e-book). ISSN 0302-9743 (print), 1611-

3349 (electronic). LCCN ???? URL http://www.springerlink.com/
content/978-3-642-30618-1.

Tan:2012:ASIa

[972] Ying Tan, Yuhui Shi, and Zhen Ji, editors. Advances in Swarm Intelligence: Third International Conference, ICSI 2012, Shenzhen, China, June 17–20, 2012 Proceedings, Part I, volume 7331 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-30975-5 (print), 3-642-30976-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.

springerlink.com/content/978-3-642-30976-2.

Tan:2012:ASIb

[973] Ying Tan, Yuhui Shi, and Zhen Ji, editors. Advances in Swarm Intelligence: Third International Conference, ICSI 2012, Shenzhen, China, June 17–20, 2012 Proceedings, Part II, volume 7332 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-31019-2 (print), 3-642-31020-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.

springerlink.com/content/978-3-642-31020-1.

Murgante:2012:CSAa

[974] Beniamino Murgante, Osvaldo Gervasi, Sanjay Misra, Nadia Nedjah, Ana Maria A. C. Rocha, David Taniar, and Bernady O. Apduhan, editors. Computational Science and Its Applications — ICCSA 2012: 12th International Conference, Salvador de Bahia, Brazil, June 18–21, 2012, Proceedings, Part I, volume 7333 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31124-5 (print), 3-642-31125-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-31125-3.

Murgante:2012:CSAb

[975] Beniamino Murgante, Osvaldo Gervasi, Sanjay Misra, Nadia Nedjah, Ana Maria A. C. Rocha, David Taniar, and Bernady O. Apduhan, editors. Computational Science and Its Applications — ICCSA 2012: 12th International Conference, Salvador de Bahia, Brazil, June 18–21, 2012, Proceedings, Part II, volume 7334 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31074-5 (print), 3-642-31075-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-31075-1.

Murgante:2012:CSAc

[976] Beniamino Murgante, Osvaldo Gervasi, Sanjay Misra, Nadia Nedjah, Ana Maria A. C. Rocha, David Taniar, and Bernady O. Apduhan, editors. Computational Science and Its Applications — ICCSA 2012: 12th International Conference, Salvador de Bahia, Brazil, June 18–21, 2012, Proceedings, Part III, volume 7335 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31136-9 (print), 3-642-31137-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-31137-6.

Murgante:2012:CSAd

[977] Beniamino Murgante, Osvaldo Gervasi, Sanjay Misra, Nadia Nedjah, Ana Maria A. C. Rocha, David Taniar, and Bernady O. Apduhan, editors. Computational Science and Its Applications — ICCSA 2012: 12th International Conference, Salvador de Bahia, Brazil, June 18–21, 2012, Proceedings, Part IV, volume 7336 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31127-X (print), 3-642-31128-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-31128-4.

Bouma:2012:NLP

[978] Gosse Bouma, Ashwin Ittoo, Elisabeth Métais, and Hans Wortmann, editors. Natural Language Processing and Information Systems: 17th International Conference on Applications of Natural Language to Information Systems, NLDB 2012, Groningen, The Netherlands, June 26–28, 2012. Proceedings, volume 7337 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31177-6 (print), 3-642-31178-4 (e-book). ISSN 0302-9743 (print), 1611-

3349 (electronic). LCCN ???? URL http://www.springerlink.com/
content/978-3-642-31178-9.

Ailamaki:2012:SSD

[979] Anastasia Ailamaki and Shawn Bowers, editors. Scientific and Statistical Database Management: 24th International Conference, SSDBM 2012, Chania, Crete, Greece, June 25–27, 2012. Proceedings, volume 7338 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31234-9 (print), 3-642-31235-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-31235-9.

Pietka:2012:ITB

[980] Ewa Pietka and Jacek Kawa, editors. Information Technologies in Biomedicine: Third International Conference, ITIB 2012, Gliwice, Poland, June 11–13, 2012. Proceedings, volume 7339 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31195-4 (print), 3-642-31196-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-31196-3.

Elmoataz:2012:ISP

[981] Abderrahim Elmoataz, Driss Mammass, Olivier Lezoray, Fathallah Nouboud, and Driss Aboutajdine, editors. Image and Signal Processing: 5th International Conference, ICISP 2012, Agadir, Morocco, June 28–30, 2012. Proceedings, volume 7340 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31253-5 (print), 3-642-31254-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-31254-0.

Bao:2012:ACN

[982] Feng Bao, Pierangela Samarati, and Jianying Zhou, editors. Applied Cryptography and Network Security: 10th International Conference, ACNS 2012, Singapore, June 26–29, 2012. Proceedings, volume 7341 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31283-7 (print), 3-642-31284-5 (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL

http://www.springerlink.com/content/978-3-642-31284-7.

Gibbons:2012:MPC

[983] Jeremy Gibbons and Pablo Nogueira, editors. Mathematics of Program Construction: 11th International Conference, MPC 2012, Madrid, Spain, June 25–27, 2012. Proceedings, volume 7342 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31112-1 (print), 3-642-31113-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.

springerlink.com/content/978-3-642-31113-0.

Dieste:2012:PFS

[984] Oscar Dieste, Andreas Jedlitschka, and Natalia Juristo, editors. Product-Focused Software Process Improvement: 13th International Conference, PROFES 2012, Madrid, Spain, June 13–15, 2012 Proceedings, volume 7343 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31062-1 (print), 3-642-31063-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-642-
31063-8.

Katzenbeisser:2012:TTC

[985] Stefan Katzenbeisser, Edgar Weippl, L. Jean Camp, Melanie Volkamer, Mike Reiter, and Xinwen Zhang, editors. Trust and Trustworthy Computing: 5th International Conference, TRUST 2012, Vienna, Austria, June 13–15, 2012. Proceedings, volume 7344 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30920-8 (print), 3-642-30921-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-30921-2.

Jiang:2012:ARA

[986] He Jiang, Wei Ding, Moonis Ali, and Xindong Wu, editors. Advanced Research in Applied Artificial Intelligence: 25th International Conference on Industrial Engineering and Other Applications of Applied Intelligent Systems, IEA/AIE 2012, Dalian, China, June 9–12, 2012. Proceedings, volume 7345 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31086-9 (print), 3-642-31087-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-642-
31087-4.

Kothe:2012:ADG

[987] Ullrich Köthe, Annick Montanvert, and Pierre Soille, editors. Applications of Discrete Geometry and Mathematical Morphology: First International Workshop, WADGMM 2010, Istanbul, Turkey, August 22, 2010, Revised Selected Papers, volume 7346 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32312-X (print), 3-642-32313-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-32313-3.

Haddad:2012:ATP

[988] Serge Haddad and Lucia Pomello, editors. Application and Theory of Petri Nets: 33rd International Conference, PETRI NETS 2012, Hamburg, Germany, June 25–29, 2012. Proceedings, volume 7347 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31130-X (print), 3-642-31131-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-31131-4.

Bodenreider:2012:DIL

[989] Olivier Bodenreider and Bastien Rance, editors. Data Integration in the Life Sciences: 8th International Conference, DILS 2012, College Park, MD, USA, June 28–29, 2012. Proceedings, volume 7348 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31039-7 (print), 3-642-31040-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-31040-9.

Vallecillo:2012:MFA

[990] Antonio Vallecillo, Juha-Pekka Tolvanen, Ekkart Kindler, Harald Störrle, and Dimitris Kolovos, editors. Modelling Foundations and Applications: 8th European Conference, ECMFA 2012, Kgs. Lyngby, Denmark, July 2– 5, 2012. Proceedings, volume 7349 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31490-2 (print), 3-642-31491-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-31491-9.